A Complete Bibliography of Scalable Computing: Practice and Experience

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

08 October 2019 Version 1.11

Title word cross-reference

3 [BDD18, MG19]. n [MRC13]. P [CI13, PSB05, LKHH05]. π [Ahm08].

-Calculus [Ahm08]. -Dimensional [MG19]. -grade [LKHH05]. -Graph [PSB05]. -manifolds [MRC13].

.NET [Woo07].

1-Output [Saq07]. 1st [PG11].

2 [Che10b]. 2nd [PVP11a, PVP11b].

3 [Che10a, RSS^{+}09].

4-Input [Saq07]. 4-Input/1-Output [Saq07].

5L [Mly09].

61131 [RSS^{+}09]. 61131-3 [RSS^{+}09].

978 [Che10b, Che10a]. 978-0-12-381472-2 [Che10b]. 978-0-13-138768-3 [Che10a].

= [SR18b].

ABD [AR09]. Abstraction [NTH11]. Abundance [CT10]. ABVE [Opr14]. ABVE-construct [Opr14]. Accelerating [CHTE15, Oro07]. Acceleration [BDD18]. Access [BCR^{+}11, HKR^{+}11, PM11c, PM11b]. according [RSS^{+}09]. Accountant [SBB^{+}07]. Achieve [FMB07]. Acid [KMMP18]. Acquaintance [PCRC07]. across [FKSP^{+}13]. ACSys2009 [NFB10]. Action [ND19]. Active [BP12, BD14]. Actuators [MBSC09]. Ad [BDY08, Doc09, DSX09, DBX09, MG19, PM05, VK19]. Ad-hoc [MG19]. Adaptation [KKL05, BAP06, GB09, MP11]. adapting [Var15]. Adaptive [ABJ05, CH15, FMR14, GTB05, HTC05, HML05, MB14, NAIC11, PBT08, Pou15, SBH05]. added [ADVM07]. Adjustable [HM07]. Administration [DM11]. ADMIRE [HST^{+}10]. Admission [YBY^{+}18]. Advanced [MS08]. Advances [DBX09, KM18]. Aerosols [DKK^{+}18]. Affine [SLK17]. Afpac [BAP06]. age [KP12]. Agent [AWLM07, ACD^{+}07, AV14, AN10, Ben13, Cab06, DPG^{+}15, EP12, FNCC17, FPSN15, GGKP05, HPS06, LEP15, MMNs12, OdGWB06, PS07, PZ12, RA06, RP05, RCW07, SMGC08, TRI08, VP07, Zha07, ACPF14, AGP^{+}13, BKG^{+}12, CPM11, GPD^{+}09, JDK17, MPG13, NKS14, Opr14, PM11c, QLC06, RF14, SCKM07]. Agent-Based [Ben13, AV14, PS07, RCW07, VP07, AGP^{+}13, MPG13, Opr14]. Agent-Client [GGKP05]. agent-team-based [GPD^{+}09]. Agents [ADVM07, BI14, MJZ06, MOB06, Mor06, Mus13, OOM^{+}06, VM10, VAM^{+}10, GKP06, RKF13]. Aggregate [JVS12]. Aggregation [FVB05]. Agile [KP12, SSE^{+}15]. Agility [APFCM12]. Aglets [QLC06]. Agreement [BSBB19, MOB06]. agreements [KW11]. Aho [OB19]. Airborne [Kor08]. AIX [Mly09]. Algebra [ANdC17, Bro08, RLA09, BCGH05]. Algebraic [MR09]. Algorithm [AA19, ABDT16, Azi12, BTGV07, BKH10, DDKK18, FB12, EHHC18, HDA18, HTC05, KGS19, KA19, MDA19, OB19, PS08, Pop11, SR18c, ZSJF15]. Algorithmic [AT10]. Algorithms [BG09, BGadH06, CM09, IRTK18, Kho07, ME08, PBP06, RNGG07, Ris18]. Alias [Woo05]. Alignment [EDS^{+}15]. Allocation [HMS17, SNKS19, MFN14, Mly09]. AMIGM [VK19]. among [GTL09]. Analysis [AB09b, AEKF16, BCR^{+}11, BKH10, CG15, EW07, FA18, HR19, JGP07, LS18, LHS07, MK17, OSSS17, PMH^{+}16, RNGG07, RWL07, SS18, Saq07, SK17, SPPGP16, TNL^{+}12, Woo05, WTLS09, dB06, LBP^{+}07, RSM07]. Analytical [AGG^{+}18, BLHK13, MG19]. Analytics [KS19]. Android [SK17]. Anharmonic [KMMP18]. Animal [VK19]. Anomaly [FB12, HU19, HDA18]. Ant [PCS^{+}11b]. Ant-inspired [PCS^{+}11b]. API [Bro08, Lou17a]. APP [TRI08]. Applicability [PSB05]. Application [CDRV16, DdSGP10, HHML06, LBP^{+}07, LKHH05, PDMP10, SSS08, WKS15, ZOBC13, CCD^{+}10, MPC09, ZDB^{+}13]. application-driven [CCD^{+}10]. Applications [AHS15, AG15, BI14, Bar10, Ben18, BG19, Ben13, BDL06, Cza11, DLT^{+}10, DHR17, FMR14, GGG^{+}18, GB12, KY19, KAG^{+}10, LLS07, MBSC09, QJZS05, Raj18, RFB^{+}14, RTV15, DMR^{+}16, RPM^{+}08, SSS08, SR18a, SK17, SGJN19, SGJ19, STA16, TNL^{+}12, VCV08, VAM^{+}10, WTLS09, BLRW09, CB07, MCCC10, dMRGJ07, Lee07]. Applied [BKL^{+}12, KSF11, Saq07]. applying [BMD10]. Approach [AGK10, Ben13, BKP^{+}08, CLM^{+}08, DCLR17, DPCA11, DFN16, GMB14, JDK17, KY19, KmWH10, KMMP18, KCD^{+}18, LEP15, Mar06, OF10, PCRC07, PCS^{+}11a, RCW07, SGJ19, TNL^{+}12, BSST13, CCD^{+}10, TV05, Che10b, Kho07]. Approaches [CG15, JVS12, Pet16, PAD11, SY16, PV14]. Approximate [WZS06]. AQsort [LTS16]. ARank [JDK17]. architectural [AHZ14]. Architecture [BZ11, FNC16, IPD16, JK17, JAKP05, LRNP06, MMNs12, PS08, PBT08, GPS14a, MFN14, MRdT07]. Architectures [AFGZ16, HYM^{+}08, VLNP^{+}16]. Area [HTC05, LLS07, PD10]. Area-Time [LLS07]. arising [AR09]. Array [LTS16, VM10, TOC18]. Arrays [Oro07]. Art [TNL^{+}12, SK17]. Artificial [Kho08, Kul09]. ARTIST [MKA^{+}14]. Aspects [BL07, Lou05, Lou06, Lou17b, MRC13, Pet07]. Assesing [ZC19]. Assessment [APFCM12, JGP07, JBW^{+}15, SK17]. Astronomical [JCCC06]. Asymmetric [BdFW^{+}09]. Asynchronous [Bro08]. Atmospheric [KMMP18]. Atom [KCD^{+}18]. Atom-centered [KCD^{+}18]. Attack [DN19]. Authenticated [BSBB19]. Authentication [AA19, PSSK19, SGG18]. Auto [SGJN19]. Auto-Scaling [SGJN19]. AutoAdmin [YBY^{+}18]. Automata [Abd06]. Automated [BCR^{+}11, BCFP13, RKF13, RF14]. Automatic [DM11, GHV^{+}12, HSB^{+}10, PXB09, PCS^{+}11b, SPPGP16, YBW^{+}17, YRL17, YBY^{+}18]. Autonomic [FNCC17]. Autonomous [EHHC18]. Autonomously [Var15]. Autonomy [HM07]. Availability [BLHK13]. Available [MG19]. AVX2 [OB19]. Aware [Ben18, MBSS17, OGF10, SBH05, SR18a, SPR10, SGJ19, ZSJF15, BSST13, DBGG19, MLO^{+}19, MR09, SY16]. Azure [HMS17].

BABD [AR09]. bacterial [BKG^{+}12]. bag [MPC09]. Balanced [KGS19, SNKS19]. Balancer [RCG16]. Balancing [BG09, FBP15, IDU11, KM08, KM19, OAT^{+}09, PVP06, PGD^{+}11, QJZS05, SNP^{+}10, BKL^{+}12, CB07]. Bandwidth [MG19]. Bargain [RF14]. Base [CMUM09]. Based [ASS08, AK19, Bar10, Ben13, BT12, CLM^{+}08, DCLR17, GMP05, HU19, Ism12, JDK17, KY19, KKKK16, KA19, MOB06, ND19, NSJ^{+}19, NDGL12, OF10, PSSK19, RP05, DMR^{+}16, RPP06, SMH06a, SH08, Var15, VRK19, Zha07, dB06, AHZ14, AASB08, ACD^{+}07, Alo18, AGPAK17, AV14, ACPF14, AN10, ANG^{+}18, AGP^{+}13, ADVM07, APFCM12, BLRW09, BCFP13, BEEY10, BDD18, BTGV07, BKG^{+}12, BR06, BKH10, Cab06, CDRV16, Cza11, Cza12, DdSGP10, DLT^{+}10, DBGG19, DPCA11, FNC16, FBLF11, GPD^{+}09, GMB14, GTB05, GB09, HDA18, HKV05, JBW^{+}15, JK16, KD19, LLS07, LEP15, MDA19, MRM^{+}16, MB14, MPG13, MP11, Mor06, MCCC10, MRdT07, Opr14, PS07, Pet06, PZ12, PM11b, Pop11, RKF13, RCW07, SE16, SSM19, SL14, SV19, SDK^{+}10, VP07, VAM^{+}10, vNMKB05]. Basic [RLA09]. BDI [MZdCML07]. Bed [AMS06]. BeesyCluster [Cza11]. Behavior [LMPQ09, LS11]. Behaviour [OGF10]. Beliefs [LS11]. Benchmarking [ARV13, HHML06]. Benchmarks [ARV13]. Between [BZ09]. Beyond [ASBW11, LB15]. Big [IRTK18, IPD16, KS19, SSBO18]. Binary [CCD^{+}09]. Bio [Kul09]. Bio-Inspired [Kul09]. Bioengineering [STA16]. Bioinformatics [Gu07, HHML06, LBP^{+}07]. biologically [SBB^{+}07]. Biology [AHS15]. biomedical [FAW^{+}16]. Biomedicine [CC15]. Biomolecular [VP07]. Blockchain [SV19]. Blocks [LS18]. Blood [GAM16]. Blue [KHV11]. boarding [FKSP^{+}13]. Bonding [NSJ^{+}19]. Book [Ahm06, Ahm08, Che10b, Che10a, Doc09, Kho07, Kho08, Kul09, Lan07, Lan08, Lee07, Mar06, Oro07, Woo07]. bootstrapping [BMD10]. Boundary [CM09]. bounded [GMW05]. BPEL [BCFP13]. BPEL-based [BCFP13]. BPM [KP12]. Broadcasting [BDaS10]. Brokering [AV14, AV15, AGP^{+}13]. BSP [Mar06]. BSPlib [Lou17a]. BSPlib-style [Lou17a]. Buffering [BCS07]. Build [JRML07]. Building [BdSI15, LS11, Ran05]. Bulk [Gav05, Lou17a]. Bulk-Synchronous [Gav05]. Bus [dB06]. Business [FBLF11, SSS08, VPPL12, RF14]. BVPs [AR09].

C [LLS07]. C-based [LLS07]. Cache [MC18, SBH05]. CAFL [MDA19]. Calculations [ANdC17]. Calculus [Ahm08, Rah06]. Capability [LH17]. Carbon [SE11]. carbonates [BFZK18]. Care [KKKK16]. Carrier [AGK10, VM10]. Carry [CG15]. Case [BKP^{+}08, DO05, KMMP18, LCV13, SST08, VLKJ17, FKSP^{+}13]. Cases [MBSC09, YBI^{+}09]. Catalogues [MRC13]. CC [SR18b]. Cell [GAM16]. centered [KCD^{+}18]. Centers [JK17]. Centre [APC^{+}13]. centred [BSST13]. centric [ZDB^{+}13]. Ceph [MM16]. Certification [Kor08]. CFD [FMR14, PVP06]. Chain [APFCM12]. chains [BSST13]. Challenges [GGJGT^{+}12, GA15, HAA19, MAJLMJ17, Pet05, Raj18, RP05, SBL^{+}11, ZC19, Cza12]. Changes [MM16]. Channel [MLO^{+}19]. Characterization [MBSC09]. Checkpointing [KKKK16, RNGG07, SY16, MB14]. Chemistry [KMMP18]. Cheque [SV19]. Chicken [CJ07]. Chip [MLO^{+}19, MR09]. Chips [FRD^{+}08]. Chord [BKH10, GMB14, GMB14]. Chord-based [BKH10, GMB14]. Churn [LVMV10]. Churn-Tolerant [LVMV10]. Circuits [IDU11]. civil [Opr14]. Class [WZS06]. Classifying [FY15]. Clearing [SV19]. Client [GGKP05]. Climate [DKK^{+}18, GGG^{+}18]. Climatology [KM18]. Clock [BDD18]. closeness [MPG13]. Cloud [Alo18, AV14, AV15, ARV13, AEKF16, APFCM12, BG19, Ben13, BBG17, BLHK13, CC15, CDRV16, CG15, DN19, DFN16, FCPV13, FY15, GA15, HMS17, IRTK18, JDK17, JK17, JBW^{+}15, KY19, KGSN19, KGS19, KM19, KSSP16, KA19, LH17, MMD^{+}15, MM16, PDMP10, Raj18, RFB^{+}14, RTV15, SS18, SGG18, SY16, SGJN19, SGJ19, SR18c, SR18b, SD19, VRK19, VCFP13, WMVP14, ZDB^{+}13, Cza12, MFN14, MB14, CGBG15, CFMN12, GGJGT^{+}12, GVJ11, HKR^{+}11, KP12, MDA12, NTH11, NDGL12, PP11, Pet12, SBL^{+}11, TNL^{+}12]. Cloud-based [JBW^{+}15, Cza12]. Cloudflow [FAW^{+}16]. cloudification [MKA^{+}14]. Cloudlet [SR18c]. CloudLightning [DFN16]. Clouds [BdSI15, BMD10, DMP13, VPPL12, FKSP^{+}13, PV14, ASBW11, ZD11]. Cluster [BLHK13, BDY08]. Clustering [KY19, KGS19, MDA19, SR18a]. Clusters [HGS17, QJZS05, RLA09, WTLS09, YBY^{+}18]. CMP [WTLS09]. co [TOC18]. co-array [TOC18]. Code [BR06, HSB^{+}10, RPP06]. Codes [PVP06]. Coding [MS08]. Cognitive [BPB^{+}12]. Coherence [JBW^{+}15]. coil [RSM07]. coiled [RSM07]. coiled-coil [RSM07]. Collaboration [FBLF11, CA18, GMPS14]. CollaborationBus [GM10]. Collaborations [CBA^{+}14]. Collaborative [IDU11]. Collective [CH15, Pou15, WKS15]. Combination [Son09]. Combining [SSE^{+}15]. Commerce [ZOBC13]. commodity [HDSY07]. Communication [Bar10, GS17, GGKP05, Kor08, RWL07, SY16, SPSK17, Zha07, BKG^{+}12]. Communication-aware [SY16]. Communications [Bar10, DLT^{+}10]. Communities [Ran05, BKG^{+}12, SDK^{+}10]. community [VE18]. Comparative [CHTE15, RSM07]. compared [HDSY07]. Comparison [EDS^{+}15, Woo07, TOC18]. Compeer [PM05]. Complex [DHR17, GM10, Kho08, LPJ16, VLZAP17]. Complexity [LS07, Saq07]. Component [BZ11, DLT^{+}10, BAP06, GB09]. Component-based [DLT^{+}10, GB09]. Components [BP12]. Composite [GB12, JVS12]. Composition [Cab06, DCLR17, PAD11, PCS^{+}11b, PCS^{+}11a]. comprehensive [GGJGT^{+}12].

Computation [CS07, Mar06, Oro07, YBW^{+}17]. Computational [BZ09, EW07, MRC13, RLA09, DMR^{+}16, SW05, XB09]. Computationally [HM07]. Computations [Pap08b, Pet05, Stp06, AASB08, CCD^{+}09]. Compute [Abd06, Cza11]. Computer [HGS17, Lan07, PM11a]. Computers [Mat17, TWC16]. Computing [AAVS07, ASBW11, APFCM12, BdFW^{+}09, BGadH06, CC15, DdSGP10, DBBR19, DPCA11, ETR11, FNCC17, FCPV13, GA15, GS17, GVJ11, GMP05, GM10, HPBG05, HMS17, IRTK18, JDK17, JK17, JAKP05, KY19, KGSN19, KGS19, Kow06, LPJ16, Lee07, Lud07, NRP19, Oro07, PVP06, Pet06, Pet07, PCMF08, PDMP10, PP11, PB08, PD09, Pou15, PPC19, PBT19, RCG16, Ris18, RSM^{+}08, SY16, SBL^{+}11, SBBS17, SNKS19, SGJ19, STA16, SR18c, SR18b, TL05, VLNP^{+}16, VLZAP17, VM10, Woo05, dB06, BDRH07, DLB^{+}10, KW11, LBP^{+}07, MPC09, MB14, MRdT07, NTH11, SDZ14]. Computing-based [APFCM12]. concept [PM11b]. Concerning [Pet05]. concurrency [GMW05]. Concurrent [Mus13, Ols17]. Condensed [JM08]. Condition [WMVP14]. Condition-Event [WMVP14]. Conditioning [CM09]. Configurations [GM10]. Confined [MBSC09]. Conflicting [SE16]. Congestion [SPR10]. Connectivity [DSX09]. considerations [AGP^{+}13]. Considering [MG19, SLK17, YBI^{+}09]. Consistency [RNGG07, BAP06]. Consolidating [APC^{+}13]. Constrained [AT10, JMQ^{+}10]. Constraint [ASS08, BCG15]. Constraint-Based [ASS08]. constraints [MPN12]. construct [Opr14]. Constructs [LCV13]. Content [Bar10, FR08, Hos05, KD19, RPP06, SSE^{+}15]. Content-Based [Bar10, RPP06, KD19]. Context [BSST13, FVB05, OGF10, OF10].

Context-Aware [OGF10, BSST13]. Continuous [NRS^{+}05, BD14]. Control [BCR^{+}11, BP16, BT12, BKL^{+}12, GHV^{+}12, MBSC09, PM11c, QJZS05, YBY^{+}18, BD14, PM11b]. controllers [RSS^{+}09]. Convolutional [MSS17]. cooperation [BKG^{+}12]. Cooperative [CJ07, BSST13]. coordinated [MB14]. Coordination [HPBG05]. Coprocessor [BB18]. coprocessors [DLB^{+}10]. Corasick [OB19]. CORBA [GTB05]. CORBA-based [GTB05]. Core [PB08, VLNP^{+}16, YRL17, HYM^{+}08]. cores [MS08]. Corporate [VCV08]. correcting [MRdT07]. Corrective [GHV^{+}12]. Correctly [Ols17]. Cost [RTV15]. Coterie [FMB07]. Course [Alo18]. CPS [SPSK17]. CPU [BB18]. CPU-MIC [BB18]. Creating [GM10]. Criteria [RNGG07]. Crowd [PPC19]. Cryptosystem [KD19]. CUDA [SK10, Che10a]. CUDAs [KMM15]. Cultural [CA18, KM18]. Current [AFGZ16, WKS15]. Custom [ARV13]. Cyber [GS17, LH17, MAJLMJ17, YBW^{+}17]. Cyber-Physical [GS17, MAJLMJ17, YBW^{+}17]. Cycles [NRS^{+}05].

D [BDD18]. Data

[AT10, APC^{+}13, ABJ05, AN10, ANG^{+}18, AGG^{+}18, CT10, CHTE15, CG15, DBGG19, DS18, DM11, ETR11, FT05, HST^{+}10, HGS17, HHK^{+}05, HKV05, IRTK18, IPD16, JK17, KGSN19, KKL05, KS19, KD19, Lan08, LMPQ09, MK17, MLO^{+}19, MSS17, MRM^{+}16, MWWW11, PIP^{+}15, SS18, SGG18, SSBO18, SLK17, SR18a, SJ07, TL05, YBW^{+}17, CC14, GMPS14, GBS18, SBH05]. Data-Flows [MLO^{+}19]. Data-Intensive [TL05]. Data-Stream [AN10]. Database [Hos05]. Datacenters [IRTK18]. Datasets [HDA18, SH08]. Datastores [CFMN12]. David [Che10b]. debugging [TV05]. decentralized [CC14]. Decision [AV14]. Declarative [BCR^{+}11]. decomposition [HDSY07]. Dedicated [NRS^{+}05]. Deep [MSS17, WH08]. DEEPG [BM11]. Default [MJZ06]. Defined [AK19]. Degree [SR18c]. Delay [HR19]. Delay-Tolerant [HR19]. Delegation [FKSP^{+}13]. deliver [ADVM07]. delivery [NKS14]. demand [MPC09]. Denial [DN19]. Density [HDA18, KCD^{+}18]. Density-based [HDA18]. Dependability [PGD^{+}11]. Dependable [DPC09, DPCA11]. Dependencies [ZLKK16, GTL09]. dependent [KCD^{+}18]. Deployment [BBG17, MPTG09]. Derivatives [BFZK18]. Derived [BCS07]. Deriving [VCV08]. Describing [PCRC07]. Description [HU19]. Design [BKH10, FR07, OSSS17]. Designing [ZOBC13]. Designs [Saq07]. Desktop [SE11]. Desktop-Grids [SE11]. Detailed [HU19]. Detection [DN19, HU19, HDA18, MSS17, MRdT07, VE18]. Developing [BP12, RFB^{+}14]. Development [ARV13, CDRV16, GPD^{+}09, LKHH05]. Developments [PB08]. Device [GAM16, PMH^{+}16]. devices [BD14, dMRGJ07]. Devoted [AB09a]. Dew [RCG16]. DHR [WS07]. DHR-Trees [WS07]. diabetic [LBP^{+}07]. Diagnostic [PSB05]. DICOM [GBS18]. different [GTL09]. Differential [YRL17]. Difficulty [Alo18]. Digital [CA18, DEBM12, KM18]. Dilemma [AFGZ16, CLM^{+}08, CJ07]. Dimensional [HDA18, MG19]. Directions [PP11, SGJN19, GPD^{+}09]. Directory [Mor06]. Disasters [SH08]. Discovering [FR08]. Discovery [BM11, BDY08, DS18, GMB14, Hos05, JAKP05, LEP15, LRNP06, SR18c]. Discrete [CS07]. Disk [CMUM09]. Distance [ND19]. distortion [MRdT07]. Distributed [APPR10, AV14, AV15, AG15, BBJMR09, BDL06, BCG15, BT12, BKL^{+}12, BP12, CS10, CHTE15, DdSGP10, DCLR17, DN19, DPC09, FT05, GP08, HST^{+}10, Hos05, HSB^{+}10, IDU11, KME17, KSSP16, LMPQ09, Lee07, MC18, Mat17, NKKV05, NRS^{+}05, Pet07, PM11c, PD09, PD10, PGD^{+}11, Pop11, RWL07, RPP06, RvSV16, STA16, SNP^{+}10, SJ07, SH06, WS07, WTV05, YBW^{+}17, Zal09, ZD11, Cza12, BS05]. Distributed-Memory [Mat17]. Distribution [ETR11, HTC05, VRK19]. Documents [MVV08]. Domain [BBM15, KSF11]. Down [PCRC07]. Download [BBJMR09]. DPSIW [DBGG19]. DR [BBG17]. DR-SWDF [BBG17]. DREAMCLIMATE [VIJ^{+}18]. Driven [GRY^{+}09, KS19, MDA12, CCD^{+}10]. Dropouts [LS07]. drug [NKS14]. Dual [BM11]. during [BAP06, TV05]. Dust [DKK^{+}18]. Dynamic [Alo18, BdFW^{+}09, BT12, CDS05, DSX09, FMB07, FPSN15, Hos05, KKKK16, KMMP18, MK17, SNKS19, YBY^{+}18, ZL06, CB07, CC14, Mly09]. Dynamically [BBG17, HYM^{+}08]. Dynamics [AAVS07, ANG^{+}18, BFZK18].

e-business [RF14]. E-CAFL [MDA19]. e-Cheque [SV19]. e-Commerce [ZOBC13]. E-DPSIW-FCA [DBGG19]. e-Health [BP16, SG16]. E-Infrastructures [KM18]. e-Services [DM11]. e-System [DM11]. E.V.E. [FS05]. EA [Ben18]. Early [RTV15]. Edge [DBBR19, NRP19, PBT19, PD09]. Editing [GM10]. Editorial [EW07, Kow06, Lud07, NRP19, Pap05, Pap08a, Pet05, Rah06, Ran05, Sed07, SH06, dMRGJ07]. Editors [RS05]. Educational [PBT08]. Edward [Che10a]. EEG [JK16]. Effect [DKK^{+}18, MPN12]. effectiveness [ACPF14]. Efficiency [DDL^{+}15, FCPV13, SE11, YBI^{+}09, YRL17]. Efficient [AK19, ANdC17, BDaS10, CCI11, EHHC18, GLD06, HYM^{+}08, KME17, Mat17, MBSS17, Sen10, SK19, SSM19, SL14, vNMKB05, PXB09]. EGEE [HHML06]. Eigenvalue [BTGV07]. EKG [DDKK18].

Elastic [SD19]. Electromagnetic [BTGV07]. Electronic [JVS12]. Element [TOWC15, TOC18]. Elements [EW07]. Elevant [KMMP18]. ELL [NRLT13]. ELL-i [NRLT13]. Embedded [GRY^{+}09, KSF11, LMPQ09, MPTG09, VLKJ17, GB09]. Emergent [OGF10, XB09]. Empirical [SMH06a]. Enable [JVS12, DPT13]. enabled [KS19, SDK^{+}10]. Enabling [CA18, FNCC17, Hos05, MRM^{+}16, MDA12, NRP19, FAW^{+}16, CBA^{+}14]. Encryption [SGG18]. End [SPR10]. End-to-End [SPR10]. Energy [ANdC17, DDL^{+}15, DBGG19, EHHC18, MBSS17, SSM19, YRL17, ZSJF15]. Energy-Aware [ZSJF15]. Energy-Efficient [EHHC18]. Enforcing [BAP06]. Engine [FP10, RPP06]. Engineering [AHZ14, GRY^{+}09, MDA12, SSE^{+}15, TNL^{+}12, Opr14, SDK^{+}10]. Enhanced [MDA19, SGG18, YRL17]. Enhancing [SCKM07]. Enol [BFZK18]. Enol-carbonates [BFZK18]. Ensemble [ND19]. Enterprise [BKP^{+}08, Woo07]. Enterprises [SST08]. entreprise [Opr14]. Environment [AEKF16, BGadH06, BKL^{+}12, CS10, CHTE15, DS18, FY15, GGG^{+}18, GM10, HPBG05, KY19, KSSP16, LdDK06, MMNs12, NRS^{+}05, PBP06, RWL07, SS18, SGG18, SD19, VRK19, ADD07, GC18, KW11, LKHH05, RF14, RSS^{+}09]. Environments [FP10, KME17, MFPL18, OdGWB06, PVP06, PGD^{+}11, SMGC08, SW05, WMVP14, SDZ14]. Equations [KM08]. equipment [BD14]. equivalent [AGP^{+}13]. Era [BKP^{+}08, KGSN19]. ERCTP [SPR10]. ERP [SCKM07]. Establishing [MPG13, SK19]. Estimation [Alo18, LCV13, LLS07, MS08, MG19]. Evaluating [BCGH05, SSBO18, VCFP13]. Evaluation [BR10, BB18, DPC09, GB12, Rah06, YBI^{+}09, ACPF14]. Event [CS07, WMVP14, YBI^{+}09, BLRW09].

Evolution [JMQ^{+}10, YRL17].

Evolutionary [LCV13]. Exact [IRTK18]. Example [Che10a, SK10, WH08]. Excellent [KM18]. Exchange [OGF10]. Executable [Pom05]. Executed [PVP06]. Execution [ARV13, BKL^{+}12, HML05, MK17, MRM^{+}16, SS18]. expandable [ADD07]. Experience [DMP13]. Experiences [CCD^{+}09]. Experimental [BB18, Pet06]. Experiments [GHJ^{+}08]. Expert [Ahm06, RNGG07]. Exploitation [BPB^{+}12, FVB05, NRS^{+}05]. Exploiting [CT10, MJZ06]. Exploration [ABDT16]. Exploratory [SST08]. Exploring [VM10]. Exposing [DFN16]. Extending [GMW05, JBW^{+}15]. Extensible [RvSV16]. Extension [BZ11, CDS05, PSB05]. External [Gav05]. Extreme [PIP^{+}15].

Facilitating [BMD10]. Facilities [GVJ11]. Factor [PSSK19, VRK19]. Factorisation [BB18]. Factorizations [AB09b]. Failures [BS05]. Family [Rah06]. Fariness [SDZ14]. Fast [JMQ^{+}10, AGK10]. faster [FAW^{+}16]. FastFix [GHV^{+}12]. Fault [FBP15, EHHC18, GLD06, IDU11, JAKP05, Mor06, RPM^{+}08]. Fault-Tolerant [EHHC18, GLD06, JAKP05, Mor06]. FCA [DBGG19]. FCA-based [DBGG19]. Federated [KSSP16, MRM^{+}16, MFPL18]. federation [FKSP^{+}13]. Feedback [QJZS05, AHZ14]. FEM [DHR17]. Fiber [dB06]. Field [Oro07]. Field-Programmable [Oro07]. File [CCI11, HTC05, LVMV10, MRM^{+}16, NRP19, PD09, PD10, VCFP13]. Filtering [AWLM07, FY15]. Finding [PS08]. Finite [TOWC15, TOC18]. Firefly [PCS^{+}11a]. Firefly-inspired [PCS^{+}11a]. First [Mar09]. fixed [NRLT13]. Flexible [CCI11, JK17]. Flocking [WKS15]. Flow [BP16, CHTE15, MK17, TWC16, TOWC15, TOC18]. Flows [CHTE15, MLO^{+}19]. Fog [DBBR19, KS19, NRP19, PBT19, SNKS19, SD19].

Fog-Cloud [SD19]. Folksonomy [VCV08]. Footprint [SE11]. Footprints [LS18]. Foraging [ZSJF15]. Forecasting [CDS05]. Form [PXB09, VCV08]. Formal [GRY^{+}09, GA15, GAM16, NKS14, PMH^{+}16, Rah06, SPSK17]. Formalization [AHS15]. Format [RPR12]. Formats [SL14]. Formic [KMMP18]. Forming [DKK^{+}18]. Formula [LCV13]. Formulation [ASS08]. FORTRAN [TOC18]. forward [SBB^{+}07]. forward-time [SBB^{+}07]. Forwarding [BDaS10, Mor06, NSJ^{+}19]. foster [GMPS14]. Foundational [PS07]. FPGA [HDSY07, LLS07]. FPGAs [Saq07]. Fragmentation [VCFP13]. Framework [Alo18, ARV13, APFCM12, BPB^{+}12, BCR^{+}11, Ben18, BBG17, BMD10, DPC09, ER18, GBS17, HU19, HML05, LH17, MBSS17, MOB06, OAT^{+}09, PCS^{+}11b, SV19, ZL06, CFMN12, GGJGT^{+}12, GB09, MKA^{+}14, MP11, QLC06]. Frameworks [CG15, IPD16, SSBO18]. Freenet [DO05]. Frequency [BBM15]. Frontiers [EW07]. Functional [EW07, GMW05]. Fusion [CHTE15]. Future [BPB^{+}12, NRP19, SGJN19, dMRGJ07]. Fuzzy [ASS08, MDA19].

G [OPM06]. G2 [OPM06]. Game [CJ07]. Games [LS07]. Gate [Oro07]. Gateway [BR10]. Gathering [Ols17]. Gene [KHV11]. Gene/P [KHV11]. General [BR10, Che10a, SK10]. General-Purpose [Che10a, SK10]. Generalized [GPS14b]. generated [MPN12]. Generating [GBS17]. Generation [HSB^{+}10, MRC13, SPPGP16]. Generative [OdGWB06]. Generic [DHR17]. Genetic [PBP06]. genetics [SBB^{+}07]. Genome [EW07, RSM07]. Genome-scale [EW07]. Genome-wide [RSM07]. Genomic [CG15]. Genomics [CHTE15, EW07]. Geographically [KM19]. Geometric [PM11a]. Geometry [DMR^{+}16]. Georgia [DKK^{+}18]. Gladwell [AB09a]. Global [BT12, BKL^{+}12, FBP15, JAKP05, MRM^{+}16, ZD11, CS07]. Go [OOM^{+}06]. Goodput [YBI^{+}09]. Gossip [Pop11]. Governance [CFMN12, KS19]. GPN [AMS06]. GPU [Che10a, BDD18, GC18, JK16, KMM15, SK10]. GPU-based [BDD18, JK16]. grade [LKHH05]. Graph [DCLR17, DSX09, OF10, PSB05]. Graph-Based [OF10]. Graphs [Alo18, JM08]. Gravity [FB12]. Great [AMS06]. Green [PPC19, SE11]. Grid [AGK10, FMR14, AGP^{+}13, HPBG05, AASB08, ATML06, ABJ05, Azi12, BM11, BEEY10, CMUM09, CS10, GPD^{+}09, GVJ11, HKV05, JCCC06, KAG^{+}10, KKL05, Kow06, LVMV10, LHS07, PDMP10, PP11, Ran05, RCW07, RPM^{+}08, SW05, SDK^{+}10, vNMKB05]. Grid-based [HKV05]. Grid-enabled [SDK^{+}10]. Grids [NAIC11, RP05, SE11, CC14, CT10, HML05, JRML07, JAKP05, Pap08b, Pet05, Stp06, vLAV05]. GridWay [HML05]. Group [BSBB19, MCCC10, VK19]. Group-based [MCCC10]. Grouping [SR18a]. Guest [RS05]. GVT [CS07].

Hadoop [LB15, YBY^{+}18]. Hands [Che10b, KmWH10]. Hands-on [Che10b, KmWH10]. Haskell [ATML06]. Having [SR18a]. HBaseSI [ZD11]. Healing [GHV^{+}12, FNC16]. Health [BP16, KKKK16, SG16, SE16, CGBG15]. Healthcare [CC15, DBBR19, MBSC09]. Heap [BM11]. Helpful [LS11]. Heritage [CA18, KM18]. Heterogeneity [ATML06]. Heterogeneous [HU19, HGS17, KAG^{+}10, OdGWB06, PVP06, RLA09, SPR10, ZLKK16]. HeteroPBLAS [RLA09]. Heuristic [IRTK18, PVP06]. Heuristics [JMQ^{+}10]. Hiding [KD19]. Hierarchical [HHK^{+}05, SR18a]. High [Bar10, BL07, BDRH07, DLB^{+}10, FMB07, GRY^{+}09, HDA18, HMS17, LPJ16, Lou05, Lou06, Lou17b, Lud07, Pom05, PD09, PPC19, VLZAP17, VAM^{+}10, Woo05, KW11, SDZ14]. High-Level [BL07, Lou05, Lou06, Lou17b, Pom05]. High-Performance [GRY^{+}09, PPC19, Woo05, KW11]. Historical [dB06]. Hoc [BDY08, Doc09, DBX09, PM05, DSX09, MG19, VK19]. Homogeneous [MDA19]. Horizontal [RCG16]. HPC [ANdC17, BCS07, DFN16, FB12, MRC13, RTV15, VLKJ17]. HPC-Embedded [VLKJ17]. HTTP [DN19]. Huge [TOWC15]. Human [PXB09]. Human-intuitive [PXB09]. Hwu [Che10b]. Hybrid [Ben13, BB18, CM09, KGS19, KHV11, PCS^{+}11a, SGG18, VRK19, ZDB^{+}13]. Hyperspectral [SPPGP16].

I/O [MC18, QJZS05]. I/O- [QJZS05]. IaaS [BdSI15, BCFP13]. Ian [AB09a]. IBM [Mly09]. IDC’2009 [NFB10]. Identification [AK19, EW07, RSM07]. Idle [NRS^{+}05]. IEC [RSS^{+}09]. Image [Bar10, GPS14a, JCCC06, MC18]. Images [BBM15]. Impact [BG09, DO05, DSX09, HMS17, HR19, ME08, MM16, SE11]. Implementation [Abd06, BTGV07, BR10, HYM^{+}08, LRNP06, Mat17, OB19, RCG16]. Implementing [BDL06, CS10, AHZ14]. implicit [Jay09]. Improve [FCPV13, RPR12, DMR^{+}16]. Improvement [AAVS07, PMH^{+}16]. Improving [MFPL18]. In-Place [LTS16]. Incorporating [MZdCML07]. Indexing [PBT08, SS18, WS07]. Industrial [FNC16]. inexpensive [NRLT13]. InfiniBand [FR07]. Influence [PBT19, Mly09]. Influenza [JGP07]. Information [AWLM07, BP16, JCCC06, KD19, MJZ06, OGF10, SG16, Woo07, GPD^{+}09, GBS18, PM11b]. Informations [PM11c]. Infrastructure [ADVM07, Pet16, SCKM07, YBW^{+}17]. Infrastructures [Cab06, KM18, Ris18, RvSV16]. initial [AGP^{+}13]. initiation [CPV13]. Input [TWC16]. Input/1 [Saq07]. Inspired [FP10, Kul09, VK19, APC^{+}13, PCS^{+}11b, PCS^{+}11a]. Instead [Pap05]. Integrated [ANG^{+}18, GHJ^{+}08]. Integrating [DHR17]. Integration [Cza12, HST^{+}10, LdDK06, MWWW11, SMGC08]. Intel [Mar09, OB19]. Intelligence [GBS17, Kho08, Kul09, SCKM07]. Intelligent [GP08, GPD^{+}09]. Intensive [Cza11, DBGG19, QJZS05, SR18a, TL05]. Intention [Var15]. Intention-Propagation-Based [Var15]. Interaction [ACD^{+}07]. interactions [MCCC10]. Interactive [ANG^{+}18, MPC09]. Intercommunicating [SD19]. Interconnected [PBP06]. Interest [SR18a]. International [VPPL12]. Internet [BPB^{+}12, BR10, CPV13, DEBM12, GGKP05, GMP05, KSF11, MFN14, SSS08, ZC19]. Internet-Based [GMP05]. Interoperability [FBLF11, PMH^{+}16]. Interprocessors [DDKK18]. Intrinsic [SLK17]. Introduction [APPR10, AB09a, AG15, BI14, BG19, BL07, BDRH07, CBA^{+}14, CH15, CGBG15, Che10a, DdSGP10, DPG^{+}15, FB13, FKK08a, For14a, For14b, Fri13, GP08, GS17, GMP05, Gu07, HAA19, HPS06, KM18, Lan07, Lan08, LPJ16, Lou05, Lou06, Lou17b, NFB10, NZ13, NKKV05, NRG13, ÕS12, PS07, Pap08b, Pet07, PDMP10, PG11, PP11, PZ12, Pet16, PB08, RA06, Raj18, RS05, SG16, SK10, SBBS17, STA16, Stp06, Tud09, VLZAP17, VPPL12, XB09, Zal09]. Intrusion [HU19]. intuitive [PXB09]. Inventory [AGPAK17]. Inverse [WZS06]. Inversion [FB12]. Investigation [BD14, BLHK13]. Invocation [Ols17]. IoT [BG19]. Irinotecan [KCD^{+}18]. IRISGrid [HHML06]. IRISGrid/EGEE [HHML06]. ISBN [Che10b, Che10a]. ISBN-13 [Che10a]. Isolation [ZD11, BD14]. Issue [APPR10, AG15, BI14, BG19, CBA^{+}14, CH15, CGBG15, DdSGP10, DPG^{+}15, FB13, For14a, For14b, Fri13, GS17, HAA19, KM18, LPJ16, Lou17b, NRP19, NZ13, NRG13, ÕS12, PG11, PP11, PZ12, Pet16, Raj18, Ris18, SG16, SBBS17, STA16, VLZAP17, VPPL12, AB09a, BL07, BDRH07, FKK08a, GP08, GMP05, Gu07, HPS06, Lou05, Lou06, NFB10, NKKV05, PS07, Pap08b, Pet07, PDMP10, PB08, RA06, Stp06, Tud09, XB09, Zal09]. Issues [KGSN19, RP05]. Iterated [CJ07].

J2EE [Woo07]. JADE [FNCC17]. Jadex [BP12]. Jason [Che10a]. Java [BCS07, BR06, BGadH06, RWL07, WTV05, Woo05, vNMKB05]. Java-based [BR06, vNMKB05]. JikesRVM [QLC06]. Job [MRM^{+}16]. Jobs [CCD^{+}10, KKL05, OPM06, MPC09]. Join [Mus13]. Joint [HHML06]. JR [Ols17]. JuxMem [ABJ05].

Kalray [VLKJ17]. Kandrot [Che10a]. Key [BSBB19, VRK19, BSST13]. Kirk [Che10b]. Klaim [BDL06]. Knowledge [AK19, BKP^{+}08, GHJ^{+}08, SST08]. Kutta [Jay09].

Lanczos [BTGV07]. Lanczos-based [BTGV07]. Languages [Lan07, Rah06, GMW05]. Large [BDD18, CMUM09, HHK^{+}05, HSB^{+}10, JCCC06, JGP07, JRML07, LSTD14, Pap08b, Pet07, PGD^{+}11, Pop11, Stp06, ZSJF15, AASB08, TOC18]. Large-Scale [CMUM09, JCCC06, JGP07, LSTD14, Pet07, AASB08]. Latency [Bar10, ME08]. Layer [BCS07, HYM^{+}08, SPR10, NTH11]. LEACH [MBSS17]. Learning [PBT08, KD19]. Level [BL07, Lou05, Lou06, Lou17b, PSB05, Pom05, TL05, CB07, KW11, ZLKK16]. Leveraging [QLC06]. Library [FS05]. Life [KM18, WKS15]. Lifecycle [CFMN12].

Lightweight [VCV08]. like [CCD^{+}09]. Linear [ANdC17, Bro08, MR09, RLA09, WZS06, AR09]. Link [BDY08]. Link-cluster [BDY08]. links [MPN12]. List [FSS08]. Literary [SK17]. Lithuania [GVJ11]. Load [BG09, BKL^{+}12, FBP15, IDU11, KM08, KM19, KA19, OAT^{+}09, PVP06, PGD^{+}11, SNP^{+}10, SR18c, CB07, GTL09]. Load-Balancing [PGD^{+}11]. Locality [SLK17, SR18a]. Localization [Son09]. Loci [ZL06]. Lock [ME08]. Logic [ACD^{+}07, MDA19]. Logic-based [ACD^{+}07]. Loops [SLK17, ZLKK16, AHZ14]. low [HDSY07]. LU [BB18, HDSY07].

M [SE16]. M-health [SE16]. Machine [KD19, PXB09, YRL17]. Machine-efficient [PXB09]. Machine-learning [KD19]. Machines [EDS^{+}15, Lan07]. Maintenance [GHV^{+}12]. Majority [BEEY10]. Maker [AFGZ16]. Malicious [BEEY10]. Malware [GBS17]. Management [APPR10, AGPAK17, BPB^{+}12, BKP^{+}08, FCPV13, FY15, MRM^{+}16, MFPL18, PM11b, Pop11, SJ07, VAM^{+}10, ZL06, GGJGT^{+}12, ZDB^{+}13]. Manager [YBW^{+}17]. Managing [ATML06, MMD^{+}15, EP12]. MANETs [BDaS10]. manifolds [MRC13]. Many [VLNP^{+}16, YRL17]. Many-Core [VLNP^{+}16, YRL17]. Many-Task [VLNP^{+}16]. Maple [BS05]. Mapping [BB18]. MapReduce [FAW^{+}16]. Marketplace [Azi12]. Marketplaces [JVS12]. Markov [SR18a]. MAS [OGF10]. Massively [Che10b, KmWH10, BLRW09]. Matching [ASS08, BDD18, CCD^{+}10, KMM15, OF10]. Mathematical [LRNP06]. Matlab [Sha09]. Matrices [LSTD14, LS18, SL14, WMVP14]. Matrix [GPS14b, KCD^{+}18, RPR12]. Maturity [LH17]. Maximum [PS08]. MC [SR18b]. MCC [SR18b]. Measurement [HPBG05]. Measurements [RSM^{+}08]. Measures [FY15]. Mechanism [HKR^{+}11, QJZS05, Mly09]. Mechanisms [RWL07, SGG18]. Media [Bar10]. mediation [BMD10]. Medical [AEKF16, GBS18]. Medigrid [SH08]. Medium [SST08]. mei [Che10b]. Membership [DSX09]. Membrane [Pet06]. Memetic [JMQ^{+}10]. Memory [Gav05, KME17, LS18, Mat17, ME08, QJZS05, ZL06]. Memory-Intensive [QJZS05]. Mendel [SBB^{+}07]. MERPSYS [GC18]. Mesh [CM09, CCD^{+}09, Sen10, GPS14a]. Mesh-like [CCD^{+}09]. mesh-of-tori [GPS14a]. Meshes [GLD06]. Message [FCPV13, VLKJ17, NSJ^{+}19, TV05]. Message-Passing [VLKJ17, TV05]. Meta [SBH05]. Meta-data [SBH05]. Metacomputing [MHPP05, PM05]. meteorology [LKHH05]. Method [KKKK16, MVV08, SSE^{+}15, VE18]. Methodology [Rah06, VCV08, MKA^{+}14]. Methods [BDD18, FMR14, GA15, Kul09, SPSK17, Jay09]. Metric [BDD18, PGD^{+}11]. Metrics [LH17, MSST08, OAT^{+}09, YBI^{+}09]. MIC [BB18]. mice [LBP^{+}07]. Microfluidic [GAM16]. microprocessor [HDSY07]. Middleware [AMS06, AN10, FCPV13, GTB05, KSSP16, KSF11, MAJLMJ17, PMH^{+}16, TL05, GPD^{+}09, DPT13]. Migration [DM11, OdGWB06, VK19, MKA^{+}14]. Migrations [DDKK18]. MIMD [PS08]. Minimization [BBJMR09]. Minimum [PS08]. Mining [FT05, HST^{+}10, Lan08]. Mirroring [GPD^{+}09]. ML [Gav05, Lou17a, SMH06b]. Mobile [ADVM07, BM11, BDL06, BR06, MOB06, Mor06, OdGWB06, Raj18, RPP06, SMGC08, SR18c, SR18b, Son09, VAM^{+}10, VK19, YBI^{+}09, Ahm08, QLC06]. Mobiles [DMP13]. Mobility [DBX09, HPS06, HR19, MG19, SH06, VK19, GKP06, QLC06]. Modbus [SPSK17]. Modbus/TCP [SPSK17]. Model [AN10, GRY^{+}09, GPS14b, GHJ^{+}08, HGS17, LH17, MDA12, OSSS17, PXB09, VK19, BKG^{+}12, Opr14, RKF13]. Model-Driven [GRY^{+}09]. Modeling [BI14, BFZK18, DLT^{+}10, PSB05, Rah06, VP07]. Modelling [EP12, FPSN15, GC18, NKS14]. Models [BZ09, DBX09, HR19, JVS12, SH08, dB06]. Modern [ME08]. modernization [MKA^{+}14]. Modified [OSSS17]. Modular [Pou15]. Molecular [AHS15, AAVS07, ANG^{+}18, BFZK18, KMMP18]. Monitor [BKH10]. Monitoring [BT12, IPD16, Pop11, RCW07]. Montoring [PBT19]. MooreCube [JK17]. mOSAIC [ARV13, MDA12]. Mosaicking [JCCC06]. Motion [MS08, MRdT07]. motion-detection [MRdT07]. Move [FKK08a, FKK08b]. Movement [WKS15]. MPI [AASB08, MC18, MK17, RPM^{+}08, TOC18, Mar06]. MPICH [OPM06]. MPICH-G2 [OPM06]. MSMAS [EP12]. Multi [AN10, Ben13, BCG15, DPG^{+}15, EP12, FMR14, HYM^{+}08, JDK17, JK17, JMQ^{+}10, KSSP16, LTS16, LEP15, MPC09, MMNs12, MS08, PSSK19, PB08, PM11c, SCKM07, VCFP13, Zha07, ZD11, CPM11, HDSY07, NKS14, ZDB^{+}13]. Multi-Agent [AN10, DPG^{+}15, LEP15, MMNs12, Zha07, JDK17, PM11c, SCKM07, CPM11, NKS14]. Multi-application [MPC09]. Multi-Array [LTS16]. Multi-Cloud [Ben13, ZDB^{+}13]. Multi-Core [PB08, HYM^{+}08]. Multi-cores [MS08]. Multi-Factor [PSSK19]. multi-FPGA [HDSY07]. Multi-Grid [FMR14]. Multi-Objective [BCG15, JMQ^{+}10, KSSP16]. multi-PaaS [ZDB^{+}13]. Multi-port [JK17]. Multi-Provider [VCFP13]. Multi-row [ZD11]. Multiagent [PCRC07]. Multicast [JAKP05]. Multicore [DDL^{+}15, FRD^{+}08, Mar09, MR09, SLK17].

Multicriteria [AT10]. Multidimensional [WS07]. Multilevel [WZS06]. Multimedia [FR08]. Multiple [AGP^{+}13, Bar10, IPD16, KMM15]. Multiplication [GPS14b]. Multiplier [Saq07]. Multiprocessors [ME08]. multispecies [BKG^{+}12]. Multithreaded [MR09]. multiuser [MCCC10]. Multivariate [LdDK06]. Museum [ER18]. MUSKEL [ADD07]. mutable [GMW05].

nano [NKS14]. nano-robotic [NKS14]. NAT [BR10]. NAT-Gateway [BR10]. Natural [SH08]. Nature [FP10]. naval [ACPF14]. Navier [KM08]. Navigation [Var15, WH08]. Negotiation [MOB06, BMD10, RKF13, RF14]. Negotiations [BMD10]. Nested [SLK17, ZLKK16]. Nested-Loops [SLK17]. Net [ND19]. nets [Pom05]. Network [APPR10, AA19, Alo18, AK19, AMS06, CLM^{+}08, DdSGP10, Gun08, HPBG05, HAA19, HTC05, LS07, MLO^{+}19, NSJ^{+}19, OSSS17, SBBS17, Son09, SW05, YBI^{+}09, VE18]. Network-based [DdSGP10]. Networked [KSF11]. Networking [GS17]. Networks [Abd06, BZ11, BCS07, BDY08, Doc09, DSX09, DBX09, FVB05, JRML07, KA19, LEP15, MBSC09, MSS17, MBSS17, MG19, RSM^{+}08, Sen10, SK19, SSM19, SV15, VK19, Zal09, MPN12]. Neural [MSS17]. NoC [FRD^{+}08, VLKJ17]. Node [AA19, PD09]. Non [NRS^{+}05]. Non-Dedicated [NRS^{+}05]. Normative [FPSN15]. Note [Ols17]. Novel [EW07]. NTRU [AA19]. Numerical [AB09b, AR09, Bro08, KM08, MR09]. NVRAM [MC18].

O [MC18]. O- [QJZS05]. Object [FS05, GPS14b, JM08, PBT08]. Objective [BCG15, JMQ^{+}10, KSSP16]. Objects [DEBM12]. Observation [Zha07]. Observation-Based [Zha07]. ODEs [Sha09]. offers [AGP^{+}13]. Offline [AT10]. Oil [Ism12]. OMTSE [PS08, PM11a]. on-boarding [FKSP^{+}13]. on-Chip [MLO^{+}19]. on-demand [MPC09]. One [LB15]. onto [LLS07]. Ontology [KS19, MJZ06, MDA12, VCV08]. Open [BdSI15, KGSN19, MSST08, RSS^{+}09, SSE^{+}15]. OpenFOAM [FA18]. OpenMP [Ben18, BB18, LTS16]. OpenStack [BCFP13]. Operating [JRML07]. operational [ACPF14]. Opportunistic [AA19, HAA19, NSJ^{+}19, SK19]. opportunities [PV14]. Optical [JBW^{+}15, dB06]. Optimal [PCS^{+}11a, SR18c, BD14]. Optimality [LS07]. Optimization [AT10, BCG15, GB12, KD19, KA19, WTLS09, GBS18]. Optimizations [SPPGP16]. Optimized [CHTE15, HDA18, KY19, RLA09]. Optimizing [Bar10, ETR11, MFN14]. Optoelectronic [PM11a]. OrcFS [CCI11]. OrganiK [BKP^{+}08]. Organisation [PCRC07]. Organising [SSE^{+}15]. Organization [CCI11]. Organized [SV15]. Organizing [APC^{+}13]. Oriented [FS05, FCPV13, FP10, GPS14b, JM08, MWWW11, PCMF08]. OSGi [DPT13]. OSyRIS [FP10]. Output [Saq07]. overlapping [VE18]. Overlay [BM11, DCLR17, JRML07, Pou15]. Overview [WKS15, GPD^{+}09].

P [KHV11]. P2P [BKH10, SV15, WS07]. PaaS [GGJGT^{+}12, ZDB^{+}13]. Package [TRI08]. Panorama [Pet12]. paper [Mus13]. Papers [NFB10, PG11, PVP11a, PVP11b, VPPL12]. Paradigm [PSB05, DPT13]. Parallel [Abd06, ATML06, ABDT16, AB09b, AR09, BL07, BTGV07, BBM15, BGadH06, BZ09, CDS05, CS07, Che10b, CI13, DdSGP10, DDL^{+}15, FA18, FMR14, Gav05, HGS17, Ism12, KME17, KAG^{+}10, Kho07, KmWH10, KHV11, LdDK06, Lou05, Lou06, Lou17a, Lou17b, MC18, Mar06, Mat17, MS08, NKKV05, Pet06, Pom05, VE18, DMR^{+}16, SMH06a, SMH06b, SMGC08, TWC16, WZS06, WTLS09, dB06, BLRW09, BCGH05, BAP06, RLA09]. Parallelization [Cza11, HGS17, SLK17]. Parameter [MM16]. Parameterised [BP16]. Parameters [TWC16]. Parrot [TL05]. Partial [MK17]. Participants [ETR11]. Particle [KA19]. Partitioned [KM19, LS18]. Partitioning [CHTE15, FSS08]. Pass [FSS08]. Passing [VLKJ17, TV05]. Pattern [EW07, PSSK19]. Pattern-Based [PSSK19]. Patterns [Mus13, AHZ14]. PC [CMUM09]. Peer [DO05]. Peer-to-Peer [DO05]. Perfectly [SLK17, ZLKK16]. perform [LBP^{+}07]. Performance [Ben18, CDS05, EDS^{+}15, FR07, GRY^{+}09, HPBG05, HMS17, HDSY07, LPJ16, LCV13, LRNP06, Lud07, Mar09, MM16, OSSS17, PXB09, PD09, PD10, PPC19, Rah06, RWL07, RPR12, RSM^{+}08, SMH06a, SPPGP16, VLZAP17, VAM^{+}10, Woo05, WTLS09, YBI^{+}09, BCGH05, BDRH07, DLB^{+}10, KW11, SDZ14]. periodic [CC14]. Personalized [AWLM07]. Perspective [DdSGP10, MRdT07]. Petri [Pom05]. Phi [BB18]. Phylogenetic [JGP07]. Physical [GS17, HYM^{+}08, MAJLMJ17, YBW^{+}17]. Pi [Rah06]. Pi-Calculus [Rah06]. PIA [AWLM07]. Pion [LdDK06]. pipeline [BCGH05]. pipeline-structured [BCGH05]. pipelines [FAW^{+}16]. PL [MRC13]. Place [LTS16]. Placement [DBGG19, IRTK18, KGSN19, KKL05, SR18a, SD19]. Plains [AMS06]. Planning [MZdCML07, TRI08]. Platform [ABJ05, ANG^{+}18, AGG^{+}18, BB18, DEBM12, FNCC17, FBLF11, GGG^{+}18, GB12, HMS17, IPD16, JBW^{+}15, KA19, LB15, MMD^{+}15, MHPP05, YRL17, NRLT13]. Platforms [BdFW^{+}09, SNKS19, LLS07]. Player [LS07, BSST13]. player-centred [BSST13].

Point [CM09]. Pointers [Mor06]. Policies [AEKF16, SE16]. Policy [NDGL12, SNKS19]. Policy-Based [NDGL12]. population [SBB^{+}07]. Port [Bar10, JK17]. Portability [PV14]. portal [AASB08, SDK^{+}10]. Portfolio [KD19]. Porting [LLS07]. position [Mus13]. PosoMAS [SSE^{+}15]. Possible [VCV08]. Potential [FRD^{+}08]. Power [JCCC06, MR09, NAIC11, SE11, HDSY07]. Power-aware [MR09]. Practical [BL07, Lou05, Lou06, Lou17b, Pet07]. Practices [DPG^{+}15]. Pragmatic [PS07]. Pravah [BP16]. Precise [Son09]. Preconditioners [WZS06]. Preconditioning [Jay09]. Prediction [CCD^{+}09, RTV15, SMH06a, SNP^{+}10, Var15]. present [dMRGJ07]. Preserving [Sen10]. Price [JVS12]. Pricing [LS07]. Principles [Ahm06, DPG^{+}15, Lee07]. Prioritization [SE16]. Prisoner [CJ07]. Privacy [SE16, Sen10, SK17]. Privacy-Preserving [Sen10]. Proactive [Zha07]. Probes [HPBG05]. Problem [BTGV07, GGKP05, HSB^{+}10, Kho08, LdDK06]. Problems [AT10, CM09, LPJ16, PM11a, PSB05, VLZAP17, CPM11]. Process [HMS17, DMR^{+}16, SPPGP16, BCGH05]. processes [Ahm08]. Processing [AT10, AN10, DCLR17, GHJ^{+}08, JK16, MC18, MRM^{+}16, PIP^{+}15, GC18, GBS18]. Processors [Che10b, HYM^{+}08, KmWH10, Mar09, MR09, SLK17]. production [SDK^{+}10]. Professor [AB09a]. profiles [RKF13]. Profiling [SMH06a]. Profit [SGJ19]. Profit-Aware [SGJ19]. Program [BT12, BKL^{+}12, MK17, SBB^{+}07]. Programmable [Oro07]. Programming [BL07, HGS17, KmWH10, KHV11, Lou05, Lou06, Lou17b, RWL07, SK10, vNMKB05, GMW05, RSS^{+}09, Ahm06, Che10a, Che10b]. Programs [BT12, DDL^{+}15, BCGH05, TV05]. Progressive [HHK^{+}05]. Project [SH08].

Proof [AK19]. Propagation [KCD^{+}18, Var15]. Properties [Gun08, BD14]. Proprietary [BCS07]. Protection [CS10, SV15]. proteins [RSM07]. proteomic [LBP^{+}07]. Protocol [BM11, BDY08, BR10, HPBG05, NAIC11, Sen10, SPR10, SPSK17, VRK19, CPV13]. Protocols [ACD^{+}07, BSBB19, HR19, SSM19, Doc09]. Provenance [KKKK16]. Provider [VCFP13]. providing [TV05]. Provisioning [BCFP13, SGJ19]. Proxy [SBH05]. pSeries [Mly09]. Public [DM11, KM19, SGG18]. Purpose [Che10a, SK10]. PVL [SLK17].

Qinna [GB09]. QoS [MP11]. QoS-based [MP11]. Quality [JBW^{+}15]. Quantum [CS07]. Quasi [AGK10]. Quasi-Random [AGK10]. Quasirandom [KAG^{+}10]. Query [Bar10]. Quo [Kow06].

Rack [LB15]. RAINBOW [FNC16]. Random [AGK10]. randomly [MPN12]. Ranking [JDK17]. Rapid [LLS07]. Reactions [VP07]. Real [ASBW11, DEBM12, DBX09, EHHC18, JK16, LMPQ09, MLO^{+}19, MPTG09, WKS15, Zal09]. Real-Time [EHHC18, JK16, MPTG09, Zal09]. Real-World [DEBM12]. Realistic [DO05, SBB^{+}07]. Reality [HKV05]. Recognition [ND19]. Reconfigurable [AAVS07, BBG17, HYM^{+}08, Oro07, Pou15, BDRH07, dMRGJ07]. Reconfiguration [BdFW^{+}09]. Reducing [DDKK18]. Reductions [KME17]. redundant [GTL09]. Reference [Woo05]. Reference-Set [Woo05]. region [CA18]. Regional [DKK^{+}18]. Regulated [WMVP14]. Relational [CCI11]. Relations [BZ09]. Reliability [PMH^{+}16, SG16, SK19]. Reliable [SPR10]. Reloaded [BCR^{+}11]. Remote [HHK^{+}05]. Replay [BLRW09].

Replay-based [BLRW09]. Replication [NAIC11, YBW^{+}17, CC14]. Repositories [KSSP16]. Representation [FVB05, Woo05]. Reputable [vLAV05]. Reputation [BEEY10]. Reputation-based [BEEY10]. Request [SD19]. Rescheduling [JMQ^{+}10, DMR^{+}16]. Research [DS18, GGG^{+}18, KGSN19, MSST08, MFPL18, RP05, SGJN19, PV14]. Reservation [YBY^{+}18]. Reservoir [Ism12]. Resolving [SE16]. Resource [AT10, BM11, Hos05, MOB06, Pop11, SGJ19, YBY^{+}18, AGP^{+}13, GMW05, GB09, Mly09]. resource-bounded [GMW05]. Resources [ANdC17, Azi12, BEEY10, CCD^{+}10, ETR11, JMQ^{+}10, MFPL18, MFN14, SDK^{+}10]. Results [Ols17]. Retinal [JBW^{+}15]. Retrieval [BDD18, HHK^{+}05]. Review [Ahm06, Ahm08, Che10b, Che10a, DBBR19, Doc09, Kho07, Kho08, Kul09, Lan07, Lan08, Lee07, Mar06, Oro07, Woo07]. RFID [AGPAK17]. Rich [Hos05, SSS08]. Ring [Ism12]. Ring-Based [Ism12]. Rings [BCG15]. Risk [FCPV13]. Risks [BR06]. Robin [SR18c]. robotic [NKS14]. Robust [BTGV07, JAKP05]. Role [Cab06, MHPP05, PM11b]. Role-based [Cab06]. Round [DMR^{+}16, SR18c]. Round-Based [DMR^{+}16]. Route [BDY08]. Routing [GLD06, HR19, MBSS17, Sen10, SK19, SSM19, YBI^{+}09, CPM11]. row [ZD11]. Rule [Pet06]. Rule-based [Pet06]. Rules [SSS08]. Run [KKL05]. Run-time [KKL05]. Runge [Jay09]. Runtime [BKH10, GB09]. RWAPI [FR07].

SAaaS4Mobile [DMP13]. SaaS [ANdC17, BCFP13]. Safe [NRS^{+}05, GB09]. Sanders [Che10a]. Sarcasm [MSS17]. Satin [vNMKB05]. Saturated [TOWC15, TOC18]. SCADA [SPSK17]. Scalability [Ben18, MVV08, PD10, SSBO18, SJ07, TOWC15]. Scalable [AV15, AG15, BKH10, CS07, IPD16, JK17, LTS16, MRM^{+}16, NAIC11, RCG16, Ris18, SBBS17, YBW^{+}17, ZOBC13, LBP^{+}07]. Scale [BDD18, CMUM09, HSB^{+}10, JCCC06, JGP07, JRML07, LSTD14, Pap08b, Pet07, PGD^{+}11, Pop11, Stp06, ZSJF15, AASB08, EW07, Ben18]. SCALE-EA [Ben18]. Scaling [LB15, SGJN19]. Scenarios [WKS15, ZDB^{+}13]. Scheduling [EHHC18, HML05, KY19, KA19, LHS07, MLO^{+}19, NDGL12, SNKS19, SW05, ZLKK16, SDZ14]. Schema [ASS08]. Scheme [AK19]. Schemes [FBP15]. Science [CT10, KM18, Lan07]. Sciences [KM18]. Scientific [BBG17, KY19, KGS19, Mar06, SE11, VAM^{+}10, WTLS09, CB07, SDK^{+}10]. scrambling [GPS14a]. Scrum [SSE^{+}15]. Search [GHJ^{+}08, RPP06]. Searching [CCI11]. Secure [CDRV16, HKR^{+}11, KD19, LVMV10, MBSS17, RFB^{+}14, SBBS17, SPSK17]. Secured [SGG18]. Securing [AK19]. Security [BSBB19, BR06, HAA19, LH17, SGG18, SG16, ZC19]. seed [VE18]. SEEM [CA18, VIJ^{+}18]. Seepage [TWC16]. Selected [CG15, NFB10, PG11, PVP11a, PVP11b, VPPL12]. Selection [Alo18, CM09, KD19, LEP15, VE18]. Selective [BDaS10]. Self [APC^{+}13, CS10, EP12, FNC16, GHV^{+}12, SV15, SSE^{+}15, Var15]. Self-adapting [Var15]. Self-Healing [GHV^{+}12, FNC16]. Self-managing [EP12]. Self-Organising [SSE^{+}15]. Self-Organized [SV15]. Self-Organizing [APC^{+}13]. Self-Protection [CS10]. Semantic [FBLF11, Gun08, KSF11, MMNs12, RP05, MPG13, PZ12, GMB14, PCS^{+}11a]. Semantics [FR08, SMH06a]. Semantics-Based [SMH06a]. Semi [BCG15, BD14]. Semi-Rings [BCG15]. Semiempirical [KCD^{+}18]. Sensing [DMP13, FVB05]. Sensitivity [LHS07, TWC16]. Sensor [BZ11, KA19, MWWW11, MBSS17, OSSS17, SSM19, Son09, YBI^{+}09, DPT13]. Sensors [AN10, MBSC09]. Sentic [KS19]. Sentic-Social [KS19]. Separation [GAM16]. Sequence [EDS^{+}15, HGS17]. Sequential [Kho07, RPM^{+}08]. Serialization [WTV05]. Server [GTB05]. Servers [APC^{+}13, JK17, Mly09]. Service [AV15, BZ11, DN19, DS18, DPCA11, FP10, FBLF11, Hos05, LEP15, LRNP06, MWWW11, MFPL18, Mor06, PCMF08, PAD11, PCS^{+}11b, PCS^{+}11a, PGD^{+}11, Pou15, RCW07, RPP06, RvSV16, SH08, VIJ^{+}18, BMD10, KW11, MP11, DPT13]. Service-Based [SH08, FBLF11, MP11]. Service-Oriented [MWWW11, PCMF08]. Services [BCR^{+}11, CFMN12, Cza11, DM11, GGJGT^{+}12, GMB14, JDK17, JRML07, JVS12, LS11, MMNs12, MPTG09, MDA12, NRP19, NDGL12, Pet12, Pet16, RCG16, VAM^{+}10, ZC19, ADVM07, BSST13, Cza12, DFN16, PZ12, KSF11, PG11, PVP11a, PVP11b]. session [CPV13]. Set [RLA09, Woo05]. Several [AT10, BZ09]. Shape [BDD18]. Shared [MJZ06, ME08, GTL09]. Sharing [ABJ05, GM10, GMPS14]. Short [SLK17]. Short-Vector [SLK17]. Shortcomings [TNL^{+}12]. Signaling [BR10]. Significance [SR18a]. SIMD [AFGZ16, FS05, OB19, PBP06]. Similarity [FY15, HGS17]. Simple [vNMKB05]. Simulation [AGK10, BI14, BKG^{+}12, DHR17, DPC09, DBX09, FA18, GMPS14, GB12, IDU11, SH08, VP07, XB09, ACPF14, GC18, NKS14]. Simulations [AAVS07, ANG^{+}18, BFZK18, CS07, DHR17, TWC16, TOWC15, TOC18]. Simulator [Ism12]. Simulators [CI13]. Simultaneous [JK16, AGP^{+}13]. Single [Bar10, FSS08, MM16]. Single-Pass [FSS08]. SIQRS [OSSS17]. Sized [LS18]. Sizing [LB15]. Skeleton [ND19, ADD07]. Skeletons [Mat17, SMH06b, BCGH05].

SLA [CDRV16]. SLA-based [CDRV16]. Slow [DN19]. Small [SST08, RSS^{+}09, TV05]. small-waiting [TV05]. Smart [AGPAK17, KKKK16, NAIC11]. Smartphone [PPC19]. Snapshot [ZD11]. SOAJA [OAT^{+}09]. SOAs [ADVM07]. SoC [DLB^{+}10]. Social [CLM^{+}08, KS19, LEP15, NSJ^{+}19, ZOBC13, VE18, Gun08]. Societal [BG19]. Societies [FPSN15]. Sociotechnical [BKP^{+}08]. SODAP [SV15]. Soft [JK16]. Software [AK19, BI14, EDS^{+}15, GKP06, GHV^{+}12, HPS06, Kor08, MKA^{+}14, PG11, PVP11a, PVP11b, RA06, RvSV16, SSE^{+}15, AHZ14, PZ12]. Software-as-a-Service [RvSV16]. Solution [Abd06, AR09, KM08, MC18, NRS^{+}05, RPM^{+}08, RvSV16, Sha09, ZDB^{+}13]. Solutions [AGPAK17, AT10, BdSI15, BG19, DS18, LPJ16, SBBS17, VLZAP17, WKS15, ZC19, Cza12, SDK^{+}10]. Solvers [MR09]. Solving [AFGZ16, HSB^{+}10, LdDK06, Kho08]. Some [PM11a]. Sorting [LTS16]. Source [BdSI15]. Space [ABDT16, MG19, SL14, dMRGJ07]. Spaces [MBSC09]. Spark [FAW^{+}16]. Sparse [LSTD14, LS18, RPR12, SL14, WZS06]. Spatial [SS18]. Spatio [ND19]. Spatio-Temporal [ND19]. Special [APPR10, AG15, BI14, BG19, CBA^{+}14, CH15, CGBG15, DdSGP10, DPG^{+}15, FB13, For14a, For14b, Fri13, GS17, HAA19, KM18, LPJ16, Lou17b, NRP19, NZ13, NRG13, ÕS12, PG11, PP11, PZ12, Pet16, Raj18, Ris18, SG16, SBBS17, STA16, VLZAP17, VPPL12, AB09a, BL07, BDRH07, FKK08a, GP08, GMP05, Gu07, HPS06, Lou05, Lou06, NFB10, NKKV05, PS07, Pap08b, Pet07, PDMP10, PB08, RA06, Stp06, Tud09, XB09, Zal09]. Species [KMMP18]. Specification [ACD^{+}07, BCR^{+}11]. Specifications [Pom05]. Spectroscopy [KMMP18, KCD^{+}18]. Speed [MR09].

Speed-up [MR09]. Speedup [FA18]. Spin [ME08]. Spin-Lock [ME08]. SPMV [RPR12]. Stability [LS07]. Standard [SBH05, RSS^{+}09, SMH06b]. Standards [MRM^{+}16]. Standards-based [MRM^{+}16]. State [ABDT16, SK17, TNL^{+}12, GMW05]. State-of-art [SK17]. Statement [Ols17]. States [BT12]. Static [KM19, KMMP18, Woo05]. Stationary [YBI^{+}09]. Stochastic [Abd06]. Stokes [KM08]. stomach [LBP^{+}07]. Stop [BS05]. Storage [BLHK13, HKR^{+}11, MM16, RPR12, SS18, SNP^{+}10, VCFP13, YBW^{+}17, FKSP^{+}13]. Storing [SL14]. Strategies [CJ07, HMS17, Kho08, PMH^{+}16, RF14]. Strategy [DBGG19, GLD06, SD19, AGP^{+}13, CB07, CC14]. Stream [AN10, DLT^{+}10]. Streaming [GTB05]. String [KMM15]. Strong [ZD11, QLC06]. Structure [KD19, SL14, WS07]. Structured [BTGV07, FMB07, Mar06, BCGH05]. Structures [Kho08]. student [GKP06]. Study [BFZK18, DO05, GTL09, JGP07, KMMP18, Mar09, MBSC09, OSSS17, SST08, TWC16, TOWC15, VLKJ17, FKSP^{+}13]. Studying [MVV08, SST08]. style [Lou17a]. Subpopulations [PBP06]. Subprograms [RLA09]. Success [CJ07]. Supercomputers [TOWC15]. Superpeer [SV15]. Supply [APFCM12]. Support [AV14, BCS07, BPB^{+}12, BCFP13, FNCC17, FBLF11, MFPL18, YRL17, GKP06, MPG13, SDK^{+}10]. supporting [ACPF14, CFMN12]. Supportive [ABJ05]. Surveillance [BdFW^{+}09, JGP07, MWWW11]. Survey [AGPAK17, CI13, KGSN19, PD10, SGJN19, SR18b]. Sustainable [PPC19]. SVM [MVV08]. Swarm [APC^{+}13, KA19, VM10]. Swarm-Array [VM10]. Swarm-inspired [APC^{+}13]. SWDF [BBG17]. Switch [LCV13]. Switch-Case [LCV13]. Symbolic [Pet05, PCMF08]. SymGrid [PCMF08]. synchronization [BLRW09]. Synchronous [Gav05, Lou17a]. System [AWLM07, ACD^{+}07, BBJMR09, CI13, CCI11, DM11, JRML07, JK16, KKKK16, LVMV10, MRM^{+}16, MWWW11, PSSK19, Pet06, PSB05, PD09, PBT08, RNGG07, SNP^{+}10, VCFP13, ACPF14, AGP^{+}13, CPM11, GPD^{+}09, HDSY07, MPG13, NKS14, PM11b, SCKM07, PD09]. System-Level [PSB05]. Systems [APPR10, Ahm06, AGPAK17, BR06, BdFW^{+}09, BKH10, BP12, CH15, DO05, DPG^{+}15, DPC09, EP12, ETR11, FCPV13, GRY^{+}09, GP08, GA15, GS17, EHHC18, KSF11, LMPQ09, LHS07, MRC13, MZdCML07, MAJLMJ17, MPTG09, NKKV05, PS07, PCRC07, PB08, Pom05, PM11c, PD10, Pop11, Pou15, SG16, SMGC08, SJ07, SSE^{+}15, SH06, WZS06, WS07, XB09, YBW^{+}17, Zal09, ZSJF15, ZLKK16, AR09, Cza12, GTL09, GBS18, GB09, MP11, MRdT07, PZ12, Woo07].

Table [AFGZ16]. Target [SPPGP16]. Task [BdFW^{+}09, KGS19, SNKS19, VLNP^{+}16, SDZ14]. Tasks [SD19, ACPF14]. Taxonomy [KGSN19, PAD11, SJ07]. TCP [SPSK17]. team [GPD^{+}09]. Teamwork [Zha07, GMPS14]. Technique [APC^{+}13, KM19, LLS07, MK17, NSJ^{+}19]. Techniques [KD19, Son09]. Technologies [CBA^{+}14, Doc09, NRP19, KP12, Kul09]. Technology [AWLM07, HST^{+}10, RA06, RPP06]. TELOS [DCLR17]. Temperature [KCD^{+}18]. Temperature-dependent [KCD^{+}18]. Temporal [ND19]. temporary [MPN12]. Term [Alo18]. Test [AMS06]. Testbed [HHML06]. Testing [CI13, GA15, LCV13, SPSK17]. Text [MSS17, MVV08]. Their [SH08, BFZK18, PDMP10]. Theoretic [GHV^{+}12]. Theoretical [KMMP18]. Theories [Kul09]. Theory [Lan07, Ahm08].

Things [KSF11, NRLT13, BPB^{+}12, CPV13, DEBM12, MFN14, ZC19]. Thread [BB18]. Threads [WTV05]. Threat [GBS17]. Threaths [ZC19]. Three [Mar09, ZLKK16]. Three-level [ZLKK16]. Throughput [AAVS07, FMB07]. Tiling [ZLKK16]. Time [BBJMR09, CS07, EHHC18, JK16, LMPQ09, LLS07, MLO^{+}19, MPTG09, RPR12, Zal09, KKL05, MB14, SBB^{+}07, TV05, CS07]. time-based [MB14]. Timed [Pom05]. timestamps [BLRW09]. tissue [LBP^{+}07]. Tolerance [BEEY10, FBP15, RPM^{+}08]. Tolerant [EHHC18, GLD06, HR19, JAKP05, LVMV10, Mor06]. Tolerating [BS05]. Tomographies [JBW^{+}15]. Tool [CDS05, MMD^{+}15, CB07]. Toolkit [SDK^{+}10]. Top [PCRC07]. Topology [SV15]. Tori [GLD06, GPS14a]. Total [RPR12]. traces [BLRW09]. Trading [Azi12]. Traffic [DBX09]. Transactions [ZD11]. Transfer [AT10, SST08]. Transformation [PXB09]. Transient [Abd06]. Transparent [SY16, TL05]. Transport [AGK10, BR10, SPR10]. travel [MPG13]. Traveling [OOM^{+}06]. Tree [FMB07, KME17, Mat17, SL14]. Tree-based [SL14]. Tree-Structured [FMB07]. Trees [CCD^{+}09, WS07]. Trends [NRP19, SGJN19]. Triangulation [SGJ19]. Trust [CLM^{+}08, FY15, LEP15, MSST08, MBSS17, VRK19]. Trust-based [LEP15]. Trusted [LS11]. Trustlet [MSST08]. TTL [GMB14]. TTL-Chord [GMB14]. Tuning [Ben18, EDS^{+}15, YRL17]. Two [CM09, CB07]. two-level [CB07]. Two-Point [CM09]. Types [BCS07].

UAVs [BdFW^{+}09]. Ubiquitous [GM10]. Ultra [AGK10]. Ultra-fast [AGK10]. Underpinnings [PS07]. Understanding [FPSN15]. Unevenness [Gun08]. Unified [Kho07]. Uniformization [Abd06]. Uniformly [LS18]. Uniformly-Sized [LS18]. Unsupervised [HU19]. Usage [FNC16, Pet06, RTV15, SE11]. use [CPM11]. User [Sen10, TL05, ZDB^{+}13]. user-centric [ZDB^{+}13]. User-Level [TL05]. Users [FY15]. Using [AA19, AAVS07, Bar10, CT10, FR08, HST^{+}10, HGS17, HSB^{+}10, JMQ^{+}10, JRML07, KW11, MK17, MVV08, DMR^{+}16, Alo18, BCG15, CCD^{+}09, ETR11, FA18, HR19, KMM15, KD19, Mar06, MSS17, MKA^{+}14, MDA12, OB19, PMH^{+}16, SLK17, SPSK17, SNKS19, TWC16, TOWC15, VAM^{+}10]. Utility [ASBW11]. Utilization [ANdC17].

Vadis [Kow06]. Validation [GRY^{+}09, SPSK17]. Value [CM09, ADVM07, BSST13]. Variability [BG09]. Variably [TOWC15, TOC18]. Variance [BBJMR09]. Vector [SLK17, YRL17]. Vectorization [SLK17]. Vectorized [Sha09]. Vehicle [CPM11]. Verification [ACD^{+}07, AEKF16, GA15, GAM16]. versus [KMMP18]. VI [VIJ^{+}18]. VI-SEEM [VIJ^{+}18]. via [Cab06, YRL17]. Vibrational [KMMP18, KCD^{+}18]. Video [MWWW11, MS08]. VieSLAF [BMD10]. Vietnam [Lud07]. View [GHV^{+}12]. Vine [SDK^{+}10]. VirMuF [ER18]. Virtual [Azi12, CMUM09, CS07, DS18, ER18, EDS^{+}15, GGG^{+}18, HKV05, MLO^{+}19, MFPL18, CA18, Opr14]. Virtualization [DPCA11, FRD^{+}08, Sed07, Mly09]. Virtualization-based [DPCA11]. Vision [BLHK13]. visual [MRdT07]. Visualization [AGG^{+}18, HHK^{+}05, HKV05, HSB^{+}10, JK16, LSTD14, SH08]. VLSI [IDU11]. VM [SNKS19]. VMI [KSSP16]. VOD [BBJMR09]. VOFS [LVMV10]. Volume [HKV05]. Volunteer [ETR11]. Voting [BEEY10].

waiting [TV05]. Watermarking [BBM15]. way [Lud07]. Weather [AGG^{+}18]. Web [AASB08, ANG^{+}18, BCR^{+}11, BGadH06, FKK08a, FKK08b, GMB14, Gun08, GHJ^{+}08, LS11, MJZ06, MPTG09, PAD11, PCS^{+}11b, PCS^{+}11a, RPP06, SGJN19, SGJ19, WH08]. Web-based [ANG^{+}18]. WebCom [MHPP05, OPM06]. WebCom-G [OPM06]. Weblogs [SST08]. WELSA [PBT08]. Wen [Che10b]. Wen-mei [Che10b]. Wide [FKK08b, HTC05, PD10, RSM07]. Wikipedia [FR08]. WiMAX [HYM^{+}08]. Wind [FA18]. Wireless [DSX09, FVB05, KA19, MBSC09, MBSS17, MG19, OSSS17, Sen10, SSM19, YBI^{+}09]. Within [MHPP05, GPD^{+}09]. Work [CHTE15]. Workflow [BCR^{+}11, DBGG19, FP10, IRTK18, LHS07, Cza12]. Workflows [BBG17, Cza11, JMQ^{+}10, KGS19, MB14]. Workload [BG09, DO05]. Workshop [PG11, PVP11a, PVP11b, VPPL12]. workshops [NFB10]. World [DEBM12, FKK08b]. WS [MOB06]. WS-Agreement [MOB06]. WSN [MDA19, SPR10].

X [BDL06]. X-Klaim [BDL06]. X1.V1 [MMD^{+}15]. X10 [FBP15]. XACML [AEKF16]. Xeon [BB18]. XML [MHPP05]. XNAT [JBW^{+}15].

YARN [YBY^{+}18]. yourSkyG [JCCC06].

Zero [AK19]. Zero-Knowledge [AK19]. Zone [SSM19]. Zone-based [SSM19]. Zoonosis [JGP07]. Zsyntax [AHS15].

References

Abouaroek:2019:NAU

[AA19] Musaeed Abouaroek and Khaleel Ahmad. Node authentication using NTRU algorithm in opportunistic network. Scalable Computing: Practice and Experience, 20(1):83–92, ????

2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1481.

Akzhalova:2008:WPL

[AASB08] Assel Zh. Akzhalova, Daniar Y.

Aizhulov, Galymzhan Seralin, and Gulnar Balakayeva. Web portal for large-scale computations based on Grid and MPI. Scalable Computing: Practice and Experience, 9(2): 135–142, June 2008. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol09/no2/SCPE_9_2_06.pdf;
http://www.scpe.org/vols/
vol09/no2/SCPE_9_2_06.zip.

Alam:2007:TIM

[AAVS07] Sadaf R. Alam, Pratul K. Agarwal, Jeffrey S. Vetter, and Melissa C. Smith. Throughput improvement of molecular dynamics simulations using reconfigurable computing. Scalable Computing: Practice and Experience, 8(4):395– 410, December 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no4/SCPE_8_4_06.pdf;
http://www.scpe.org/vols/
vol08/no4/SCPE_8_4_06.zip.

Amodio:2009:ISI

[AB09a] Pierluigi Amodio and Luigi Brugnano. Introduction to the special issue: Special issue devoted to Professor Ian Gladwell. Scalable Computing: Practice and Experience, 10(4):i–ii, December 2009. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/vols/vol10/
no4/vol10no4introduction.
html.

Amodio:2009:PFN

[AB09b] Pierluigi Amodio and Luigi Brugnano. Parallel factorizations in numerical analysis. Scalable Computing: Practice and Experience, 10(4):385– 396, December 2009. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol10/no4/SCPE_10_4_
05.pdf;   http://www.scpe.
org/vols/vol10/no4/SCPE_
10_4_05.zip.

Abdallah:2006:PIU

[Abd06] Haïscam Abdallah. Parallel implementation of uniformization to compute the transient solution of stochastic automata networks. Scalable Computing: Practice and Experience, 7(2): 53–64, June 2006. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol07/no2/SCPE_7_2_05.pdf;
 http://www.scpe.org/vols/
vol07/no2/SCPE_7_2_05.zip.

Allal:2016:PAS

[ABDT16] Lamia Allal, Ghalem Belalem, Philippe Dhaussy, and Ciprian Teodorov. A parallel algorithm for the state space exploration. Scalable Computing: Practice and Experience, 17(2):129–142, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1161.

Antoniu:2005:JAS

[ABJ05] Gabriel Antoniu, Luc Bougé, and Mathieu Jan. JuxMem: An adaptive supportive platform for data sharing on the Grid. Scalable Computing: Practice and Experience, 6(3):45– 55, September 2005. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol06/no3/SCPE_6_3_05.pdf;
http://www.scpe.org/vols/
vol06/no3/SCPE_6_3_05.zip.

Alberti:2007:SVA

[ACD^{+}07] Marco Alberti, Federico Chesani, Davide Daolio, Marco Gavanelli, Evelina Lamma, Paola Mello, and Paolo Torroni. Specification and verification of agent interaction protocols in a logic-based system. Scalable Computing: Practice and Experience, 8(1):1– 13, March 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no1/SCPE_8_1_01.pdf;
http://www.scpe.org/vols/
vol08/no1/SCPE_8_1_01.zip.

Anghinolfi:2014:SSE

[ACPF14] Davide Anghinolfi, Alberto Capogrosso, Massimo Paolucci, and Perra Francesco. A system supporting the evaluation of the operational effectiveness of naval tasks based on agent simulation. Scalable Computing: Practice and Experience, 15(3):

??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/1016.

Aldinucci:2007:MES

[ADD07] Marco Aldinucci, Marco Danelutto, and Patrizio Dazzi. MUSKEL: an expandable skeleton environment. Scalable Computing: Practice and Experience, 8(4):325–341, December 2007. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol08/no4/
SCPE_8_4_01.pdf;   http:/
/www.scpe.org/vols/vol08/
no4/SCPE_8_4_01.zip.

Aversa:2007:MAB

[ADVM07] Rocco Aversa, Beniamino Di Martino, Salvatore Venticinque, and Nicola Mazzocca. A mobile agents based infrastructure to deliver value added services in SOAs. Scalable Computing: Practice and Experience, 8(3):313–324, September 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no3/SCPE_8_3_09.pdf;
 http://www.scpe.org/vols/
vol08/no3/SCPE_8_3_09.zip.

Ayache:2016:AVX

[AEKF16] Meryeme Ayache, Mohammed Erradi, Ahmed Khoumsi, and Bernd Freisleben. Analysis and verification of XACML policies in a medical cloud environment. Scalable Computing: Practice and Experience, 17(3):189–206, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1180.

Avenel:2016:STM

[AFGZ16] Christophe Avenel, Pierre Fortin, Mourad Gouicem, and Samia Zaidi. Solving the Table Maker’s dilemma on current SIMD architectures. Scalable Computing: Practice and Experience, 17(3):237–250, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1183.

Anderson:2015:ISI

[AG15] David Anderson and Marjan Gusev. Introduction to the special issue on scalable and distributed applications. Scalable Computing: Practice and Experience, 16(2):iii, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1084.

Astsatryan:2018:WDV

[AGG^{+}18] Hrachya Astsatryan, Hayk Grogoryan, Eliza Gyulgyulyan, Anush Hakobyan, Aram Kocharyan, Wahi Narsisian, Vladimir Sahakyan, Yuri Shoukourian, Rita Abrahamyan, Zarmandukht Petrosyan, and Julien Aligon. Weather data visualization and analytical platform. Scalable Computing: Practice and Experience, 19(2):79–86, ???? 2018. CODEN ????

ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1351.

Atanassov:2010:UFC

[AGK10] Emanouil Atanassov, Todor Gurov, and Aneta Karaivanova. Ultra-fast carrier transport simulation on the grid. quasirandom approach. Scalable Computing: Practice and Experience, 11(2):137–147, June 2010. CODEN ???? ISSN 1895-1767. URL http://www.

                                        scpe.org/vols/vol11/no2/
                                        SCPE_11_2_05.pdf;   http:
                                        //www.scpe.org/vols/vol11/
                                        no2/SCPE_11_2_05.zip.

Attaoui:2013:MES

[AGP^{+}13] Naoual Attaoui, Maria Ganzha, Marcin Paprzycki, Katarzyna Wasielewska, and Mohammad Essaaidi. Multiple equivalent simultaneous offers strategy in an agent-based grid resource brokering system — initial considerations. Scalable Computing: Practice and Experience, 14(2):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/index.
                                        php/scpe/article/view/842.

Alwadi:2017:SSR

[AGPAK17] Ali Alwadi, Amjad Gawanmeh, Sazia Parvin, and Jamal N. Al-Karaki. Smart solutions for RFID based inventory management systems: A survey. Scalable Computing: Practice and Experience, 18(4):347–360, ???? 2017. CODEN ????

ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1333.

Ahmad:2006:BRE

[Ahm06]  Raheel Ahmad.  Book re-

view: Expert Systems: Principles and Programming. Scalable Computing: Practice and Experience, 7(4):119, December 2006. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol07/no4/
vol07no4bookreview.html.

Ahmad:2008:BRC

[Ahm08] Raheel Ahmad. Book review:

The π-Calculus: A theory of mobile processes. Scalable Computing: Practice and Experience, 9(2):151–152, June 2008. CODEN ???? ISSN 1895-

1767. URL http://www.scpe.
org/vols/vol09/no2/BR_100.
html.

Ahmad:2015:FZA

[AHS15] Sohaib Ahmad, Osman Hasan, and Umair Siddique. On the formalization of Zsyntax with applications in molecular biology. Scalable Computing: Practice and Experience, 16(1):??, ???? 2015. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/1059.

Abeywickrama:2014:EIS

[AHZ14] Dhaminda B. Abeywickrama, Nicklas Hoch, and Franco Zambonelli. Engineering and implementing software architectural patterns based on feedback loops. Scalable Computing: Practice and Experience, 15(4): ??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/1052.

Alshameri:2019:EZK

[AK19] Hamza Mutaher Alshameri and Pradeep Kumar. An efficient zero-knowledge proof based identification scheme for securing software defined network. Scalable Computing: Practice and Experience, 20(1):181–189, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1473.

Alostad:2018:CBD

[Alo18] Jasem M Alostad. Cloud based dynamic course selection framework using network graphs with term difficulty estimation. Scalable Computing: Practice and Experience, 19(4):361–373, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1425.

Apon:2006:GPN

[AMS06] Amy W. Apon, Gregory E. Monaco, and Gordon K. Springer. The Great Plains Network (GPN) middleware test bed. Scalable Computing: Practice and Experience, 7(3): 95–108, September 2006. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/

vols/vol07/no3/SCPE_7_3_
09.pdf;   http://www.scpe.
org/vols/vol07/no3/SCPE_7_
3_09.zip.

Aritoni:2010:SDS

[AN10] Ovidiu Aritoni and Viorel Negru. Sensors data-stream processing middleware based on multi-agent model. Scalable Computing: Practice and Experience, 11(1):19–32, March 2010. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol11/no1/
SCPE_11_1_02.pdf;   http:
//www.scpe.org/vols/vol11/
no1/SCPE_11_1_02.zip.

Astsatryan:2017:SEE

[ANdC17] Hrachya Astsatryan, Wahi Narsisian, and Georges da Costa. SaaS for energy efficient utilization of HPC resources of linear algebra calculations. Scalable Computing: Practice and Experience, 18(2):145–150, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1286.

Astsatryan:2018:IWB

[ANG^{+}18] Hrachya Astsatryan, Wahi Narsisian, Eliza Gyulgyulyan, Vardan Baghdasaryan, Armen Poghosyan, Yevgeni Mamasakhlisov, and Peter Wittenburg. An integrated Webbased interactive data platform for molecular dynamics simulations. Scalable Computing: Practice and Experience, 19(2):131–138, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1337.

Anghel:2013:SIT

[APC^{+}13] Ionut Anghel, Cristina Bianca Pop, Tudor Cioara, Ioan Salomie, and Iulia Vartic. A swarm-inspired technique for self-organizing and consolidating data centre servers. Scalable Computing: Practice and Experience, 14(2):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/841.

Azevedo:2012:ASC

[APFCM12] Susana Azevedo, Paula Prata, Paulo Fazendeiro, and V. Cruz-Machado. Assessment of supply chain agility in a cloud computing-based framework. Scalable Computing: Practice and Experience, 13(4):??, ???? 2012. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/811.

Abawajy:2010:ISI

[APPR10] Jemal H. Abawajy, Mukaddim Pathan, Al-Sakib Khan Pathan, and Mustafizur Rahman. Introduction to the special issue: Network management in distributed systems. Scalable Computing: Practice and Experience, 11(4):i–ii, December 2010. CODEN ????

ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/664.

Amodio:2009:PNS

[AR09] Pierluigi Amodio and Giuseppe Romanazzi. Parallel numerical solution of ABD and BABD linear systems arising from BVPs. Scalable Computing: Practice and Experience, 10(4): 373–383, December 2009. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol10/no4/SCPE_10_4_
04.pdf;   http://www.scpe.
org/vols/vol10/no4/SCPE_
10_4_04.zip.

Aversano:2013:MBF

[ARV13] Giuseppe Aversano, Massimiliano Rak, and Umberto Villano. The mOSAIC benchmarking framework: Development and execution of custom cloud benchmarks. Scalable Computing: Practice and Experience, 14(1):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/825.

Assel:2011:BCT

[ASBW11] Matthias Assel, Lutz Schubert, Daniel Rubio Bonilla, and Stefan Wesner. Beyond Clouds — towards real utility computing. Scalable Computing: Practice and Experience, 12(2): 179–191, June 2011. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/713.

Algergawy:2008:FCB

[ASS08] Alsayed Algergawy, Eike Schallehn, and Gunter Saake. Fuzzy constraint-based schema matching formulation. Scalable Computing: Practice and Experience, 9(4):303–314, December 2008. CODEN ???? ISSN 1895-1767. URL http://www.

                                        scpe.org/vols/vol09/no4/
                                        SCPE_9_4_07.pdf;   http:/
                                        /www.scpe.org/vols/vol09/
                                        no4/SCPE_9_4_07.zip.

Andreica:2010:ASS

[AT10] Mugurel Ionut Andreica and Nicolae Tapus. Algorithmic solutions for several offline constrained resource processing and data transfer multicriteria optimization problems. Scalable Computing: Practice and Experience, 11(1):1–17, March 2010. CODEN ???? ISSN 1895-1767. URL http://www.

                                        scpe.org/vols/vol11/no1/
                                        SCPE_11_1_01.pdf;   http:
                                        //www.scpe.org/vols/vol11/
                                        no1/SCPE_11_1_01.zip.

AlZain:2006:MHG

[ATML06] A. D. Al Zain, Philip W. Trinder, Greg Michaelson, and Hans-Wolfgang Loidl. Managing heterogeneity in a Grid Parallel Haskell. Scalable Computing: Practice and Experience, 7 (3):9–25, September 2006. CO-DEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/
                                        vols/vol07/no3/SCPE_7_3_
                                        02.pdf;   http://www.scpe.

org/vols/vol07/no3/SCPE_7_
3_02.zip.

Amato:2014:DAB

[AV14] Alba Amato and Salvatore Venticinque. A distributed agentbased decision support for cloud brokering. Scalable Computing: Practice and Experience, 15(1):??, ???? 2014. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/966.

Amato:2015:SDC

[AV15] Alba Amato and Salvatore Venticinque. A scalable and distributed cloud brokering service. Scalable Computing: Practice and Experience, 16(2): 139–152, ???? 2015. CO-DEN ???? ISSN 1895-

1767.  URL https://www.
scpe.org/index.php/scpe/
article/view/1086.

Albayrak:2007:ATP

[AWLM07] Sahin Albayrak, Stefan Wollny, Andreas Lommatzsch, and Dragan Milosevic. Agent technology for personalized information filtering: The PIA system. Scalable Computing: Practice and Experience, 8(1): 29–40, March 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no1/SCPE_8_1_03.pdf;
http://www.scpe.org/vols/
vol08/no1/SCPE_8_1_03.zip.

Aziz:2012:ATG

[Azi12]  Benjamin Aziz.  An algo-

rithm for trading Grid resources in a virtual marketplace. Scalable Computing: Practice and Experience, 13(4): ??, ???? 2012. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/815.

Buisson:2006:AEC

[BAP06] J. Buisson, F. André, and J.- L. Pazat. Afpac: Enforcing consistency during the adaptation of a parallel component. Scalable Computing: Practice and Experience, 7(3):61– 73, September 2006. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol07/no3/SCPE_7_3_06.pdf;
 http://www.scpe.org/vols/
vol07/no3/SCPE_7_3_06.zip.

Barlas:2010:OIC

[Bar10] Gerassimos Barlas. Optimizing image content-based query applications over high latency communication media, using single and multiple port communications. Scalable Computing: Practice and Experience, 11(3):221–237, September 2010. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/656.

Bylina:2018:EEO

[BB18] Beata Bylina and Jaroslaw Bylina. An experimental evaluation of the OpenMP thread mapping for LU factorisation on Xeon Phi coprocessor and on hybrid CPU-MIC platform. Scalable Computing: Practice and Experience, 19(3):259–274, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1373.

Bousselmi:2017:DSD

[BBG17] Khadija Bousselmi, Zaki Brahmi, and Mohamed Mohsen Gammoudi. DR-SWDF: A dynamically reconfigurable framework for scientific workflows deployment in the cloud. Scalable Computing: Practice and Experience, 18(2):177–193, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1289.

Baert:2009:MDT

[BBJMR09] Anne-Elisabeth Baert, Vincent Boudet, Alain Jean-Marie, and Xavier Roche. Minimization of download time variance in a distributed VOD system. Scalable Computing: Practice and Experience, 10(1):75–86, March 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no1/
SCPE_10_1_07.pdf;   http:
//www.scpe.org/vols/vol10/
no1/SCPE_10_1_07.zip.

Bollman:2015:PWI

[BBM15] Dorothy Bollman, Alcibiades Bustillo, and Einstein Morales. Parallel watermarking of images in the frequency domain. Scalable Computing: Practice and Experience, 16(2):205–217, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1090.

Bellavista:2013:TAB

[BCFP13] Paolo Bellavista, Antonio Corradi, Luca Foschini, and Alessandro Pernafini. Towards an automated BPELbased SaaS provisioning support for OpenStack IaaS. Scalable Computing: Practice and Experience, 14(4):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/930.

Billiau:2015:MOD

[BCG15] Graham Billiau, Chee Fon Chang, and Aditya Ghose. Multi-objective distributed constraint optimization using semirings. Scalable Computing: Practice and Experience, 16(4): 403–422, ???? 2015. CO-DEN ???? ISSN 1895-

                                        1767.  URL https://www.
                                        scpe.org/index.php/scpe/
                                        article/view/1130.

Benoit:2005:EPP

[BCGH05] Anne Benoît, Murray Cole, Stephen Gilmore, and Jane Hillston. Evaluating the performance of pipeline-structured parallel programs with skeletons and process algebra. Scalable Computing: Practice and Experience, 6(4):1– 16, December 2005. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol06/no4/SCPE_6_4_01.pdf;
http://www.scpe.org/vols/
vol06/no4/SCPE_6_4_01.zip.

Barletta:2011:WAC

[BCR^{+}11] Michele Barletta, Alberto Calvi, Silvio Ranise, Luca Vigano, and Luca Zanetti. Workflow and access control reloaded: a declarative specification framework for the automated analysis of Web services. Scalable Computing: Practice and Experience, 12(1): 1–20, March 2011. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/684.

Baker:2007:BLS

[BCS07] Mark Baker, Bryan Carpenter, and Aamir Shafi. A buffering layer to support derived types and proprietary networks for Java HPC. Scalable Computing: Practice and Experience, 8(4):343–358, December 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no4/SCPE_8_4_02.pdf;
http://www.scpe.org/vols/
vol08/no4/SCPE_8_4_02.zip.

Basili:2014:IOP

[BD14] Michela Basili and Maurizio De Angelis. Investigation on the optimal properties of semi active control devices with continuous control for equipment isolation. Scalable Computing:

Practice and Experience, 15(4): ??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/1054.

Bein:2010:EBM

[BDaS10] Doina Bein, Ajoy K. Datta, and Balaji ashok Sathyanarayanan. Efficient broadcasting in MANETs by selective forwarding. Scalable Computing: Practice and Experience, 11(1):43–52, March 2010. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol11/no1/SCPE_11_1_
04.pdf;   http://www.scpe.
org/vols/vol11/no1/SCPE_
11_1_04.zip.

Benjelloun:2018:GBA

[BDD18] Mohammed Benjelloun, El Wardani Dadi, and El Mostafa Daoudi. GPU-based acceleration of methods based on clock matching metric for large scale 3D shape retrieval. Scalable Computing: Practice and Experience, 19(1):31–38, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1391.

Binotto:2009:TTD

[BdFW^{+}09] Alécio P. D. Binotto, Edison P. de Freitas, Marco A. Wehrmeister, Carlos E. Pereira, André Stork, and Tony Larsson. Towards task dynamic reconfiguration over asymmetric computing platforms for UAVs surveillance systems. Scalable Computing: Practice and Experience, 10(3):277–289, September 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no3/
SCPE_10_3_05.pdf;   http:
//www.scpe.org/vols/vol10/
no3/SCPE_10_3_05.zip.

Bettini:2006:IMD

[BDL06] Lorenzo Bettini, Rocco De Nicola, and Michele Loreti. Implementing mobile and distributed applications in X-Klaim. Scalable Computing: Practice and Experience, 7(4): 13–35, December 2006. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol07/no4/SCPE_7_4_
02.pdf;   http://www.scpe.
org/vols/vol07/no4/SCPE_7_
4_02.zip.

Bollman:2007:ISI

[BDRH07] Dorothy Bollman, Javier Díaz, and Francisco Rodriguez-Henriquez. Introduction to the special issue: High performance reconfigurable computing. Scalable Computing: Practice and Experience, 8(4):vii–viii, December 2007. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/vols/vol08/
no4/vol08no4introduction2.
html.

Barkat:2015:OSS

[BdSI15] Amine Barkat, Alysson Diniz dos Santos, and Sonia Ikken. Open source solutions for building IaaS clouds.

Scalable Computing: Practice and Experience, 16(2):187–204, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1089.

Bein:2008:LCR

[BDY08] Doina Bein, Ajoy K. Datta, and Shashirekha Yellenki. A link-cluster route discovery protocol for ad hoc networks. Scalable Computing: Practice and Experience, 9(1):21– 28, March 2008. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol09/no1/SCPE_9_1_03.pdf;
                                         http://www.scpe.org/vols/
                                        vol09/no1/SCPE_9_1_03.zip.

Bendahmane:2010:RBM

[BEEY10] Ahmed Bendahmane, Mohammad Essaaidi, Ahmed El Moussaoui, and Ali Younes. Reputation-based majority voting for malicious Grid resources tolerance. Scalable Computing: Practice and Experience, 11(4):385–392, December 2010. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/669.

Benmerzoug:2013:ABA

[Ben13]  Djamel Benmerzoug.  An

agent-based approach for hybrid multi-cloud applications. Scalable Computing: Practice and Experience, 14(2):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/843.

Benedict:2018:SES

[Ben18] Shajulin Benedict. SCALE-EA: A scalability aware performance tuning framework for OpenMP applications. Scalable Computing: Practice and Experience, 19(1):15–30, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1390.

Bigovic:2018:MMD

[BFZK18] Miljan Bigovic, Luka Filipovic, Zarko Zecevic, and Bozo Krstajic. Modeling and molecular dynamics simulations study of enol-carbonates and their derivatives. Scalable Computing: Practice and Experience, 19(2):139–148, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1353.

Beltran:2009:IWV

[BG09] Marta Beltrán and Antonio Guzmán. The impact of workload variability on load balancing algorithms. Scalable Computing: Practice and Experience, 10(2):131–146, June 2009. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol10/no2/SCPE_10_2_
01.pdf;   http://www.scpe.
org/vols/vol10/no2/SCPE_
10_2_01.zip.

Benedict:2019:ISI

[BG19] Shajulin Benedict and Michael Gerndt. Introduction to the special issue on IoT cloud solutions for societal applications. Scalable Computing: Practice and Experience, 20(3): iii, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1323.

Bonorden:2006:WCE

[BGadH06] Olaf Bonorden, Joachim Gehweiler, and Friedhelm Meyer auf der Heide. A Web computing environment for parallel algorithms in Java. Scalable Computing: Practice and Experience, 7(2): 1–14, June 2006. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol07/no2/SCPE_7_2_01.pdf;
                                         http://www.scpe.org/vols/
                                        vol07/no2/SCPE_7_2_01.zip.

Badica:2014:ISI

[BI14] Costin Badica and Mirjana Ivanovic. Introduction to the special issue on modeling and simulation applications of software agents. Scalable Computing: Practice and Experience, 15(3):??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1015.

Bihary:2012:SCC

[BKG^{+}12] Döra Bihary, Ádám Kerényi, Zsolt Gelencsér, Sergiu Netotea, Attila Kertész-Farkas, Vittorio Venturi, and Sándor Pongor. Simulation of communication and cooperation in multispecies bacterial communities with an agent based model. Scalable Computing: Practice and Experience, 13(1): 21–28, March 2012. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/764.

Binzenhofer:2010:DAS

[BKH10] Andreas Binzenhöfer, Gerald Kunzmann, and Robert Henjes. Design and analysis of a scalable algorithm to monitor chord-based P2P systems at runtime. Scalable Computing: Practice and Experience, 11(3):239–249, September 2010. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/657.

Borkowski:2012:DPG

[BKL^{+}12] Janusz Borkowski, Damian Kopański, Eryk Laskowski, Richard Olejnik, and Marek Tudruj. A distributed program global execution control environment applied to load balancing. Scalable Computing: Practice and Experience, 13(3):269– 284, September 2012. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/797.

Bibikas:2008:SAK

[BKP^{+}08] Dimitris Bibikas, Dimitrios Kourtesis, Iraklis Paraskakis, Ansgar Bernardi, Leo Sauermann, Dimitris Apostolou, Gregoris Mentzas, and Ana Cristina Vasconcelos. A sociotechnical approach to knowledge management in the era of enterprise 2.0: the case of OrganiK. Scalable Computing: Practice and Experience, 9(4):315– 327, December 2008. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol09/no4/SCPE_9_4_08.pdf;
                                         http://www.scpe.org/vols/
                                        vol09/no4/SCPE_9_4_08.zip.

Benoit:2007:ISI

[BL07] Anne Benoît and Frédéric Loulergue. Introduction to the special issue: Practical aspects of high-level parallel programming. Scalable Computing: Practice and Experience, 8(4):v, December 2007. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/vols/vol08/
                                        no4/vol08no4introduction1.
                                        html.

Bruneo:2013:AIA

[BLHK13] Dario Bruneo, Francesco Longo, David Hadas, and Elliot K. Kolodner. Analytical investigation of availability in a vision cloud storage cluster. Scalable Computing: Practice and Experience, 14(4):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/933.

Becker:2009:RBS

[BLRW09] Daniel Becker, John C. Linford, Rolf Rabenseifner, and Felix Wolf. Replay-based synchronization of timestamps in event traces of massively parallel applications. Scalable Computing: Practice and Experience, 10(1):49–60, March 2009. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol10/no1/SCPE_10_1_
05.pdf;   http://www.scpe.
org/vols/vol10/no1/SCPE_
10_1_05.zip.

Begum:2011:DDH

[BM11] Yusuf Mohamadi Begum and M. A. Maluk Mohamed. DEEPG: Dual heap overlay resource discovery protocol for mobile Grid. Scalable Computing: Practice and Experience, 12(2):239–255, June 2011. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/index.php/scpe/
article/view/698;   http:
//www.scpe.org/index.php/
scpe/article/view/717.

Brandic:2010:VFF

[BMD10] Ivona Brandic, Dejan Music, and Schahram Dustdar. VieS-LAF framework: Facilitating negotiations in clouds by applying service mediation and negotiation bootstrapping. Scalable Computing: Practice and Experience, 11(2):189–204, June 2010. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol11/no2/
SCPE_11_2_09.pdf;   http:

                                        //www.scpe.org/vols/vol11/
                                        no2/SCPE_11_2_09.zip.

Braubach:2012:DDS

[BP12] Lars Braubach and Alexander Pokahr. Developing distributed systems with active components and Jadex. Scalable Computing: Practice and Experience, 13(2):100–119, June 2012. CODEN ???? ISSN 1895-

                                        1767. URL http://jadex.
                                        sourcefourge.net/;   http:
                                        //www.scpe.org/index.php/
                                        scpe/article/view/773.

Bhardwaj:2016:PPI

[BP16] Chandrika Bhardwaj and Sanjiva Prasad. Pravah: Parameterised information flow control in e-health. Scalable Computing: Practice and Experience, 17(3):171–188, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1179.

Baldini:2012:CMF

[BPB^{+}12] Gianmarco Baldini, Ranga Rao Vencatesha Prasad, Abdur Rahim Biswas, Klaus Moessner, Matti Etelapera, Juha-Pekka Soininen, Septimiu-Cosmin Nechifor, Vera Stavroulaki, and Panagiotis Vlacheas. A cognitive management framework to support exploitation of the future Internet of Things. Scalable Computing: Practice and Experience, 13(2):139–147, June 2012. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/775.

Binder:2006:SRJ

[BR06] Walter Binder and Volker Roth. Security risks in Javabased mobile code systems. Scalable Computing: Practice and Experience, 7(4):1– 11, December 2006. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol07/no4/SCPE_7_4_01.pdf;
http://www.scpe.org/vols/
vol07/no4/SCPE_7_4_01.zip.

Bless:2010:IEN

[BR10] Roland Bless and Martin Röhricht. Implementation and evaluation of a NAT-gateway for the general Internet signaling transport protocol. Scalable Computing: Practice and Experience, 11(4):329–343, December 2010. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/665.

Brodtkorb:2008:AAN

[Bro08] André Rigland Brodtkorb. An asynchronous API for numerical linear algebra. Scalable Computing: Practice and Experience, 9(3):153–163, September 2008. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol09/no3/SCPE_9_3_01.pdf;
http://www.scpe.org/vols/
vol09/no3/SCPE_9_3_01.zip.

Bosa:2005:TSF

[BS05] Károly Bósa and Wolfgang Schreiner. Tolerating stop failures in Distributed Maple. Scalable Computing: Practice and Experience, 6(2):59– 70, June 2005. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol06/no2/SCPE_6_2_06.pdf;
 http://www.scpe.org/vols/
vol06/no2/SCPE_6_2_06.zip.

Bala:2019:SAG

[BSBB19] Suman Bala, Gaurav Sharma, Himani Bansal, and Tarunpreet Bhatia. On the security of authenticated group key agreement protocols. Scalable Computing: Practice and Experience, 20(1):93–99, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1440.

Bauer:2013:CAS

[BSST13] Christine Bauer, Christine Strauss, Christian Stummer, and Alexander Trieb. Contextaware services in cooperative value chains: a key playercentred approach. Scalable Computing: Practice and Experience, 14(3):??, ???? 2013. CODEN ???? ISSN 1895-

1767.  URL http://www.
scpe.org/index.php/scpe/
article/view/856.

Borkowski:2012:DDP

[BT12] Janusz Borkowski and Marek Tudruj. Dynamic distributed programs control based on global program states monitoring. Scalable Computing: Practice and Experience, 13(2): 173–186, June 2012. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/778.

Bernabeu:2007:RPI

[BTGV07] Miguel O. Bernabeu, Máriam Taroncher, Víctor M. García, and Ana Vidal. Robust parallel implementation of a Lanczosbased algorithm for an structured electromagnetic eigenvalue problem. Scalable Computing: Practice and Experience, 8(3):263–270, September 2007. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol08/no3/
SCPE_8_3_04.pdf;   http:/
/www.scpe.org/vols/vol08/
no3/SCPE_8_3_04.zip.

Bruda:2009:RBS

[BZ09] Stefan D. Bruda and Yuanqiao Zhang. Relations between several parallel computational models. Scalable Computing: Practice and Experience, 10 (2):163–172, June 2009. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol10/no2/SCPE_10_2_
03.pdf;   http://www.scpe.
org/vols/vol10/no2/SCPE_
10_2_03.zip.

Bachara:2011:SCA

[BZ11] Pawel Bachara and Krzysztof Zielinski. Service component architecture extension for sensor networks. Scalable Computing: Practice and Experience, 12(1):121–135, March 2011. CODEN ???? ISSN 1895-1767. URL http://www.

                                        scpe.org/index.php/scpe/
                                        article/view/692.

Charalambous:2018:EVC

[CA18] Panayiotis Charalambous and George Artopoulos. Enabling virtual collaboration in digital cultural heritage in the SEEM region. Scalable Computing: Practice and Experience, 19(2):161–174, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1348.

Cabri:2006:ACR

[Cab06] Giacomo Cabri. Agent composition via role-based infrastructures. Scalable Computing: Practice and Experience, 7(1): 37–47, March 2006. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol07/no1/SCPE_7_1_04.pdf;
                                         http://www.scpe.org/vols/
                                        vol07/no1/SCPE_7_1_04.zip.

Carino:2007:TTL

[CB07] Ricolindo L. Carino and Ioana Banicescu. A tool for a twolevel dynamic load balancing strategy in scientific applications. Scalable Computing: Practice and Experience, 8(3): 249–261, September 2007. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol08/no3/SCPE_8_3_
03.pdf;   http://www.scpe.
org/vols/vol08/no3/SCPE_8_
3_03.zip.

Cabri:2014:ISI

[CBA^{+}14] Giacomo Cabri, Federico Bergenti, Marco Aiello, Sumitra Reddy, and Ramada Reddy. Introduction to the special issue on Enabling Technologies for Collaborations. Scalable Computing: Practice and Experience, 15(4): ??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/1051.

Chettaoui:2014:NDP

[CC14] Hanene Chettaoui and Faouzi Ben Charrada. A new decentralized periodic replication strategy for dynamic data grids. Scalable Computing: Practice and Experience, 15(1):??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/969.

Calabrese:2015:CCH

[CC15] Barbara Calabrese and Mario Cannataro. Cloud computing in healthcare and biomedicine. Scalable Computing: Practice and Experience, 16(1):??, ???? 2015. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/1057.

Cordasco:2009:EML

[CCD^{+}09] Gennaro Cordasco, Biagio Cosenza, Rosario De Chiara, Ugo Erra, and Vittorio Scarano. Experiences with mesh-like computations using prediction binary trees. Scalable Computing: Practice and Experience, 10(2):173–187, June 2009. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol10/no2/SCPE_10_2_
04.pdf;   http://www.scpe.
org/vols/vol10/no2/SCPE_
10_2_04.zip.

Clematis:2010:MJR

[CCD^{+}10] A. Clematis, A. Corana, D. D’Agostino, A. Galizia, and A. Quarati. Matching jobs with resources: an applicationdriven approach. Scalable Computing: Practice and Experience, 11(2):109–120, June 2010. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol11/no2/SCPE_11_2_
02.pdf;   http://www.scpe.
org/vols/vol11/no2/SCPE_
11_2_02.zip.

Colesa:2011:FOO

[CCI11] Adrian Coleşa, Alexandra Coldea, and Iosif Ignat. Flexible organization in the OrcFS relational file system for efficient file searching. Scalable Computing: Practice and Experience, 12(1):79–92, March 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/689.

Casola:2016:SBS

[CDRV16] Valentina Casola, Alessandra De Benedictis, Massimiliano Rak, and Umberto Villano. SLA-based secure cloud application development. Scalable Computing: Practice and Experience, 17(4):271–284, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1200.

Caron:2005:PED

[CDS05] Eddy Caron, Frédéric Desprez, and Frédéric Suter. Parallel extension of a dynamic performance forecasting tool. Scalable Computing: Practice and Experience, 6(1):57–69, March 2005. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol06/SCPE_
6_1_05.pdf;   http://www.
scpe.org/vols/vol06/SCPE_
6_1_05.zip.

Copie:2012:DSS

[CFMN12] Adrian Copie, Teodor Florin Fortis, Victor Ion Munteanu, and Viorel Negru. Datastores supporting services lifecycle in the framework of Cloud governance. Scalable Computing: Practice and Experience, 13(3):251–267, September 2012. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/796.

Church:2015:SAF

[CG15] Philip Church and Andrzej Goscinski. Selected approaches and frameworks to carry out genomic data analysis on the cloud. Scalable Computing: Practice and Experience, 16(1): ??, ???? 2015. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1058.

Carretero:2015:ISI

[CGBG15] Jesus Carretero, Javier Garcia-Blas, and Sandra Gesing. Introduction to the special issue on Cloud for Health. Scalable Computing: Practice and Experience, 16(1):??, ???? 2015. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1056.

Cabri:2015:ISI

[CH15] Giacomo Cabri and Emma Hart. Introduction to the special issue on collective adaptive systems. Scalable Computing: Practice and Experience, 16(3): iii, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1097.

Cheng:2010:BRBb

[Che10a]  Jie Cheng.  Book review:

CUDA by Example: An Introduction to General-Purpose GPU Programming, by Jason Sanders and Edward Kandrot, ISBN-13 978-0-13-138768-3. Scalable Computing: Practice and Experience, 11(4):401, December 2010. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/663. See

[SK10].

Cheng:2010:BRBa

[Che10b] Jie Cheng. Book review: Programming Massively Parallel Processors. A Hands-on Approach, by David Kirk and Wen-mei Hwu, ISBN 978-0-12-381472-2. Scalable Computing: Practice and Experience, 11(3): 327, September 2010. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/654.

See [KmWH10].

Choudhury:2015:ACG

[CHTE15] Olivia Choudhury, Nicholas Hazekamp, Douglas Thain, and Scott Emrich. Accelerating comparative genomics work flows in a distributed environment with optimized data partitioning and work flow fusion. Scalable Computing: Practice and Experience, 16(1): ??, ???? 2015. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/1060.

Ciobanu:2013:STP

[CI13] Alex Ciobanu and Florentin Ipate. P system testing with parallel simulators — a survey. Scalable Computing: Practice and Experience, 14(3):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/858.

Carlsson:2007:SCS

[CJ07] Bengt Carlsson and K. Ingemar Jönsson. The success of cooperative strategies in the Iterated Prisoner’s Dilemma and the Chicken Game. Scalable Computing: Practice and Experience, 8(1): 87–100, March 2007. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol08/no1/SCPE_8_1_08.pdf;
                                         http://www.scpe.org/vols/
                                        vol08/no1/SCPE_8_1_08.zip.

Carchiolo:2008:DTS

[CLM^{+}08] V. Carchiolo, A. Longheu, M. Malgeri, G. Mangioni, and V. Nicosia. The dilemma of trust: a social network based approach. Scalable Computing: Practice and Experience, 9(1): 29–37, March 2008. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol09/no1/SCPE_9_1_04.pdf;
                                         http://www.scpe.org/vols/
                                        vol09/no1/SCPE_9_1_04.zip.

Cash:2009:CHM

[CM09] Jeff R. Cash and Francesca Mazzia. Conditioning and hybrid mesh selection algorithms for two-point boundary value problems. Scalable Computing: Practice and Experience, 10(4): 347–361, December 2009. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol10/no4/SCPE_10_4_
02.pdf;   http://www.scpe.
org/vols/vol10/no4/SCPE_
10_4_02.zip.

Chai:2009:VLS

[CMUM09] Erianto Chai, Katsuyoshi Matsumoto, Minoru Uehara, and Hideki Mori. Virtual large-scale disk base on PC Grid. Scalable Computing: Practice and Experience, 10(1):87–98, March 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no1/
SCPE_10_1_08.pdf;   http:
//www.scpe.org/vols/vol10/
no1/SCPE_10_1_08.zip.

Chomatek:2011:VRP

[CPM11] Lukasz Chomatek and Aneta Poniszewska-Maranda. Vehicle routing problems with the use of multi-agent system. Scalable Computing: Practice and Experience, 12(3):283–291, September 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/724.

Cirani:2013:SIP

[CPV13] Simone Cirani, Marco Picone, and Luca Veltri. A session initiation protocol for the Internet of Things. Scalable Computing: Practice and Experience, 14(4): ??, ???? 2013. CODEN ???? ISSN 1895-1767.

Chen:2007:TQG

[CS07] Gilbert G. Chen and Boleslaw K.

Szymanski. Time quantum GVT: A scalable computation of the Global Virtual Time in parallel discrete event simulations. Scalable Computing: Practice and Experience, 8(4): 423–435, December 2007. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol08/no4/SCPE_8_4_
08.pdf;   http://www.scpe.
org/vols/vol08/no4/SCPE_8_
4_08.zip.

Chopra:2010:ISP

[CS10] Inderpreet Chopra and Maninder Singh. Implementing selfprotection in distributed Grid environment. Scalable Computing: Practice and Experience, 11(4):373–383, December 2010. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/668.

Cesario:2010:UGE

[CT10] Eugenio Cesario and Domenico Talia. Using Grids for exploiting the abundance of data in science. Scalable Computing: Practice and Experience, 11(3):251–262, September 2010. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/658.

Czarnul:2011:PCI

[Cza11] Pawel Czarnul. Parallelization of compute intensive applications into workflows based on services in BeesyCluster. Scalable Computing: Practice and Experience, 12(2):227–238, June 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/716.

Czarnul:2012:ICB

[Cza12]  Pawel Czarnul.  Integra-

tion of cloud-based services into distributed workflow systems: challenges and solutions. Scalable Computing: Practice and Experience, 13(4): ??, ???? 2012. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/814.

dAuriol:2006:HAF

[dB06] Brian J. d’Auriol and Maria Beltran. A historical analysis of fiber based optical bus parallel computing models. Scalable Computing: Practice and Experience, 7(1):77– 93, March 2006. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol07/no1/SCPE_7_1_07.pdf;
http://www.scpe.org/vols/
vol07/no1/SCPE_7_1_07.zip.

Dash:2019:EFC

[DBBR19] Sujata Dash, Sitanath Biswas, Debajit Banerjee, and Atta UR Rahman. Edge and fog computing in healthcare — a review. Scalable Computing: Practice and Experience, 20(2):191–206, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1504.

Derouiche:2019:DFE

[DBGG19] Rihab Derouiche, Zaki Brahmi, Mohammed Mohsen Gammoundi, and Sebastian Garcia Galan. E-DPSIW-FCA: Energy aware FCA-based data placement strategy for intensive workflow. Scalable Computing: Practice and Experience, 20(3):541–562, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1556.

Doci:2009:RAS

[DBX09] Arta Doci, Leonard Barolli, and Fatos Xhafa. Recent advances on the simulation models for ad hoc networks: Real traffic and mobility models. Scalable Computing: Practice and Experience, 10(1):1–11, March 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no1/
SCPE_10_1_01.pdf;   http:
//www.scpe.org/vols/vol10/
no1/SCPE_10_1_01.zip.

Dazzi:2017:TAD

[DCLR17] Patrizio Dazzi, Emanuele Carlini, Alessandro Lulli, and Laura Ricci. TELOS: An approach for distributed graph processing based on overlay composition. Scalable Computing: Practice and Experience, 18(1):35–49, ???? 2017. CODEN ???? ISSN 1895-

1767.  URL https://www.
scpe.org/index.php/scpe/
article/view/1232.

Daoudi:2018:RIM

[DDKK18] El Mostafa Daoudi, Abdelmajid Dargham, Aicha Kerfali, and Mohammed Khatiri. Reducing the interprocessors migrations of the EKG algorithm. Scalable Computing: Practice and Experience, 19(3):293–300, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1397.

Davidovic:2015:EEP

[DDL^{+}15] Davor Davidovic, Matjaz Depolli, Tomislav Lipic, Karolj Skala, and Roman Trobec. Energy efficiency of parallel multicore programs. Scalable Computing: Practice and Experience, 16(4):437–448, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1132.

DAmbra:2010:ISI

[DdSGP10] Pasqua D’Ambra, Daniela di Serafino, Mario Rosario Guarracino, and Francesca Perla. Introduction to the special issue: Parallel, distributed and network-based computing: an application perspective. Scalable Computing: Practice and Experience, 11(3):i, September 2010. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/655.

De:2012:ITP

[DEBM12] Suparna De, Tarek Elsaleh, Payam Barnaghi, and Stefan Meissner. An Internet of Things platform for real-world and digital objects. Scalable Computing: Practice and Experience, 13(1):45–57, March 2012. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/766.

Dragan:2016:EHS

[DFN16] Ioan Dragan, Teodor Florin Fortis, and Marian Neagul. Exposing HPC services in the cloud: the CloudLightning approach. Scalable Computing: Practice and Experience, 17(4):323–330, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1204.

Dietze:2017:IGF

[DHR17] Robert Dietze, Michael Hofmann, and Gudula Ruenger. Integrating generic FEM simulations into complex simulation applications. Scalable Computing: Practice and Experience, 18(2):133–144, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1285.

Davitashvili:2018:EDA

[DKK^{+}18] Teimuraz Davitashvili, Nato Kutaladze, Ramaz Kvatadze, Ramaz Alexander Kvatadze, and George Mikuchadze. Effect of dust aerosols in forming the regional climate of Georgia. Scalable Computing: Practice and Experience, 19(2):87– 96, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1398.

Danese:2010:HPC

[DLB^{+}10] Gianni Danese, Francesco Leporati, Marco Bera, Mauro Giachero, Nelson Nazzicari, and Alvaro Spelgatti. High performance computing through SoC coprocessors. Scalable Computing: Practice and Experience, 11(3):277–288, September 2010. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/660.

Danelutto:2010:MSC

[DLT^{+}10] Marco Danelutto, D. Laforenza, N. Tonellotto, M. Vanneschi, and C. Zoccolo. Modeling stream communications in component-based applications. Scalable Computing: Practice and Experience, 11(3):263–275, September 2010. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/659.

Dobre:2011:ADM

[DM11] Ciprian Dobre and Andreea Marin. Automatic data migration e-system for public administration e-services.

Scalable Computing: Practice and Experience, 12(1):51–62, March 2011. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/687.

Distefano:2013:MSC

[DMP13] Salvatore Distefano, Giovanni Merlino, and Antonio Puliafito. Mobiles for sensing clouds: the SAaaS4Mobile experience. Scalable Computing: Practice and Experience, 14(4): ??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/929.

Righi:2016:UCG

[DMR^{+}16] Rodrigo Da Rosa Righi, Vladimir Magalhaes Guerreiro, Gustavo Rostirolla, Vinicius Facco Rodrigues, Cristiano Andre Da Costa, and Leonardo Dagnino Chiwiacowsky. Using computational geometry to improve process rescheduling on round-based parallel applications. Scalable Computing: Practice and Experience, 17(1): 13–32, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1147.

delMoral:2007:EPF

[dMRGJ07] Beatriz Aparicio del Moral, Julio Rodríguez-Gómez, and Antonio C. López Jiménez. Editorial: The present and the future of reconfigurable devices for space applications. Scalable Computing: Practice and Experience, 8(4):i–iv, December 2007. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol08/no4/
vol08no4editorial.html.

Dhanapal:2019:SHD

[DN19] A. Dhanapal and P. Nithyanandam. The slow HTTP distributed denial of service attack detection in cloud. Scalable Computing: Practice and Experience, 20(2):285–298, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1501.

DaCosta:2005:IRW

[DO05] Georges Da Costa and Richard Olivier. Impact of realistic workload in peer-to-peer systems a case study: Freenet. Scalable Computing: Practice and Experience, 6(1):45–56, March 2005. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/vols/vol06/
SCPE_6_1_04.pdf;   http:/
/www.scpe.org/vols/vol06/
SCPE_6_1_04.zip.

Doci:2009:BRA

[Doc09] Arta Doci. Book review: Ad Hoc Networks: Technologies and Protocols. Scalable Computing: Practice and Experience, 10(1):111–112, March 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no1/
SCPE_10_1_10.pdf;   http:

                                        //www.scpe.org/vols/vol10/
                                        no1/SCPE_10_1_10.zip.

Dobre:2009:SFE

[DPC09] Ciprian Dobre, Florin Pop, and Valentin Cristea. Simulation framework for the evaluation of dependable distributed systems. Scalable Computing: Practice and Experience, 10 (1):13–23, March 2009. CO-DEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/
                                        vols/vol10/no1/SCPE_10_1_
                                        02.pdf;   http://www.scpe.
                                        org/vols/vol10/no1/SCPE_
                                        10_1_02.zip.

Dobre:2011:VBA

[DPCA11] Ciprian Dobre, Florin Pop, Valentin Cristea, and Ovidiu-Marian Achim. A virtualizationbased approach to dependable service computing. Scalable Computing: Practice and Experience, 12(3):337–350, September 2011. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/728.

Dam:2015:ISI

[DPG^{+}15] Hoa Khanh Dam, Jeremy Pitt, Guido Governatori, Takayuki Ito, and Yang Xu. Introduction to the special issue on principles and practices in multi-agent systems. Scalable Computing: Practice and Experience, 16(4): iii, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1126.

Modica:2013:OME

[DPT13] Giuseppe Di Modica, Francesco Pantano, and Orazio Tomarchio. An OSGi middleware to enable the Sensor as a Service paradigm. Scalable Computing: Practice and Experience, 14(4):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/927.

Dimitrov:2018:SDD

[DS18] Vladimir Dimitrov and Stiliyan Stoyanov. Solutions for data discovery service in a virtual research environment. Scalable Computing: Practice and Experience, 19(2):181–187, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1350.

Doci:2009:IDM

[DSX09] Arta Doci, William Springer, and Fatos Xhafa. Impact of the dynamic membership in the connectivity graph of the wireless ad hoc networks. Scalable Computing: Practice and Experience, 10(1):25–34, March 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no1/
SCPE_10_1_03.pdf;   http:
//www.scpe.org/vols/vol10/
no1/SCPE_10_1_03.zip.

Estrada:2015:PCT

[EDS^{+}15] Zachary John Estrada, Fei Deng, Zachary Stephens, Cuong Pham, Zbigniew Kalbarczyk, and Ravishankar Iyer. Performance comparison and tuning of virtual machines for sequence alignment software. Scalable Computing: Practice and Experience, 16(1):??, ???? 2015. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/1061.

Ghor:2018:EER

[EHHC18] Hussein El Ghor, Julia Hage, Nizar Hamadeh, and Rafic Hage Chehade. Energy-efficient realtime scheduling algorithm for fault-tolerant autonomous systems. Scalable Computing: Practice and Experience, 19(4): 387–400, ???? 2018. CO-DEN ???? ISSN 1895-

1767.  URL https://www.
scpe.org/index.php/scpe/
article/view/1438.

Elakehal:2012:MMS

[EP12] Emad Eldeen Elakehal and Julian Padget. MSMAS: Modelling self-managing multi agent systems. Scalable Computing: Practice and Experience, 13(2): 121–137, June 2012. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/774.

Elfarargy:2018:VVM

[ER18] Mohammed Elfarargy and Amr Rizq. VirMuF: The virtual museum framework. Scalable Computing: Practice and Experience, 19(2):175–180, ????

2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1349.

Elwaer:2011:ODD

[ETR11] Abdelhamid Elwaer, Ian Taylor, and Omer Rana. Optimizing data distribution in volunteer computing systems using resources of participants. Scalable Computing: Practice and Experience, 12(2):193–208, June 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/714.

Elnitski:2007:ENF

[EW07] Laura Elnitski and Lonnie R.

Welch. Editorial: New frontiers in computational genomics: Identification of novel functional elements and genomescale pattern analysis. Scalable Computing: Practice and Experience, 8(3):i–vii, September 2007. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol08/no3/
vol08no3editorial.html.

Frasheri:2018:APW

[FA18] Neki Frasheri and Emanouil Atanassov. An analysis for parallel wind simulation speedup using OpenFOAM. Scalable Computing: Practice and Experience, 19(2):97–105, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1342.

Forer:2016:CEF

[FAW^{+}16] Lukas Forer, Enis Afgan, Hansi Weissensteiner, Davor Davidovic, Guenther Specht, Florian Kronenberg, and Sebastian Schoenherr. Cloudflow — enabling faster biomedical pipelines with MapReduce and Spark. Scalable Computing: Practice and Experience, 17(2):103–114, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1159.

Frasheri:2012:AGA

[FB12] Neki Frasheri and Salvatore Bushati. An algorithm for gravity anomaly inversion in HPC. Scalable Computing: Practice and Experience, 13(2): 149–157, June 2012. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/index.
                                        php/scpe/article/view/776.

Fazio:2013:ISI

[FB13] Maria Fazio and Nik Bessis. Introduction to the special issue. Scalable Computing: Practice and Experience, 14(4):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/926.

Furdik:2011:SSI

[FBLF11] Karol Furdík, Peter Bednăr, Gabriel Lukăč, and Christoph Fritsch. Support of semantic interoperability in a servicebased business collaboration platform. Scalable Computing: Practice and Experience, 12(3):293–305, September 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/725.

Fohry:2015:FTS

[FBP15] Claudia Fohry, Marco Bungart, and Jonas Posner. Fault tolerance schemes for global load balancing in X10. Scalable Computing: Practice and Experience, 16(2):169–186, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1088.

Fazio:2013:MOM

[FCPV13] Maria Fazio, Antonio Celesti, Antonio Puliafito, and Massimo Villari. A message oriented middleware for cloud computing to improve efficiency in risk management systems. Scalable Computing: Practice and Experience, 14(4):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/928.

Flejter:2008:ISI

[FKK08a] Dominik Flejter, Tomasz Kaczmarek, and Marek Kowalkiewicz. Introduction to the special issue: The Web on the move. Scalable Computing: Practice and Experience, 9(4):iii, December 2008. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/vols/vol09/
                                        no4/vol09no4introduction.
                                        html.

Flejter:2008:WWW

[FKK08b] Dominik Flejter, Tomasz Kaczmarek, and Marek Kowalkiewicz. World Wide Web on the move. Scalable Computing: Practice and Experience, 9(4):219– 241, December 2008. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol09/no4/SCPE_9_4_01.pdf;
                                         http://www.scpe.org/vols/
                                        vol09/no4/SCPE_9_4_01.zip.

Formisano:2013:DAS

[FKSP^{+}13] Ciro Formisano, Elliot K. Kolodner, Alexandra Shulman-Peleg, Ermanno Travaglino, Gil Vernik, and Massimo Villari. Delegation across storage clouds: on-boarding federation as a case study. Scalable Computing: Practice and Experience, 14(4):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/934.

Frain:2007:HAH

[FMB07] Ivan Frain, Abdelaziz M’zoughi, and Jean-Paul Bahsoun. How to achieve high throughput with dynamic tree-structured coterie. Scalable Computing: Practice and Experience, 8(3): 241–248, September 2007. CO-DEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/
                                        vols/vol08/no3/SCPE_8_3_

02.pdf;   http://www.scpe.
org/vols/vol08/no3/SCPE_8_
3_02.zip.

Frisch:2014:AMG

[FMR14] Jerome Frisch, Ralf-Peter Mundani, and Ernst Rank. Adaptive multi-grid methods for parallel CFD applications. Scalable Computing: Practice and Experience, 15(1):??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/964.

Farahani:2016:SHA

[FNC16] Ali Farahani, Eslam Nazemi, and Giacomo Cabri. A selfhealing architecture based on RAINBOW for industrial usage. Scalable Computing: Practice and Experience, 17(4):351– 368, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1206.

Farahani:2017:EAC

[FNCC17] Ali Farahani, Eslam Nazemi, Giacomo Cabri, and Nicola Capodieci. Enabling autonomic computing support for the JADE agent platform. Scalable Computing: Practice and Experience, 18(1):91–103, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1235.

Fortis:2014:ISIa

[For14a]  Teodor Florin Fortis.  In-

troduction to the special issue. Scalable Computing: Practice and Experience, 15(1):??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/961.

Fortis:2014:ISIb

[For14b]  Teodor Florin Fortis.  In-

troduction to the special issue. Scalable Computing: Practice and Experience, 15(2):??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/978.

Frincu:2010:ONI

[FP10] Marc Eduard Frîncu and Dana Petcu. OSyRIS: a nature inspired workflow engine for service oriented environments. Scalable Computing: Practice and Experience, 11(1): 81–97, March 2010. CO-DEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/
                                        vols/vol11/no1/SCPE_11_1_
                                        08.pdf;   http://www.scpe.
                                        org/vols/vol11/no1/SCPE_
                                        11_1_08.zip.

Frantz:2015:MDN

[FPSN15] Christopher Konstantin Frantz, Martin K. Purvis, Bastin Tony Roy Savarimuthu, and Mariusz Nowostawski. Modelling dynamic normative understanding in agent societies.

Scalable Computing: Practice and Experience, 16(4):355–380, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1128.

Fredj:2007:RID

[FR07] Ouissem Ben Fredj and Éric Renault. RWAPI over Infini-Band: Design and performance. Scalable Computing: Practice and Experience, 8(3):271–280, September 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no3/SCPE_8_3_05.pdf;
 http://www.scpe.org/vols/
vol08/no3/SCPE_8_3_05.zip.

Fogarolli:2008:DSM

[FR08] Angela Fogarolli and Marco Ronchetti. Discovering semantics in multimedia content using Wikipedia. Scalable Computing: Practice and Experience, 9(4):259–269, December 2008. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol09/no4/SCPE_9_4_03.pdf;
 http://www.scpe.org/vols/
vol09/no4/SCPE_9_4_03.zip.

Flich:2008:PNV

[FRD^{+}08] J. Flich, S. Rodrigo, J. Duato, T. Sødring, Å. G. Solheim, T. Skeie, and O. Lysne. On the potential of NoC virtualization for multicore chips. Scalable Computing: Practice and Experience, 9(3):165–177, September 2008. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol09/no3/SCPE_9_3_02.pdf;
                                         http://www.scpe.org/vols/
                                        vol09/no3/SCPE_9_3_02.zip.

Frincu:2013:ISI

[Fri13]  Marc Eduard Frincu.  In-

troduction to the special issue. Scalable Computing: Practice and Experience, 14(1):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/822.

Falcou:2005:VOO

[FS05] Joel Falcou and Jocelyn Serot. E.V.E., an object oriented SIMD library. Scalable Computing: Practice and Experience, 6 (4):31–42, December 2005. CO-DEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/
                                        vols/vol06/no4/SCPE_6_4_
                                        03.pdf;   http://www.scpe.
                                        org/vols/vol06/no4/SCPE_6_
                                        4_03.zip.

Frias:2008:SPL

[FSS08] Leonor Frias, Johannes Singler, and Peter Sanders. Singlepass list partitioning. Scalable Computing: Practice and Experience, 9(3):179–184, September 2008. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol09/no3/SCPE_9_3_03.pdf;
                                         http://www.scpe.org/vols/
                                        vol09/no3/SCPE_9_3_03.zip.

Fiolet:2005:DDM

[FT05] Valerie Fiolet and Bernard Toursel. Distributed data mining. Scalable Computing: Practice and Experience, 6(1):99– 109, March 2005. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/vols/vol06/
SCPE_6_1_08.pdf;   http:/
/www.scpe.org/vols/vol06/
SCPE_6_1_08.zip.

Ferscha:2005:CSA

[FVB05] Alois Ferscha, Simon Vogl, and Wolfgang Beer. Context sensing, aggregation, representation and exploitation in wireless networks. Scalable Computing: Practice and Experience, 6 (2):71–81, June 2005. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol06/no2/SCPE_6_2_07.pdf;
http://www.scpe.org/vols/
vol06/no2/SCPE_6_2_07.zip.

Filali:2015:CFU

[FY15] Fatima Zohra Filali and Belabbas Yagoubi. Classifying and filtering users by similarity measures for trust management in cloud environment. Scalable Computing: Practice and Experience, 16(3):289–302, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1102.

Gawanmeh:2015:CFM

[GA15] Amjad Gawanmeh and Ahmad Alomari. Challenges in formal methods for testing and verification of cloud computing systems. Scalable Computing: Practice and Experience, 16(3):321–332, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1104.

Gawanmeh:2016:FVM

[GAM16] Amjad Gawanmeh, Anas Alazzam, and Bobby Mathew. Formal verification of a microfluidic device for blood cell separation. Scalable Computing: Practice and Experience, 17(3):227–236, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1182.

Gava:2005:EMB

[Gav05] Frédéric Gava. External memory in bulk-synchronous parallel ML. Scalable Computing: Practice and Experience, 6(4): 43–70, December 2005. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol06/no4/SCPE_6_4_
04.pdf;   http://www.scpe.
org/vols/vol06/no4/SCPE_6_
4_04.zip.

Gonnord:2009:QCB

[GB09] Laure Gonnord and Jean-Philippe Babau. Qinna: a component-based framework for runtime safe resource adaptation of embedded systems. Scalable Computing: Practice and Experience, 10(3):253– 264, September 2009. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol10/no3/SCPE_10_3_
03.pdf;   http://www.scpe.
org/vols/vol10/no3/SCPE_
10_3_03.zip.

Grabis:2012:SPE

[GB12] Janis Grabis and Martins Bonders. A simulation platform for evaluation and optimization of composite applications. Scalable Computing: Practice and Experience, 13(4):??, ???? 2012. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/813.

Gandotra:2017:FGM

[GBS17] Ekta Gandotra, Divya Bansal, and Sanjeev Sofat. A framework for generating malware threat intelligence. Scalable Computing: Practice and Experience, 18(3):195–206, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1300.

Golubev:2018:DDP

[GBS18] Alexandr Golubev, Peter Bogatencov, and Grigore Secrieru. DICOM data processing optimization in medical information systems. Scalable Computing: Practice and Experience, 19(2):189–201, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1399.

Gajger:2018:MSG

[GC18] Tomasz Gajger and Pawel Czarnul. Modelling and simulation of GPU processing in the MERPSYS environment. Scalable Computing: Practice and Experience, 19(4):401–422, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1439.

Gadzhev:2018:CAV

[GGG^{+}18] Georgi Gadzhev, Ivelina Georgieva, Kostadin Ganev, Vladimir Ivanov, Nikolay Miloshev, Hristo Chervenkov, and Dimiter Syrakov. Climate applications in a virtual research environment platform. Scalable Computing: Practice and Experience, 19(2):107–118, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1347.

Garcia-Gomez:2012:CCM

[GGJGT^{+}12] Sergio García-Gómez, Miguel Jiménez-Gañán, Yehia Taher, Christof Momm, Frederic Junker, József Biro, Andreas Menychtas, Vasilios Andrikopoulos, and Steve Strauch. Challenges for the comprehensive management of Cloud services in a PaaS framework. Scalable Computing: Practice and Experience, 13(3):201–213, September 2012. CODEN ????

ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/793.

Gawinecki:2005:PAC

[GGKP05] Maciej Gawinecki, Minor Gordon, Pawel Kaczmarek, and Marcin Paprzycki. The problem of agent-client communication on the Internet. Scalable Computing: Practice and Experience, 6(1):111–123, March 2005. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol06/SCPE_
6_1_09.pdf;   http://www.
scpe.org/vols/vol06/SCPE_
6_1_09.zip.

Gursky:2008:KPW

[GHJ^{+}08] P. Gurský, T. Horváth, J. Jirásek, R. Novotný, J. Pribolová, V. Vaneková, and P. Vojtás. Knowledge processing for Web search—an integrated model and experiments. Scalable Computing: Practice and Experience, 9(1):51–59, March 2008. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol09/no1/SCPE_9_1_
06.pdf;   http://www.scpe.
org/vols/vol09/no1/SCPE_9_
1_06.zip.

Gaudin:2012:FCT

[GHV^{+}12] Benoit Gaudin, Mike H.

Hinchey, Emil Vassev, Paddy Nixon, Joao Coelho Garcia, and Walid Maalej. FastFix: a control theoretic view of self-healing for automatic corrective software maintenance.

Scalable Computing: Practice and Experience, 13(1):5–20, March 2012. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/763.

Ganzha:2006:SAS

[GKP06] Maria Ganzha, Wojciech Kuranowski, and Marcin Paprzycki. Software agents in support of student mobility. Scalable Computing: Practice and Experience, 7(1):49– 58, March 2006. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol07/no1/SCPE_7_1_05.pdf;
                                         http://www.scpe.org/vols/
                                        vol07/no1/SCPE_7_1_05.zip.

Gomez:2006:EFT

[GLD06] M. E. Gómez, P. López, and J. Duato. An efficient faulttolerant routing strategy for tori and meshes. Scalable Computing: Practice and Experience, 7(3):51–59, September 2006. CODEN ???? ISSN 1895-1767. URL http://www.

                                        scpe.org/vols/vol07/no3/
                                        SCPE_7_3_05.pdf;   http:/
                                        /www.scpe.org/vols/vol07/
                                        no3/SCPE_7_3_05.zip.

Gross:2010:CES

[GM10] Tom Gross and Nicolai Marquardt. Creating, editing, and sharing complex ubiquitous computing environment configurations with CollaborationBus. Scalable Computing: Practice and Experience, 11(3):289–303, September 2010. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/661.

Gharzouli:2014:TCC

[GMB14] Mohamed Gharzouli, Youcef Messelem, and Mohamed Elhadi Bounas. TTL-Chord: A chord-based approach for Semantic Web services discovery. Scalable Computing: Practice and Experience, 15(1):??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/968.

Grigoras:2005:ISI

[GMP05] Dan Grigoras, John P. Morrison, and Marcin Paprzycki. Introduction to the special issue: Internet-based computing. Scalable Computing: Practice and Experience, 6(1):iii, March 2005. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/vols/vol06/
vol06no1introduction.html.

Gargiulo:2014:SDS

[GMPS14] Claudio Gargiulo, Delfina Malandrino, Donato Pirozzi, and Vittorio Scarano. Simulation data sharing to foster teamwork collaboration. Scalable Computing: Practice and Experience, 15(4):??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/1053.

Gilmore:2005:ERB

[GMW05] Stephen Gilmore, Kenneth MacKenzie, and Nicholas Wolverson. Extending resourcebounded functional programming languages with mutable state and concurrency. Scalable Computing: Practice and Experience, 6(4):17– 30, December 2005. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol06/no4/SCPE_6_4_02.pdf;
                                         http://www.scpe.org/vols/
                                        vol06/no4/SCPE_6_4_02.zip.

Ganzha:2008:ISI

[GP08] Maria Ganzha and Marcin Paprzycki. Introduction to the special issue: Distributed intelligent systems. Scalable Computing: Practice and Experience, 9(1):i, March 2008. CODEN ???? ISSN 1895-

                                        1767.  URL http://www.
                                        scpe.org/vols/vol09/no1/
                                        vol09no1introduction.html.

Ganzha:2009:MIW

[GPD^{+}09] Maria Ganzha, Marcin Paprzycki, Michal Drozdowicz, Mehrdad Senobari, Ivan Lirkov, Sofiya Ivanovska, Richard Olejnik, and Pavel Telegin. Mirroring information within an agent-team-based intelligent Grid middleware; an overview and directions for system development. Scalable Computing: Practice and Experience, 10(4): 397–411, December 2009. CO-DEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/

vols/vol10/no4/SCPE_10_4_
06.pdf;   http://www.scpe.
org/vols/vol10/no4/SCPE_
10_4_06.zip.

Ganzha:2014:IS

[GPS14a] Maria Ganzha, Marcin Paprzycki, and Stanislav Sedukhin. Image scrambling on a “mesh-of-tori” architecture. Scalable Computing: Practice and Experience, 15(1):??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/967.

Ganzha:2014:GMM

[GPS14b] Maria Ganzha, Marcin Paprzycki, and Stanislav G. Sedukhin. Generalized matrix multiplication and its object oriented model. Scalable Computing: Practice and Experience, 15(2): ??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/983.

Gamatie:2009:MDE

[GRY^{+}09] Abdoulaye Gamatié, Éric Rutten, Huafeng Yu, Pierre Boulet, and Jean-Luc Dekeyser. Modeldriven engineering and formal validation of high-performance embedded systems. Scalable Computing: Practice and Experience, 10(2):147–161, June 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no2/
SCPE_10_2_02.pdf;   http:

                                        //www.scpe.org/vols/vol10/
                                        no2/SCPE_10_2_02.zip.

Gawanmeh:2017:ISI

[GS17] Amjad Gawanmeh and Kashif Saleem. Introduction to the special issue on communication, computing, and networking in cyber-physical systems. Scalable Computing: Practice and Experience, 18(4):iii, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1328.

Goldschmidt:2005:CBM

[GTB05] Balázs Goldschmidt, Roland Tusch, and László Böszörményi. A CORBA-based middleware for an adaptive streaming server. Scalable Computing: Practice and Experience, 6(2): 83–92, June 2005. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol06/no2/SCPE_6_2_08.pdf;
                                         http://www.scpe.org/vols/
                                        vol06/no2/SCPE_6_2_08.zip.

Galdun:2009:SDL

[GTL09] Jàn Galdun, Jean-Marc Thiriet, and Jàn Ligus. Study of different load dependencies among shared redundant systems. Scalable Computing: Practice and Experience, 10(3): 241–252, September 2009. CO-DEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/
                                        vols/vol10/no3/SCPE_10_3_
                                        02.pdf;   http://www.scpe.

org/vols/vol10/no3/SCPE_
10_3_02.zip.

Gu:2007:ISI

[Gu07] Dazhang Gu. Introduction to the special issue: Bioinformatics. Scalable Computing: Practice and Experience, 8(2): iii, June 2007. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/vols/vol08/
no2/vol08no2introduction.
html.

Guns:2008:UNP

[Gun08] Raf Guns. Unevenness in network properties on the Social Semantic Web. Scalable Computing: Practice and Experience, 9(4):271–279, December 2008. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol09/no4/SCPE_9_4_04.pdf;
http://www.scpe.org/vols/
vol09/no4/SCPE_9_4_04.zip.

Giedrimas:2011:GCC

[GVJ11] Vaidas Giedrimas, Audrius Varoneckas, and Algimantas Juozapavicius. The Grid and Cloud computing facilities in Lithuania. Scalable Computing: Practice and Experience, 12(4):417–421, December 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/750.

Hassan:2019:ISI

[HAA19] Rosilah Hassan, Khairol Amali Bin Ahmad, and Khaleel Ahmad.

Introduction to the special issue on opportunistic network and its security challenges. Scalable Computing: Practice and Experience, 20(1):iii– v, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1519.

Hashmi:2018:ODB

[HDA18] Adeel Shiraz Hashmi, Mohammad Najmud Doja, and Tanvir Ahmad. An optimized densitybased algorithm for anomaly detection in high dimensional datasets. Scalable Computing: Practice and Experience, 19(1): 69–77, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1394.

Hauser:2007:PLD

[HDSY07] T. Hauser, A. Dasu, A. Sudarsanam, and S. Young. Performance of a LU decomposition on a multi-FPGA system compared to a low power commodity microprocessor system. Scalable Computing: Practice and Experience, 8(4):373– 385, December 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no4/SCPE_8_4_04.pdf;
 http://www.scpe.org/vols/
vol08/no4/SCPE_8_4_04.zip.

Hajibaba:2017:SSP

[HGS17] Majid Hajibaba, Saed Gorgin, and Mohsen Sharifi. Sequence similarity parallelization over heterogeneous computer clusters using data parallel programming model. Scalable Computing: Practice and Experience, 18(1):51–66, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1233.

Hege:2005:PRH

[HHK^{+}05] Hans-Christian Hege, Andrei Hutanu, Ralf Kähler, Andre Merzky, Thomas Radke, Edward Seidel, and Brygg Ullmer. Progressive retrieval and hierarchical visualization of large remote data. Scalable Computing: Practice and Experience, 6(3): 57–66, September 2005. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol06/no3/SCPE_6_3_
06.pdf;   http://www.scpe.
org/vols/vol06/no3/SCPE_6_
3_06.zip.

Herrera:2006:BJI

[HHML06] Jose Herrera, Eduardo Huedo, Rubén S. Montero, and Ignacio Martín Llorente. Benchmarking of a joint IRISGrid/ EGEE testbed with a bioinformatics application. Scalable Computing: Practice and Experience, 7(2):25–34, June 2006. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol07/no2/SCPE_7_2_
03.pdf;   http://www.scpe.
org/vols/vol07/no2/SCPE_7_
2_03.zip.

Harnik:2011:SAM

[HKR^{+}11] Danny Harnik, Elliot K. Kolodner, Shahar Ronen, Julian Satran, Alexandra Shulman-Peleg, and Sivan Tal. Secure access mechanism for Cloud storage. Scalable Computing: Practice and Experience, 12(3):317– 336, September 2011. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/index.
                                        php/scpe/article/view/727.

Heinzlreiter:2005:GBV

[HKV05] Paul Heinzlreiter, Dieter Kranzlmüller, and Jens Volkert. Grid-based virtual reality visualization of volume data. Scalable Computing: Practice and Experience, 6(2):35– 44, June 2005. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol06/no2/SCPE_6_2_04.pdf;
                                         http://www.scpe.org/vols/
                                        vol06/no2/SCPE_6_2_04.zip.

Hexmoor:2007:CAA

[HM07] Henry Hexmoor and Brian Mclaughlan. Computationally adjustable autonomy. Scalable Computing: Practice and Experience, 8(1):41–48, March 2007. CODEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/
                                        vols/vol08/no1/SCPE_8_1_
                                        04.pdf;   http://www.scpe.
                                        org/vols/vol08/no1/SCPE_8_
                                        1_04.zip.

Huedo:2005:GFA

[HML05] Eduardo Huedo, Rubén S. Montero, and Ignacio Martín Llorente. The GridWay framework for adaptive scheduling and execution on Grids. Scalable Computing: Practice and Experience, 6(3):1– 8, September 2005. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol06/no3/SCPE_6_3_01.pdf;
http://www.scpe.org/vols/
vol06/no3/SCPE_6_3_01.zip.

Hassan:2017:IPA

[HMS17] Hanan A. Hassan, Aya I.

Maiyza, and Walaa M. Sheta. Impact of process allocation strategies in high performance cloud computing on Azure platform. Scalable Computing: Practice and Experience, 18(2):161–176, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1288.

Hoschek:2005:ERS

[Hos05] Wolfgang Hoschek. Enabling rich service and resource discovery with a database for dynamic distributed content. Scalable Computing: Practice and Experience, 6(1):1–16, March 2005. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol06/SCPE_
6_1_01.pdf;   http://www.
scpe.org/vols/vol06/SCPE_
6_1_01.zip.

Harakaly:2005:PCP

[HPBG05] Robert Harakaly, Pascale Primet, Franck Bonnassieux, and Benjamin Gaidioz. Probes coordination protocol for network performance measurement in GRID computing environment. Scalable Computing: Practice and Experience, 6(1):71–80, March 2005. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/vols/vol06/
SCPE_6_1_06.pdf;   http:/
/www.scpe.org/vols/vol06/
SCPE_6_1_06.zip.

Hexmoor:2006:ISI

[HPS06] Henry Hexmoor, Marcin Paprzycki, and Niranjan Suri. Introduction to the special issue: Software agent mobility. Scalable Computing: Practice and Experience, 7(4):iii, December 2006. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/vols/vol07/
no4/vol07no4introduction.
html.

Hossen:2019:ADT

[HR19] Md. Sharif Hossen and Muhammad Sajjadur Rahim. Analysis of delay-tolerant routing protocols using the impact of mobility models. Scalable Computing: Practice and Experience, 20(1):17–26, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1450.

Hutanu:2010:LSP

[HSB^{+}10] Andrei Hutanu, Erik Schnetter, Werner Benger, Eloisa Bentivegna, Alex Clary, Peter Diener, Jinghua Ge, Robert Kooima, Oleg Korobkin, Kexi Liu, Frank Löffler, Ravi Paruchuri, Jian Tao, Cornelius Toole, Adam Yates, and Gabrielle Allen. Large scale problem solving using automatic code generation and distributed visualization. Scalable Computing: Practice and Experience, 11(2):205–220, June 2010. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol11/no2/
SCPE_11_2_10.pdf;   http:
//www.scpe.org/vols/vol11/
no2/SCPE_11_2_10.zip.

Habala:2010:DDI

[HST^{+}10] Ondrej Habala, Martin Seleng, Viet Tran, Ladislav Hluchý, Martin Kremler, and Martin Gera. Distributed data integration and mining using AD-MIRE technology. Scalable Computing: Practice and Experience, 11(2):131–136, June 2010. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol11/no2/
SCPE_11_2_04.pdf;   http:
//www.scpe.org/vols/vol11/
no2/SCPE_11_2_04.zip.

Hoshino:2005:AFD

[HTC05] Takashi Hoshino, Kenjiro Taura, and Takashi Chikayama. An adaptive file distribution algorithm for wide area network. Scalable Computing: Practice and Experience, 6(3):67– 84, September 2005. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/

                                        vol06/no3/SCPE_6_3_07.pdf;
                                         http://www.scpe.org/vols/
                                        vol06/no3/SCPE_6_3_07.zip.

Hajamydeen:2019:DDU

[HU19] Asif Iqbal Hajamydeen and Nur Izura Udzir. A detailed description on unsupervised heterogeneous anomaly based intrusion detection framework. Scalable Computing: Practice and Experience, 20(1):113–160, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1465.

Han:2008:EIW

[HYM^{+}08] Wei Han, Ying Yi, Mark Muir, Ioannis Nousias, Tughrul Arslan, and Ahmet T. Erdogan. Efficient implementation of WiMAX physical layer on multi-core architectures with dynamically reconfigurable processors. Scalable Computing: Practice and Experience, 9(3): 185–196, September 2008. CO-DEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/
                                        vols/vol09/no3/SCPE_9_3_
                                        04.pdf;   http://www.scpe.
                                        org/vols/vol09/no3/SCPE_9_
                                        3_04.zip.

Ivask:2011:DFS

[IDU11] Eero Ivask, Sergei Devadze, and Raimund Ubar. Distributed fault simulation with collaborative load balancing for VLSI circuits. Scalable Computing: Practice and Experience, 12(1): 153–163, March 2011. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/694.

Iuhasz:2016:ASP

[IPD16] Gabriel Iuhasz, Daniel Pop, and Ioan Dragan. Architecture of a scalable platform for monitoring multiple big data frameworks. Scalable Computing: Practice and Experience, 17(4):313–321, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1203.

Ikken:2018:EHD

[IRTK18] Sonia Ikken, Eric Renault, Abdelkamel Tari, and Tahar Kechadi. Exact and heuristic data workflow placement algorithms for big data computing in cloud datacenters. Scalable Computing: Practice and Experience, 19(3):223–244, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1365.

Ismail:2012:RBP

[Ism12] Leila Ismail. A ring-based parallel oil reservoir simulator. Scalable Computing: Practice and Experience, 13(1):85– 98, March 2012. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/769.

Juhas:2005:TRF

[JAKP05] Zoltan Juhas, Arpad Andics, Krisztian Kuntner, and Szabolcs Pota. Towards a robust and fault-tolerant multicast discovery architecture for global computing Grids. Scalable Computing: Practice and Experience, 6(2):23– 33, June 2005. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol06/no2/SCPE_6_2_03.pdf;
 http://www.scpe.org/vols/
vol06/no2/SCPE_6_2_03.zip.

Jay:2009:PIR

[Jay09] Laurent O. Jay. Preconditioning of implicit Runge–Kutta methods. Scalable Computing: Practice and Experience, 10(4): 363–372, December 2009. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol10/no4/SCPE_10_4_
03.pdf;   http://www.scpe.
org/vols/vol10/no4/SCPE_
10_4_03.zip.

Jansen:2015:EXT

[JBW^{+}15] Christoph Jansen, Maximilian Beier, Michael Witt, Jie Wu, and Dagmar Krefting. Extending XNAT towards a cloud-based quality assessment platform for retinal optical coherence tomographies. Scalable Computing: Practice and Experience, 16(1): ??, ???? 2015. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/1062.

Jacob:2006:YLS

[JCCC06] Joseph C. Jacob, James B.

Collier, Loring G. Craymer, and David W. Curkendall. yourSkyG: Large-scale astronomical image mosaicking on the information power Grid. Scalable Computing: Practice and Experience, 7(1):59– 75, March 2006. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol07/no1/SCPE_7_1_06.pdf;
 http://www.scpe.org/vols/
vol07/no1/SCPE_7_1_06.zip.

Jahani:2017:AMA

[JDK17] Arezoo Jahani, Farnaz Derakhshan, and Leyli Mohammad Khanli. ARank: A multiagent based approach for ranking of cloud computing services. Scalable Computing: Practice and Experience, 18(2):105–116, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1283.

Janies:2007:LSP

[JGP07] Daniel A. Janies, Pablo A.

Goloboff, and Diego Pol. Largescale phylogenetic analysis for the study of zoonosis and assessment of influenza surveillance. Scalable Computing: Practice and Experience, 8(2): 143–146, June 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no2/SCPE_8_2_01.pdf;
 http://www.scpe.org/vols/
vol08/no2/SCPE_8_2_01.zip.

Juhasz:2016:GBS

[JK16] Zoltan Juhasz and Gyorgy Kozmann. A GPU-based soft realtime system for simultaneous EEG processing and visualization. Scalable Computing: Practice and Experience, 17(2): 61–78, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1156.

Jahani:2017:MSF

[JK17] Arezoo Jahani and Leyli Mohammad Khanli. MooreCube: A scalable and flexible architecture for cloud computing data centers on multi-port servers. Scalable Computing: Practice and Experience, 18(4):361–374, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1334.

John:2008:OOC

[JM08] Sunil John and John P. Morrison. Object oriented condensed graphs. Scalable Computing: Practice and Experience, 9(2): 83–94, June 2008. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol09/no2/SCPE_9_2_01.pdf;
                                         http://www.scpe.org/vols/
                                        vol09/no2/SCPE_9_2_01.zip.

Jakob:2010:FMO

[JMQ^{+}10] Wilfried Jakob, Florian Möser, Alexander Quinte, Karl-Uwe Stucky, and Wolfgang Süß. Fast multi-objective rescheduling of workflows to constrained resources using heuristics and memetic evolution. Scalable Computing: Practice and Experience, 11(2):173–188, June 2010. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol11/no2/
SCPE_11_2_08.pdf;   http:
//www.scpe.org/vols/vol11/
no2/SCPE_11_2_08.zip.

Jeanvoine:2007:UON

[JRML07] Emmanuel Jeanvoine, Louis Rilling, Christine Morin, and Daniel Leprince. Using overlay networks to build operating system services for large scale Grids. Scalable Computing: Practice and Experience, 8(3):229–239, September 2007. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol08/no3/SCPE_8_3_
01.pdf;   http://www.scpe.
org/vols/vol08/no3/SCPE_8_
3_01.zip.

Junker:2012:AAP

[JVS12] Frederic Junker, Jürgen Vogel, and Katarina Stanoevska. Approaches to aggregate price models to enable composite services on electronic marketplaces. Scalable Computing: Practice and Experience, 13(3):233–249, September 2012. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/795.

Kushwaha:2019:PSO

[KA19] Arvinda Kushwaha and Mohd Amjad. A particle swarm optimization based load scheduling algorithm in cloud platform for wireless sensor networks. Scalable Computing: Practice and Experience, 20(1): 71–82, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1464.

Karaivanova:2010:PQA

[KAG^{+}10] Aneta Karaivanova, Emanouil Atanassov, Todor Gurov, Sofiya Ivanovska, and Mariya Durchova. Parallel quasirandom applications on heterogeneous Grid. Scalable Computing: Practice and Experience, 11 (1):73–80, March 2010. CO-DEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/
                                        vols/vol11/no1/SCPE_11_1_
                                        07.pdf;   http://www.scpe.
                                        org/vols/vol11/no1/SCPE_
                                        11_1_07.zip.

Koteska:2018:SAC

[KCD^{+}18] Bojana Koteska, Maja Simonoska Crcarevska, Marija Glavas Dodov, Jasmina Tonic Ribarska, and Ljupco Pejov. Semiempirical atom-centered density matrix propagation approach to temperaturedependent vibrational spectroscopy of irinotecan. Scalable Computing: Practice and Experience, 19(2):149–159, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1344.

Kumar:2019:SSH

[KD19] Chanchal Kumar and Mohammad Najmud Doja. A secure structure for hiding information in a cryptosystem based on machine-learning techniques and content-based optimization using portfolio selection data. Scalable Computing: Practice and Experience, 20(1):161–180, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1488.

Kaur:2019:HBT

[KGS19] Avinash Kaur, Pooja Gupta, and Manpreet Singh. Hybrid balanced task clustering algorithm for scientific workflows in cloud computing. Scalable Computing: Practice and Experience, 20(2):237–258, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1515.

Kaur:2019:DPE

[KGSN19] Avinash Kaur, Pooja Gupta, Manpreet Singh, and Anand Nayyar. Data placement in era of cloud computing: a survey, taxonomy and open research issues. Scalable Computing: Practice and Experience, 20(2):377–398, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1530.

Khorasani:2007:BRA

[Kho07] Elham S. Khorasani. Book review: Algorithms Sequential & Parallel: A Unified Approach. Scalable Computing: Practice and Experience, 8(1): 141–142, March 2007. CO-DEN ???? ISSN 1895-

                                        1767.  URL http://www.
                                        scpe.org/vols/vol08/no1/
                                        vol08no1bookreview.html.

Khorasani:2008:BRA

[Kho08] Elham S. Khorasani. Book review: Artificial Intelligence: Structures and Strategies for Complex Problem Solving. Scalable Computing: Practice and Experience, 9(3):219–220, September 2008. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/vols/vol09/
                                        no3/BR_100.html.

Kristensen:2011:HPP

[KHV11] Mads Kristensen, Hans Happe, and Brian Vinter. Hybrid parallel programming for Blue Gene/ P. Scalable Computing: Practice and Experience, 12(2):265– 274, June 2011. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/719.

Kail:2016:PBC

[KKKK16] Eszter Kail, Krisztian Karoczkai, Peter Kacsuk, and Miklos Kozlovszky. Provenance based checkpointing method for dynamic health care smart system. Scalable Computing: Practice and Experience, 17(2): 143–153, ???? 2016. CO-DEN ???? ISSN 1895-

1767.  URL https://www.
scpe.org/index.php/scpe/
article/view/1162.

Kola:2005:RTA

[KKL05] George Kola, Tevfik Kosar, and Miron Livny. Run-time adaptation of Grid data placement jobs. Scalable Computing: Practice and Experience, 6(3): 33–43, September 2005. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol06/no3/SCPE_6_3_
04.pdf;   http://www.scpe.
org/vols/vol06/no3/SCPE_6_
3_04.zip.

Karagiorgos:2008:LBN

[KM08] Gregory Karagiorgos and Nikolaos M. Missirlis. Load balancing for the numerical solution of the Navier–Stokes equations. Scalable Computing: Practice and Experience, 9(1): 61–68, March 2008. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol09/no1/SCPE_9_1_07.pdf;
 http://www.scpe.org/vols/
vol09/no1/SCPE_9_1_07.zip.

Karaivanova:2018:ISI

[KM18] Aneta Karaivanova and Anastas Mishev. Introduction to the special issue on einfrastructures for excellent science: Advances in life sciences, digital cultural heritage and climatology. Scalable Computing: Practice and Experience, 19(2): iii–iv, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1401.

Kaur:2019:SLB

[KM19] Mandeep Kaur and Rajni Mohana. Static load balancing technique for geographically partitioned public cloud. Scalable Computing: Practice and Experience, 20(2):299–316, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1520.

Kakehi:2017:EPT

[KME17] Kazuhiko Kakehi, Kiminori Matsuzaki, and Kento Emoto. Efficient parallel tree reductions on distributed memory environments. Scalable Computing: Practice and Experience, 18(1): 1–15, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1230.

Kouzinopoulos:2015:MSM

[KMM15] Charalampos S. Kouzinopoulos, Panagiotis D. Michailidis, and Konstantinos G. Margaritis. Multiple string matching on a GPU using CUDAs. Scalable Computing: Practice and Experience, 16(2):121–138, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1085.

Koteska:2018:DVS

[KMMP18] Bojana Koteska, Verce Manevska, Anastas Mishev, and Ljupco Pejov. Dynamic versus static approach to theoretical anharmonic vibrational spectroscopy of molecular species elevant to atmospheric chemistry: A case study of formic acid. Scalable Computing: Practice and Experience, 19(2):119–130, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1352.

Kirk:2010:PMP

[KmWH10] David B. Kirk and Wen mei W. Hwu. Programming Massively Parallel Processors: a Hands-on Approach. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2010. ISBN 0-12-381472-3. xviii + 258 pp. LCCN QA76.642 .K57 2010. Chapter 7 (pages 125– 140) discusses GPU floatingpoint considerations.

Kornecki:2008:ASC

[Kor08] Andrew J. Kornecki. Airborne software: Communication and certification. Scalable Computing: Practice and Experience, 9(1):77–82, March 2008. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol09/no1/SCPE_9_1_
09.pdf;   http://www.scpe.

                                        org/vols/vol09/no1/SCPE_9_
                                        1_09.zip.

Kowalik:2006:EQV

[Kow06] Janusz S. Kowalik. Editorial: Quo vadis Grid computing. Scalable Computing: Practice and Experience, 7(3):i, September 2006. CODEN ???? ISSN 1895-1767. URL http://www.

                                        scpe.org/vols/vol07/no3/
                                        vol07no3editorial.html.

Kolar:2012:ABA

[KP12] Jiri Kolar and Tomas Pitner. Agile BPM in the age of Cloud technologies. Scalable Computing: Practice and Experience, 13(4):??, ???? 2012. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/index.
                                        php/scpe/article/view/810.

Kumar:2019:ODS

[KS19] Akshi Kumar and Abhilasha Sharma. Ontology driven social big data analytics for fog enabled sentic-social governance. Scalable Computing: Practice and Experience, 20(2):223–236, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1513.

Kostelnik:2011:SMN

[KSF11] Peter Kostelník, Martin Sarnovský, and Karol Furdík. The semantic middleware for networked embedded systems applied in the Internet of Things and Services domain. Scalable Computing: Practice and Experience, 12(3):307–315, September 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/726.

Kimovski:2016:MOM

[KSSP16] Dragi Kimovski, Nishant Saurabh, Vlado Stankovski, and Radu Prodan. Multi-objective middleware for distributed VMI repositories in federated cloud environment. Scalable Computing: Practice and Experience, 17(4):299–312, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1202.

Kulis:2009:BRB

[Kul09] Anthony Kulis. Book review:

Bio-Inspired Artificial Intelligence: Theories, Methods, and Technologies. Scalable Computing: Practice and Experience, 10(4):443–444, December 2009. CODEN ???? ISSN 1895-

1767. URL http://www.scpe.
org/vols/vol10/no4/BR_100.
html.

Kubert:2011:USL

[KW11] Roland Kübert and Stefan Wesner. Using service level agreements in a high-performance computing environment. Scalable Computing: Practice and Experience, 12(2):164–177, June 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/712.

Kadri:2019:OSA

[KY19] Walid Kadri and Belabbas Yagoubi. Optimized scheduling approach for scientific applications based on clustering in cloud computing environment. Scalable Computing: Practice and Experience, 20(3):527–540, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1548.

Langin:2007:BRL

[Lan07] Chet Langin. Book review: Languages and Machines: An Introduction to the Theory of Computer Science. Scalable Computing: Practice and Experience, 8(4):437, December 2007. CODEN ???? ISSN 1895-

                                        1767. URL http://www.scpe.
                                        org/vols/vol08/no4/BR_100.
                                        html.

Langin:2008:BRI

[Lan08] Chet Langin. Book review: Introduction to Data Mining. Scalable Computing: Practice and Experience, 9(4):353–354, December 2008. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/vols/vol09/
                                        no4/BR_100.html.

Litke:2015:SBO

[LB15] Wieslawa Litke and Marcin Budka. Scaling beyond one rack and sizing of Hadoop platform. Scalable Computing: Practice and Experience, 16(4):423–436, ???? 2015. CODEN ????

ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1131.

List:2007:ABS

[LBP^{+}07] Edward O. List, Darlene E.

Berryman, Amanda J. Palmer, Elahu Gosney, Shigeru Okada, Bruce Kelder, Jens Lichtenberg, Lonnie R. Welch, and John J. Kopchick. Application of bioinformatics and scalable computing to perform proteomic analysis of stomach tissue from diabetic mice. Scalable Computing: Practice and Experience, 8(2):173– 183, June 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no2/SCPE_8_2_04.pdf;
http://www.scpe.org/vols/
vol08/no2/SCPE_8_2_04.zip.

Latiu:2013:NPE

[LCV13] Gentiana Ioana Latiu, Octavian Augustin Cret, and Lucia Vacariu. New performance estimation formula for evolutionary testing of switch-case constructs. Scalable Computing: Practice and Experience, 14(1):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/827.

Li:2006:PPS

[LdDK06] Shujun Li, Elise de Doncker, and Karlis Kaugars. Pion: A problem solving environment for parallel multivariate integration. Scalable Computing:

Practice and Experience, 7(3): 87–94, September 2006. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol07/no3/SCPE_7_3_
08.pdf;   http://www.scpe.
org/vols/vol07/no3/SCPE_7_
3_08.zip.

Lee:2007:BRD

[Lee07] Yung-Chuan Lee. Book review: Distributed Computing: Principles and Applications. Scalable Computing: Practice and Experience, 8(2):227–228, June 2007. CODEN ???? ISSN 1895-

1767. URL http://www.scpe.
org/vols/vol08/no2/BR_100.
html.

Louati:2015:MAA

[LEP15] Amine Louati, Joyce El Haddad, and Suzanne Pinson. A multi-agent approach for trustbased service discovery and selection in social networks. Scalable Computing: Practice and Experience, 16(4):381–402, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1129.

Le:2017:CMM

[LH17] Ngoc T. Le and Doan B. Hoang. Capability maturity model and metrics framework for cyber cloud security. Scalable Computing: Practice and Experience, 18(4):277–290, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1329.

Lopez:2007:SAW

[LHS07] María M. López, Elisa Heymann, and Miquel A. Senar. Sensitivity analysis of workflow scheduling on Grid systems. Scalable Computing: Practice and Experience, 8(3):301–311, September 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no3/SCPE_8_3_08.pdf;
 http://www.scpe.org/vols/
vol08/no3/SCPE_8_3_08.zip.

Lovas:2005:AGD

[LKHH05] Róbert Lovas, Péter Kacsuk, Ákos Horváth, and András Horányi. Application of p-grade development environment in meteorology. Scalable Computing: Practice and Experience, 6 (2):13–22, June 2005. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol06/no2/SCPE_6_2_02.pdf;
 http://www.scpe.org/vols/
vol06/no2/SCPE_6_2_02.zip.

Lieu:2007:RAT

[LLS07] My Chuong Lieu, Siew Kei Lam, and Thambipillai Srikanthan. Rapid area-time estimation technique for porting Cbased applications onto FPGA platforms. Scalable Computing: Practice and Experience, 8(4): 359–371, December 2007. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol08/no4/SCPE_8_4_
03.pdf;   http://www.scpe.
org/vols/vol08/no4/SCPE_8_
4_03.zip.

LeBerre:2009:RTB

[LMPQ09] Tanguy Le Berre, Philippe Mauran, Gérard Padiou, and Philippe Quéinnec. Real time behavior of data in distributed embedded systems. Scalable Computing: Practice and Experience, 10(3):229–239, September 2009. CODEN ???? ISSN 1895-1767. URL http://www.

                                        scpe.org/vols/vol10/no3/
                                        SCPE_10_3_01.pdf;   http:
                                        //www.scpe.org/vols/vol10/
                                        no3/SCPE_10_3_01.zip.

Loulergue:2005:ISI

[Lou05]  Frédéric Loulergue.  Intro-

duction to the special issue: Practical aspects of high-level parallel programming. Scalable Computing: Practice and Experience, 6(4):iii, December 2005. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/vols/vol06/
                                        no4/vol06no4introduction.
                                        html.

Loulergue:2006:ISI

[Lou06]  Frédéric Loulergue.  Intro-

duction to the special issue: Practical aspects of high-level parallel programming. Scalable Computing: Practice and Experience, 7(3):iii, September 2006. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/vols/vol07/
                                        no3/vol07no3introduction.
                                        html.

Loulergue:2017:BSA

[Lou17a] Frederic Loulergue. A BSPlibstyle API for bulk synchronous parallel ML. Scalable Computing: Practice and Experience, 18(3):261–274, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1306.

Loulergue:2017:ISI

[Lou17b]  Frederic Loulergue.  Intro-

duction to the special issue on practical aspects of highlevel parallel programming. Scalable Computing: Practice and Experience, 18(1):iii, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1229.

Lara:2016:ISI

[LPJ16] Pedro Valero Lara, Fernando L.

Pelayo, and Johan Jansson. Introduction to the special issue on high performance computing solutions for complex problems. Scalable Computing: Practice and Experience, 17(1): iii–iv, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1145.

Ludwig:2006:MSD

[LRNP06] Simone A. Ludwig, Omer F.

Rana, William Naylor, and Julian Padget. Mathematical service discovery: Architecture, implementation and performance. Scalable Computing: Practice and Experience, 7(2): 35–52, June 2006. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol07/no2/SCPE_7_2_04.pdf;
                                         http://www.scpe.org/vols/
                                        vol07/no2/SCPE_7_2_04.zip.

Lomonosov:2007:SOC

[LS07] Andrew Lomonosov and Meera Sitharam. Stability, optimality and complexity of network games with pricing and player dropouts. Scalable Computing: Practice and Experience, 8(1): 79–86, March 2007. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol08/no1/SCPE_8_1_07.pdf;
                                         http://www.scpe.org/vols/
                                        vol08/no1/SCPE_8_1_07.zip.

Letia:2011:TBH

[LS11] Ioan Alfred Leţia and Radu Răzvan Slăvescu. Trusted beliefs for helpful behavior when building Web services. Scalable Computing: Practice and Experience, 12(1):105–120, March 2011. CODEN ???? ISSN 1895-1767. URL http://www.

                                        scpe.org/index.php/scpe/
                                        article/view/691.

Langr:2018:AMF

[LS18] Daniel Langr and Ivan Simecek. Analysis of memory footprints of sparse matrices partitioned into uniformly-sized blocks. Scalable Computing: Practice and Experience, 19(3):275–292, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1358.

Langr:2014:LSV

[LSTD14] D. Langr, I. Simecek, P. Tvrdik, and T. Dytrych. Large-scale visualization of sparse matrices. Scalable Computing: Practice and Experience, 15(1):??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/963.

Langr:2016:ASM

[LTS16] Daniel Langr, Pavel Tvrdik, and Ivan Simecek. AQsort: Scalable multi-array inplace sorting with OpenMP. Scalable Computing: Practice and Experience, 17(4):369–391, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1207.

Ludwig:2007:EVW

[Lud07] Thomas Ludwig. Editorial:

Vietnam on its way to high performance computing. Scalable Computing: Practice and Experience, 8(1):i, March 2007. CODEN ???? ISSN 1895-

1767.  URL http://www.
scpe.org/vols/vol08/no1/
vol08no1editorial.html.

Lindback:2010:VSC

[LVMV10] Leif Lindbäck, Vladimir Vlassov, Shahab Mokarizadeh, and Gabriele Violino. VOFS: A secure churn-tolerant Grid file system. Scalable Computing: Practice and Experience, 11 (2):99–108, June 2010. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol11/no2/SCPE_11_2_
01.pdf;   http://www.scpe.
org/vols/vol11/no2/SCPE_
11_2_01.zip.

Mohamed:2017:MCC

[MAJLMJ17] Nader Mohamed, Jameela Al-Jaroodi, Sanja Lazarova-Molnar, and Imad Jawhar. Middleware challenges for cyber-physical systems. Scalable Computing: Practice and Experience, 18(4):331–346, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1332.

Marowka:2006:BRP

[Mar06] Ami Marowka. Book review: Parallel Scientific Computation: A Structured Approach using BSP and MPI. Scalable Computing: Practice and Experience, 7(2):107–108, June 2006. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol07/no2/
vol07no2bookreview.html.

Marowka:2009:PSF

[Mar09] Ami Marowka. Performance study of the first three Intel multicore processors. Scalable Computing: Practice and Experience, 10(4):429–441, December 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no4/
SCPE_10_4_09.pdf;   http:
//www.scpe.org/vols/vol10/
no4/SCPE_10_4_09.zip.

Matsuzaki:2017:EIT

[Mat17] Kiminori Matsuzaki. Efficient implementation of tree skeletons on distributed-memory parallel computers. Scalable Computing: Practice and Experience, 18(1):17–34, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1231.

Meroufel:2014:ATB

[MB14] Bakhta Meroufel and Ghalem Belalem. Adaptive time-based coordinated checkpointing for cloud computing workflows. Scalable Computing: Practice and Experience, 15(2):??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/981.

Martinez:2009:WSA

[MBSC09] Diego Martínez, Francisco Blanes, Jose Simo, and Alfons Crespotica. Wireless sensors and actuators networks: Characterization and cases study for confined spaces healthcare and control applications. Scalable Computing: Practice and Experience, 10(3):291–305, September 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no3/
SCPE_10_3_06.pdf;   http:
//www.scpe.org/vols/vol10/
no3/SCPE_10_3_06.zip.

Miglani:2017:EET

[MBSS17] Arzoo Miglani, Tarunpreet Bhatia, Gaurav Sharma, and Gulshan Shrivastava. An energy efficient and trust aware framework for secure routing in LEACH for wireless sensor networks. Scalable Computing: Practice and Experience, 18(3):207–218, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1301.

Malinowski:2018:SIP

[MC18] Artur Malinowski and Pawel Czarnul. A solution to image processing with parallel MPI I/O and distributed NVRAM cache. Scalable Computing: Practice and Experience, 19(1): 1–14, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1389.

Morgado:2010:GBI

[MCCC10] Carmen Morgado, José C. Cunha, Nuno Correia, and Jorge Custódio. Group-based interactions for multiuser applications. Scalable Computing: Practice and Experience, 11(1):63–71, March 2010. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol11/no1/SCPE_11_1_
06.pdf;   http://www.scpe.
org/vols/vol11/no1/SCPE_
11_1_06.zip.

Moscato:2012:EMD

[MDA12] Francesco Moscato, Beniamino Di Martino, and Rocco Aversa. Enabling model driven engineering of Cloud services by using mOSAIC ontology. Scalable Computing: Practice and Experience, 13(1):29–44, March 2012. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/765.

Mehra:2019:ECA

[MDA19] Pawan Singh Mehra, Mohammad Najmud Doja, and Bashir Alam. Enhanced clustering algorithm based on fuzzy logic (E-CAFL) for WSN E-CAFL for homogeneous WSN. Scalable Computing: Practice and Experience, 20(1):41–54, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1443.

Meyer:2008:LIS

[ME08] Jan Christian Meyer and Anne C. Elster. Latency impact on spin-lock algorithms for modern shared memory multiprocessors. Scalable Computing: Practice and Experience, 9(3):197–206, September 2008. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol09/no3/SCPE_9_3_
05.pdf;   http://www.scpe.
org/vols/vol09/no3/SCPE_9_
3_05.zip.

Manate:2014:OCR

[MFN14] Bogdan Manate, Teodor-Florin Fortis, and Viorel Negru. Optimizing cloud resources allocation for an Internet of Things architecture. Scalable Computing: Practice and Experience, 15(4):??, ???? 2014. CODEN ???? ISSN 1895-

                                        1767.  URL http://www.
                                        scpe.org/index.php/scpe/
                                        article/view/1055.

Mishev:2018:ISM

[MFPL18] Anastas Mishev, Sonja Filiposka, Ognjen Prnjat, and Ioannis Liabotis. Improving service management for federated resources to support virtual research environments. Scalable Computing: Practice and Experience, 19(2):203–214, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1354.

Mukta:2019:AAB

[MG19] . Mukta and Neeraj Gupta. Analytical available bandwidth estimation in wireless ad-hoc networks considering mobility in 3-Dimensional space. Scalable Computing: Practice and Experience, 20(2):317–334, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1525.

Morrison:2005:RXW

[MHPP05] John P. Morrison, Philip D. Healy, David A. Power, and Keith J. Power. The role of XML within the WebCom metacomputing platform. Scalable Computing: Practice and Experience, 6(1):33–43, March 2005. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol06/SCPE_
6_1_03.pdf;   http://www.
scpe.org/vols/vol06/SCPE_
6_1_03.zip.

Ma:2006:ESO

[MJZ06] Yinglong Ma, Beihong Jin, and Mingquan Zhou. Exploiting shared ontology with default information for Web agents. Scalable Computing: Practice and Experience, 7(1): 11–22, March 2006. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol07/no1/SCPE_7_1_02.pdf;
http://www.scpe.org/vols/
vol07/no1/SCPE_7_1_02.zip.

Manwade:2017:DFA

[MK17] Karveer B. Manwade and Dinesh B. Kulkarni. Data flow analysis of MPI program using dynamic analysis technique with partial execution. Scalable Computing: Practice and Experience, 18(4):375–385, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1335.

Menychtas:2014:SMC

[MKA^{+}14] Andreas Menychtas, Kleopatra Konstanteli, Juncal Alonso, Leire Orue-Echevarria, Jesus Gorronogoitia, George Kousiouris, Christina Santzaridou, Hugo Bruneliere, Bram Pellens, Peter Stuer, Oliver Strauss, Tatiana Senkova, and Theodora Varvarigou. Software modernization and cloudification using the ARTIST migration methodology and framework. Scalable Computing: Practice and Experience, 15(2): ??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/980.

Meghabber:2019:VCA

[MLO^{+}19] Amine Meghabber, Lakhdar Loukil, Richard Olejnik, Abou El Hassan Benyamina, and Abdelkader Aroui. Virtual channel aware scheduling for real time data-flows on network on-chip. Scalable Computing: Practice and Experience, 20(3):495–510, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1423.

Mlynski:2009:IIP

[Mly09] Maciej Mlynski. The influence of the IBM pSeries servers virtualization mechanism on dynamic resource allocation in AIX 5L. Scalable Computing: Practice and Experience, 10(2):189–199, June 2009. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol10/no2/SCPE_10_2_
05.pdf;   http://www.scpe.

org/vols/vol10/no2/SCPE_
10_2_05.zip.

Meyer:2016:ISP

[MM16] Stefan Meyer and John P. Morrison. Impact of single parameter changes on Ceph cloud storage performance. Scalable Computing: Practice and Experience, 17(4):285–298, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1201.

Marzini:2015:TMX

[MMD^{+}15] Emanuel Marzini, Paolo Mori, Sergio Di Bona, Davide Guerri, Marco Lettere, and Laura Ricci. A tool for managing the X1.V1 platform on the cloud. Scalable Computing: Practice and Experience, 16(1): ??, ???? 2015. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/1063.

Mindruta:2012:MAA

[MMNs12] Cristina Mindruta, Victor Ion Munteanu, Viorel Negru, and Calin sandru. Multi-agent architecture in semantic services environment. Scalable Computing: Practice and Experience, 13(1):73–83, March 2012. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/index.php/scpe/
article/view/768.

Mobach:2006:WAB

[MOB06] D. G. A. Mobach, B. J. Overeinder, and F. M. T. Brazier. A WS-agreement based resource negotiation framework for mobile agents. Scalable Computing: Practice and Experience, 7(1):23–36, March 2006. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol07/no1/SCPE_7_1_
03.pdf;   http://www.scpe.
org/vols/vol07/no1/SCPE_7_
1_03.zip.

Moreau:2006:FTD

[Mor06] Luc Moreau. A fault-tolerant directory service for mobile agents based on forwarding pointers. Scalable Computing: Practice and Experience, 7(4): 53–87, December 2006. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol07/no4/SCPE_7_4_
04.pdf;   http://www.scpe.
org/vols/vol07/no4/SCPE_7_
4_04.zip.

Mirandola:2011:QBF

[MP11] Raffaela Mirandola and Pasqualina Potena. A QoS-based framework for the adaptation of service-based systems. Scalable Computing: Practice and Experience, 12(1):63–78, March 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/688.

Marovic:2009:MAB

[MPC09] Branko Marović, Milan Potocnik, and Branislav Cukanović. Multi-application bag of jobs for interactive and on-demand computing. Scalable Computing: Practice and Experience, 10(4):413–418, December 2009. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol10/no4/SCPE_10_4_
07.pdf;   http://www.scpe.
org/vols/vol10/no4/SCPE_
10_4_07.zip.

Mesjasz:2013:ESC

[MPG13] Mariusz Mesjasz, Marcin Paprzycki, and Maria Ganzha. Establishing semantic closeness in an agent-based travel support system. Scalable Computing: Practice and Experience, 14(2):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/844.

Muscalagiu:2012:ETL

[MPN12] Ionel Muscalagiu, Horia Emil Popa, and Viorel Negru. The effect of temporary links in randomly generated networks of constraints. Scalable Computing: Practice and Experience, 13(1):59–71, March 2012. CODEN ???? ISSN 1895-

1767.  URL http://www.
scpe.org/index.php/scpe/
article/view/767.

Moritz:2009:DEW

[MPTG09] Guido Moritz, Steffen Prueter, Dirk Timmermann, and Frank Golatowski. Deployment of embedded Web services in realtime systems. Scalable Computing: Practice and Experience, 10(3):265–275, September 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no3/
SCPE_10_3_04.pdf;   http:
//www.scpe.org/vols/vol10/
no3/SCPE_10_3_04.zip.

Mukherjee:2009:PAS

[MR09] Jayanta Mukherjee and Soumyendu Raha. Power-aware speedup for multithreaded numerical linear algebraic solvers on chip multicore processors. Scalable Computing: Practice and Experience, 10(2):217–228, June 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no2/
SCPE_10_2_07.pdf;   http:
//www.scpe.org/vols/vol10/
no2/SCPE_10_2_07.zip.

Marani:2013:GCP

[MRC13] Alessandro Marani, Marzia Rivi, and Paola Cristofori. Generation of catalogues of PL n-manifolds: Computational aspects on HPC systems. Scalable Computing: Practice and Experience, 14(1): ??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/823.

Mota:2007:CAC

[MRdT07] Sonia Mota, Eduardo Ros, and Francisco de Toro. A computing architecture for correcting perspective distortion in motiondetection based visual systems. Scalable Computing: Practice and Experience, 8(4):387– 394, December 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no4/SCPE_8_4_05.pdf;
 http://www.scpe.org/vols/
vol08/no4/SCPE_8_4_05.zip.

Memon:2016:ESD

[MRM^{+}16] Shahbaz Memon, Morris Riedel, Shiraz Memon, Chris Koeritz, Andrew Grimshaw, and Helmut Neukirchen. Enabling scalable data processing and management through standardsbased job execution and the global federated file system. Scalable Computing: Practice and Experience, 17(2):115–128, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1160.

Momcilovic:2008:PAV

[MS08] Svetislav Momcilovic and Leonel Sousa. Parallel advanced video coding: Motion estimation on multi-cores. Scalable Computing: Practice and Experience, 9(3):207–218, September 2008. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol09/no3/SCPE_9_3_
06.pdf;   http://www.scpe.
org/vols/vol09/no3/SCPE_9_
3_06.zip.

Mehndiratta:2017:DST

[MSS17] Pulkit Mehndiratta, Shelly Sachdeva, and Devpriya Soni. Detection of sarcasm in text data using deep convolutional neural networks. Scalable Computing: Practice and Experience, 18(3):219–228, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1302.

Massa:2008:TOR

[MSST08] Paolo Massa, Kasper Souren, Martino Salvetti, and Danilo Tomasoni. Trustlet, open research on trust metrics. Scalable Computing: Practice and Experience, 9(4):341– 351, December 2008. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol09/no4/SCPE_9_4_10.pdf;
                                         http://www.scpe.org/vols/
                                        vol09/no4/SCPE_9_4_10.zip.

Muscar:2013:JPC

[Mus13] Alex Muscar. Join patterns for concurrent agents (position paper). Scalable Computing: Practice and Experience, 14(3): ??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/859.

Morariu:2008:SSM

[MVV08] Daniel Morariu, Maria Vintan, and Lucian Vintan. Studying SVM method’s scalability using text documents.

Scalable Computing: Practice and Experience, 9(1):1– 10, March 2008. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol09/no1/SCPE_9_1_01.pdf;
http://www.scpe.org/vols/
vol09/no1/SCPE_9_1_01.zip.

Mierzwinski:2011:VSD

[MWWW11] Damian Mierzwinski, Dariusz Walczak, Marcin Wolski, and Marcin Wrzos. Video and sensor data integration in a serviceoriented surveillance system. Scalable Computing: Practice and Experience, 12(1):93–103, March 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/690.

Meneguzzi:2007:IPB

[MZdCML07] Felipe Rech Meneguzzi, Avelino Francisco Zorzo, Michael da Costa Móra, and Michael Luck. Incorporating planning into BDI systems. Scalable Computing: Practice and Experience, 8(1): 15–28, March 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no1/SCPE_8_1_02.pdf;
http://www.scpe.org/vols/
vol08/no1/SCPE_8_1_02.zip.

Navarro:2011:ASR

[NAIC11] Joan Navarro, José Enrique Armendariz-Inigo, and August Climent. An adaptive and scalable replication protocol on power smart grids. Scalable Computing: Practice and Experience, 12(3):351–362, September 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/729.

Naveenkumar:2019:EST

[ND19] M Naveenkumar and S Domnic. Ensemble spatio-temporal distance net for skeleton based action recognition. Scalable Computing: Practice and Experience, 20(3):485–494, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1545.

Nizamic:2012:PBS

[NDGL12] Faris Nizamic, Viktoriya Degeler, Rix Groenboom, and Alexander Lazovik. Policy-based scheduling of Cloud services. Scalable Computing: Practice and Experience, 13(3):187–199, September 2012. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/792.

Negru:2010:ISI

[NFB10] Viorel Negru, Adina Magda Florea, and Costin Badica. Introduction to the special issue: Selected papers from workshops ACSys2009 and IDC’2009. Scalable Computing: Practice and Experience, 11(1): i, March 2010. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/vols/vol11/
no1/vol11no1introduction.
html.

Nemeth:2005:ISI

[NKKV05] Zsolt Németh, Dieter Kranzlmüller, Péter Kacsuk, and Jens Volkert. Introduction to the special issue: Distributed and parallel systems. Scalable Computing: Practice and Experience, 6(2):v, June 2005. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/vols/vol06/
no2/vol06no2introduction.
html.

Ntika:2014:FMS

[NKS14] Marina Ntika, Petros Kefalas, and Ioanna Stamatopoulou. Formal modelling and simulation of a multi-agent nanorobotic drug delivery system. Scalable Computing: Practice and Experience, 15(3):??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/1017.

Nigussie:2013:ISI

[NRG13] Ethiopia Nigussie, Andreas Riener, and Liang Guang. Introduction to the special issue. Scalable Computing: Practice and Experience, 14(3):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/855.

Nikander:2013:EIP

[NRLT13] Pekka Nikander, Vaddina Kameswar Rao, Petri Liuha, and Hannu Tenhunen. ELL-i: An inexpensive platform for fixed things. Scalable Computing: Practice and Experience, 14(3): ??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/857.

Nayyar:2019:SIR

[NRP19] Anand Nayyar, Rudra Rameshwar, and Pijush Kanti Dutta Pramanik. Special issue on recent trends and future of fog and edge computing, services and enabling technologies editorial file. Scalable Computing: Practice and Experience, 20(2): iii–vi, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1558.

Novaes:2005:NDD

[NRS^{+}05] Reynaldo C. Novaes, Paulo Roisenberg, Roque Scheer, Caio Northfleet, J. H. Jornada, and Walfredo Cirne. Non-dedicated distributed environment: A solution for safe and continuous exploitation of idle cycles. Scalable Computing: Practice and Experience, 6(3):107– 115, September 2005. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol06/no3/SCPE_6_3_10.pdf;
 http://www.scpe.org/vols/
vol06/no3/SCPE_6_3_10.zip.

Nigam:2019:BBT

[NSJ^{+}19] Ritu Nigam, Deepak Kumar Sharma, Satbir Jain, Sarthak Gupta, and Shilpa Ghosh. Bonding based technique for message forwarding in social opportunistic network. Scalable Computing: Practice and Experience, 20(1):1–15, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1469.

Nguyen:2011:ALC

[NTH11] Binh Minh Nguyen, Viet Tran, and Ladislav Hluchy. Abstraction layer for Cloud computing. Scalable Computing: Practice and Experience, 12(3):371– 374, September 2011. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/731.

Negru:2013:ISI

[NZ13] Viorel Negru and Daniela Zaharie. Introduction to the special issue. Scalable Computing: Practice and Experience, 14(2):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/840.

Olejnik:2009:LBM

[OAT^{+}09] Richard Olejnik, Iyad Alshabani, Bernard Toursel, Eryk Laskowski, and Marek Tudruj. Load balancing metrics for the SOAJA framework. Scalable Computing: Practice and Experience, 10(4):419–428, December 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no4/
SCPE_10_4_08.pdf;   http:

                                        //www.scpe.org/vols/vol10/
                                        no4/SCPE_10_4_08.zip.

Ourlis:2019:SIA

[OB19] Lazhar Ourlis and Djamel Bellala. SIMD implementation of the Aho–Corasick algorithm using Intel AVX2. Scalable Computing: Practice and Experience, 20(3):563–576, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1572.

Overeinder:2006:GMA

[OdGWB06] B. J. Overeinder, D. R. A. de Groot, N. J. E. Wijngaards, and F. M. T. Brazier. Generative mobile agent migration in heterogeneous environments. Scalable Computing: Practice and Experience, 7(4):89– 99, December 2006. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol07/no4/SCPE_7_4_05.pdf;
                                         http://www.scpe.org/vols/
                                        vol07/no4/SCPE_7_4_05.zip.

Olaru:2010:GBA

[OF10] Andrei Olaru and Adina Magda Florea. A graph-based approach to context matching. Scalable Computing: Practice and Experience, 11(4):393–399, December 2010. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/670.

Olaru:2010:CAE

[OGF10] Andrei Olaru, Cristian Gratie, and Adina Magda Florea. Context-aware emergent behaviour in a MAS for information exchange. Scalable Computing: Practice and Experience, 11(1):33–42, March 2010. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol11/no1/SCPE_11_1_
03.pdf;   http://www.scpe.
org/vols/vol11/no1/SCPE_
11_1_03.zip.

Olsson:2017:NCG

[Ols17] Ronald A. Olsson. A note on correctly gathering results from JR’s concurrent invocation statement. Scalable Computing: Practice and Experience, 18(3):253–260, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1305.

OKane:2006:AGT

[OOM^{+}06] Donal O’Kane, G. M. P.

O’Hare, David Marsh, Song Shen, and Richard Tynan. Agents go traveling. Scalable Computing: Practice and Experience, 7(1):1–10, March 2006. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol07/no1/SCPE_7_1_
01.pdf;   http://www.scpe.
org/vols/vol07/no1/SCPE_7_
1_01.zip.

ODowd:2006:WGM

[OPM06] Padraig J. O’Dowd, Adarsh Patil, and John P. Morrison. WebCom-G and MPICH-G2 jobs. Scalable Computing: Practice and Experience, 7(3): 75–86, September 2006. CO-DEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/
                                        vols/vol07/no3/SCPE_7_3_
                                        07.pdf;   http://www.scpe.
                                        org/vols/vol07/no3/SCPE_7_
                                        3_07.zip.

Oprea:2014:ACA

[Opr14]  Mihaela Oprea.  ABVE-

construct: an agent-based virtual entreprise model for civil engineering. Scalable Computing: Practice and Experience, 15(3):??, ???? 2014. CODEN ???? ISSN 1895-

                                        1767.  URL http://www.
                                        scpe.org/index.php/scpe/
                                        article/view/1018.

Orozco:2007:BRR

[Oro07] Edusmildo Orozco. Book review: Reconfigurable Computing. Accelerating Computation with Field-Programmable Gate Arrays. Scalable Computing: Practice and Experience, 8 (4):437–438, December 2007. CODEN ???? ISSN 1895-

                                        1767. URL http://www.scpe.
                                        org/vols/vol08/no4/BR_101.
                                        html.

Ounapuu:2012:ISI

[ÕS12] Enn Õunapuu and Vlado Stankovski. Introduction to the special issue. Scalable Computing: Practice and Experience, 13(4):??, ???? 2012. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/809.

Ojha:2017:DAM

[OSSS17] Rudra Pratap Ojha, Goutam Sanyal, Pramod Kumar Srivastava, and Kavita Sharma. Design and analysis of modified SIQRS model for performance study of wireless sensor network. Scalable Computing: Practice and Experience, 18(3):229–242, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1303.

Petrova-Antonova:2011:TTW

[PAD11] Dessislava Petrova-Antonova and Aleksandar Dimov. Towards a taxonomy of Web service composition approaches. Scalable Computing: Practice and Experience, 12(4):377–384, December 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/746.

Paprzycki:2005:IE

[Pap05] Marcin Paprzycki. Instead of the editorial. Scalable Computing: Practice and Experience, 6(1):ii, March 2005. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol06/vol06no1editorial.
html.

Paprzycki:2008:E

[Pap08a] Marcin Paprzycki. Editorial. Scalable Computing: Practice and Experience, 9(4):i, December 2008. CODEN ???? ISSN 1895-1767. URL http://www.

                                        scpe.org/vols/vol09/no4/
                                        vol09no4editorial.html.

Paprzycki:2008:ISI

[Pap08b]  Marcin Paprzycki.  Intro-

duction to the special issue: Large scale computations on Grids. Scalable Computing: Practice and Experience, 9(2): i, June 2008. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/vols/vol09/
                                        no2/vol09no2introduction.
                                        html.

Pllana:2008:ISI

[PB08] Sabri Pllana and Siegfried Benkner. Introduction to the special issue: Recent developments in multi-core computing systems. Scalable Computing: Practice and Experience, 9(3):i, September 2008. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/vols/vol09/
                                        no3/vol09no3introduction.
                                        html.

Prabhu:2006:SEG

[PBP06] Devaraya Prabhu, Bill P. Buckles, and Frederick E. Petry. A SIMD environment for genetic algorithms with interconnected subpopulations. Scalable Computing: Practice and Experience, 7(2):65–86, June 2006. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol07/no2/SCPE_7_2_
06.pdf;   http://www.scpe.
org/vols/vol07/no2/SCPE_7_
2_06.zip.

Popescu:2008:LOA

[PBT08] Elvira Popescu, Costin Badica, and Philippe Trigano. Learning objects’ architecture and indexing in WELSA adaptive educational system. Scalable Computing: Practice and Experience, 9(1):11– 20, March 2008. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol09/no1/SCPE_9_1_02.pdf;
 http://www.scpe.org/vols/
vol09/no1/SCPE_9_1_02.zip.

Prasad:2019:IMF

[PBT19] Vivek Kumar Prasad, Madhuri D Bhavsar, and Sudeep Tanwar. Influence of montoring: Fog and edge computing. Scalable Computing: Practice and Experience, 20(2):365– 376, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1533.

Petcu:2008:SOS

[PCMF08] Dana Petcu, Alexandru Cârstea, Georgiana Macariu, and Marc Frîncu. Service-oriented symbolic computing with Sym-Grid. Scalable Computing: Practice and Experience, 9(2): 111–125, June 2008. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/

                                        vol09/no2/SCPE_9_2_04.pdf;
                                         http://www.scpe.org/vols/
                                        vol09/no2/SCPE_9_2_04.zip.

Pena:2007:TAD

[PCRC07] Joaquín Peña, Rafael Corchuelo, and Antonio Ruiz-Cortés. A top down approach for describing the acquaintance organisation of multiagent systems. Scalable Computing: Practice and Experience, 8(1): 49–62, March 2007. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol08/no1/SCPE_8_1_05.pdf;
                                         http://www.scpe.org/vols/
                                        vol08/no1/SCPE_8_1_05.zip.

Pop:2011:HFI

[PCS^{+}11a] Cristina Bianca Pop, Viorica Rozina Chifu, Ioan Salomie, Ramona Bianca Baico, Mihaela Dinsoreanu, and Georgiana Copil. A hybrid firefly-inspired approach for optimal Semantic Web service composition. Scalable Computing: Practice and Experience, 12(3):363–369, September 2011. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/730.

Pop:2011:AIF

[PCS^{+}11b] Cristina Bianca Pop, Viorica Rozina Chifu, Ioan Salomie, Mihaela Dinsoreanu, Tudor David, and Vlad Acretoaie. Ant-inspired framework for automatic Web service composition. Scalable Computing: Practice and Experience, 12(1):

137–151, March 2011. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/693.

Ponnavaikko:2009:ENF

[PD09] Kovendhan Ponnavaikko and Janakiram D. The Edge Node File System: A distributed file system for high performance computing. Scalable Computing: Practice and Experience, 10(1):111–114, March 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no1/
SCPE_10_1_11.pdf;   http:
//www.scpe.org/vols/vol10/
no1/SCPE_10_1_11.zip.

Ponnavaikko:2010:WAD

[PD10] Kovendhan Ponnavaikko and Janakiram Dharanipragada. Wide area distributed file systems — a scalability and performance survey. Scalable Computing: Practice and Experience, 11(3):305–325, September 2010. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/662.

Petcu:2010:ISI

[PDMP10] Dana Petcu, Ewa Deelman, Norbert Meyer, and Marcin Paprzycki. Introduction to the special issue: Grid and cloud computing and their application. Scalable Computing: Practice and Experience, 11(2): i–ii, June 2010. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/vols/vol11/
                                        no2/vol11no2introduction.
                                        html.

Petcu:2005:ECC

[Pet05] Dana Petcu. Editorial: Challenges concerning symbolic computations on Grids. Scalable Computing: Practice and Experience, 6(3):iii–iv, September 2005. CODEN ???? ISSN 1895-1767. URL http://www.

                                        scpe.org/vols/vol06/no3/
                                        vol06no3editorial.html.

Petcu:2006:PRB

[Pet06] Dana Petcu. A parallel rulebased system and its experimental usage in membrane computing. Scalable Computing: Practice and Experience, 7(3):39–49, September 2006. CODEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/
                                        vols/vol07/no3/SCPE_7_3_
                                        04.pdf;   http://www.scpe.
                                        org/vols/vol07/no3/SCPE_7_
                                        3_04.zip.

Petcu:2007:ISI

[Pet07]  Dana Petcu.  Introduction

to the special issue: Practical aspects of large-scale distributed computing. Scalable Computing: Practice and Experience, 8(3):vii, September 2007. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/vols/vol08/
                                        no3/vol08no3introduction.
                                        html.

Petcu:2012:PCS

[Pet12] Dana Petcu. A panorama of Cloud services. Scalable Computing: Practice and Experience, 13(4):??, ???? 2012. CODEN ???? ISSN 1895-

1767.  URL http://www.
scpe.org/index.php/scpe/
article/view/812.

Petcu:2016:ISI

[Pet16] Dana Petcu. Introduction to the special issue on new approaches for infrastructure services. Scalable Computing: Practice and Experience, 17(4): iii, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1199.

Petcu:2011:ISIa

[PG11] Dana Petcu and Alex Galis.

Introduction to the special issue: Selected papers from 1st Workshop on Software Services. Scalable Computing: Practice and Experience, 12(1): iii, March 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/683.

Pop:2011:LBM

[PGD^{+}11] Florin Pop, Marius-Viorel Grigoras, Ciprian Dobre, Ovidiu Achim, and Valentin Cristea. Load-balancing metric for service dependability in large scale distributed environments. Scalable Computing: Practice and Experience, 12(4):391–401, December 2011. CODEN ????

ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/748.

Petcu:2015:PED

[PIP^{+}15] Dana Petcu, Gabriel Iuhasz, Daniel Pop, Domenico Talia, Jesus Carretero, Radu Prodan, Thomas Fahringer, Ivan Grasso, Ramon Doallo, Maria J. Martin, Basilio B. Fraguela, Roman Trobec, Matjaz Depolli, Francisco Almeida Rodriguez, Francisco de Sande, Georges Da Costa, Jean-Marc Pierson, Stergios Anastasiadis, Aristides Bartzokas, Christos Lolis, Pedro Goncalves, Fabrice Brito, and Nick Brown. On processing extreme data. Scalable Computing: Practice and Experience, 16(4):467–490, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1134.

Power:2005:AHM

[PM05] Keith Power and John P. Morrison. Ad hoc metacomputing with Compeer. Scalable Computing: Practice and Experience, 6(1):17–32, March 2005. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol06/SCPE_
6_1_02.pdf;   http://www.
scpe.org/vols/vol06/SCPE_
6_1_02.zip.

Panigrahi:2011:SGP

[PM11a] Satish Panigrahi and Asish Mukhopadhyay. Some geometric problems on OMTSE optoelectronic computer. Scalable Computing: Practice and Experience, 12(2):275–282, June 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/720.

Poniszewska-Maranda:2011:MAC

[PM11b] Aneta Poniszewska-Maranda.

Management of access control in information system based on role concept. Scalable Computing: Practice and Experience, 12(1):35–49, March 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/686.

Poniszewska-Maranda:2011:MAS

[PM11c] Aneta Poniszewska-Maranda.

Multi-agent systems for access control in distributed informations systems. Scalable Computing: Practice and Experience, 12(4):403–415, December 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/749.

Pervez:2016:ISD

[PMH^{+}16] Usman Pervez, Asiah Mahmood, Osman Hasan, Khalid Latif, and Amjad Gawanmeh. Improvement strategies for device interoperability middleware using formal reliability analysis. Scalable Computing: Practice and Experience, 17(3):150–170, ????

2016. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1178.

Pommereau:2005:PNE

[Pom05] Franck Pommereau. Petri nets as executable specifications of high-level timed parallel systems. Scalable Computing: Practice and Experience, 6(4): 71–82, December 2005. CO-DEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/
                                        vols/vol06/no4/SCPE_6_4_
                                        05.pdf;   http://www.scpe.
                                        org/vols/vol06/no4/SCPE_6_
                                        4_05.zip.

Pop:2011:RMB

[Pop11] Florin Pop. Resource management based on gossip monitoring algorithm for large scale distributed systems. Scalable Computing: Practice and Experience, 12(1):21–34, March 2011. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/685.

Pournaras:2015:OSC

[Pou15] Evangelos Pournaras. Overlay service computing — modular and reconfigurable collective adaptive systems. Scalable Computing: Practice and Experience, 16(3):249–270, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1100.

Petcu:2011:ISIb

[PP11] Dana Petcu and Marcin Paprzycki. Introduction to special issue: New directions in Cloud and Grid computing. Scalable Computing: Practice and Experience, 12(2):i, June 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/711.

Pramanik:2019:GSH

[PPC19] Pijush Kanti Dutta Pramanik, Saurabh Pal, and Prasenjit Choudhury. Green and sustainable high-performance computing with Smartphone crowd computing. Scalable Computing: Practice and Experience, 20(2):259–284, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1517.

Paprzycki:2007:ISI

[PS07] Marcin Paprzycki and Niranjan Suri. Introduction to the special issue: Foundational underpinnings for pragmatic agent-based systems. Scalable Computing: Practice and Experience, 8(1): iii, March 2007. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/vols/vol08/
no1/vol08no1introduction.
html.

Panigrahi:2008:MAF

[PS08] S. C. Panigrahi and G. Sahoo. An MIMD algorithm for finding maximum and minimum on OMTSE architecture. Scalable Computing: Practice and Experience, 9(1): 69–75, March 2008. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol09/no1/SCPE_9_1_08.pdf;
 http://www.scpe.org/vols/
vol09/no1/SCPE_9_1_08.zip.

Polgar:2005:EAG

[PSB05] Balázs Polgár, Endre Selényi, and Tamás Bartha. On the extension and applicability of the P -graph modeling paradigm to system-level diagnostic problems. Scalable Computing: Practice and Experience, 6(2): 45–57, June 2005. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol06/no2/SCPE_6_2_05.pdf;
 http://www.scpe.org/vols/
vol06/no2/SCPE_6_2_05.zip.

Pankhuri:2019:PBM

[PSSK19] Pankhuri, Akash Sinha, Gulshan Shrivastava, and Prabhat Kumar. A pattern-based multifactor authentication system. Scalable Computing: Practice and Experience, 20(1):101–112, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1460.

Petcu:2014:PCA

[PV14] Dana Petcu and Athanasios V. Vasilakos. Portability in clouds: approaches and research opportunities. Scalable Computing:

Practice and Experience, 15(3): ??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/1019.

Petcu:2006:HLB

[PVP06] Dana Petcu, Daniel Vizman, and Marcin Paprzycki. Heuristic load balancing for CFD codes executed in heterogeneous computing environments. Scalable Computing: Practice and Experience, 7(2):15– 24, June 2006. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol07/no2/SCPE_7_2_02.pdf;
http://www.scpe.org/vols/
vol07/no2/SCPE_7_2_02.zip.

Petcu:2011:SPWb

[PVP11a] Dana Petcu and Jose Luis Vazquez-Poletti. Selected papers from the 2nd Workshop on Software Services. Scalable Computing: Practice and Experience, 12(3):iii, September 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/723.

Petcu:2011:SPWc

[PVP11b] Dana Petcu and Jose Luis Vazquez-Poletti. Selected papers from the 2nd Workshop on Software Services. Scalable Computing: Practice and Experience, 12(4):iii–??, December 2011. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/745.

Pllana:2009:APM

[PXB09] Sabri Pllana, Fatos Xhafa, and Leonard Barolli. Automatic performance model transformation from a human-intuitive to a machine-efficient form. Scalable Computing: Practice and Experience, 10(1):35–47, March 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no1/
SCPE_10_1_04.pdf;   http:
//www.scpe.org/vols/vol10/
no1/SCPE_10_1_04.zip.

Petcu:2012:ISI

[PZ12] Dana Petcu and Daniela Zaharie. Introduction to the special issue: Agent based systems & semantic software services. Scalable Computing: Practice and Experience, 13(1): iii, March 2012. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/762.

Qin:2005:FCM

[QJZS05] Xiao Qin, Hong Jiang, Yifeng Zhu, and David R. Swanson. A feedback control mechanism for balancing I/O- and memoryintensive applications on clusters. Scalable Computing: Practice and Experience, 6(4):95– 107, December 2005. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol06/no4/SCPE_6_4_07.pdf;

http://www.scpe.org/vols/
vol06/no4/SCPE_6_4_07.zip.

Quitadamo:2006:LSA

[QLC06] Raffaele Quitadamo, Letizia Leonardi, and Giacomo Cabri. Leveraging strong agent mobility for Aglets with the Mobile JikesRVM framework. Scalable Computing: Practice and Experience, 7(4):37– 51, December 2006. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol07/no4/SCPE_7_4_03.pdf;
http://www.scpe.org/vols/
vol07/no4/SCPE_7_4_03.zip.

Rahimi:2006:ISI

[RA06] Shahram Rahimi and Raheel Ahmad. Introduction to the special issue: Software agent technology. Scalable Computing: Practice and Experience, 7(1):v, March 2006. CODEN ???? ISSN 1895-

1767.  URL http://www.
scpe.org/vols/vol07/no1/
vol07no1introduction.html.

Rahimi:2006:ETP

[Rah06] Shahram Rahimi. Editorial:

Toward a performance evaluation methodology for the pi-calculus family of formal modeling languages. Scalable Computing: Practice and Experience, 7(1):i–iii, March 2006. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol07/no1/
vol07no1editorial.html.

Rajasekaran:2018:ISI

[Raj18] Rajkumar Rajasekaran. Introduction to the special issue on mobile cloud applications and challenges. Scalable Computing: Practice and Experience, 19(4): iii, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1499.

Rana:2005:EBG

[Ran05] Omer F. Rana. Editorial: Building Grid communities. Scalable Computing: Practice and Experience, 6(2):iii–iv, June 2005. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol06/no2/
vol06no2editorial.html.

Ristov:2016:IHS

[RCG16] Sasko Ristov, Kiril Cvetkov, and Marjan Gusev. Implementation of a horizontal scalable balancer for dew computing services. Scalable Computing: Practice and Experience, 17(2): 79–90, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1157.

Rochford:2007:ABA

[RCW07] Keith Rochford, Brian Coghlan, and John Walsh. An agentbased approach to Grid service monitoring. Scalable Computing: Practice and Experience, 8(3):281–290, September 2007. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol08/no3/SCPE_8_3_

06.pdf;   http://www.scpe.
org/vols/vol08/no3/SCPE_8_
3_06.zip.

Radu:2014:BSA

[RF14] Serban Radu and Adina Magda Florea. Bargain strategies for agent automated negotiation in an e-business environment. Scalable Computing: Practice and Experience, 15(2):??, ???? 2014. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/979.

Rak:2014:DSC

[RFB^{+}14] Massimiliano Rak, Massimo Ficco, Ermanno Battista, Valentina Casola, and Nicola Mazzocca. Developing secure cloud applications. Scalable Computing: Practice and Experience, 15(1):??, ???? 2014. CODEN ???? ISSN 1895-

1767.  URL http://www.
scpe.org/index.php/scpe/
article/view/965.

Ristov:2018:SII

[Ris18] Sasko Ristov. Special issue on infrastructures and algorithms for scalable computing. Scalable Computing: Practice and Experience, 19(3):iii– iv, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1441.

Radu:2013:MAN

[RKF13] Serban Radu, Eugenia Kalisz, and Adina Magda Florea. A model of automated negotiation based on agents profiles. Scalable Computing: Practice and Experience, 14(1):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/826.

Reddy:2009:HSP

[RLA09] Ravi Reddy, Alexey Lastovetsky, and Pedro Alonso. HeteroPBLAS: A set of Parallel Basic Linear Algebra Subprograms optimized for heterogeneous computational clusters. Scalable Computing: Practice and Experience, 10(2):201–216, June 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no2/
SCPE_10_2_06.pdf;   http:
//www.scpe.org/vols/vol10/
no2/SCPE_10_2_06.zip.

Rahimi:2007:ESA

[RNGG07] Shahram Rahimi, Guruprasad Nagaraja, Lisa Gandy, and Bidyut Gupta. An expert system for analysis of consistency criteria in checkpointing algorithms. Scalable Computing: Practice and Experience, 8(2): 197–208, June 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no2/SCPE_8_2_06.pdf;
 http://www.scpe.org/vols/
vol08/no2/SCPE_8_2_06.zip.

Rana:2005:ABS

[RP05] Omer F. Rana and Line Pouchard. Agent based semantic grids: Research issues and challenges. Scalable Computing: Practice and Experience, 6 (4):83–94, December 2005. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol06/no4/SCPE_6_4_
06.pdf;   http://www.scpe.
org/vols/vol06/no4/SCPE_6_
4_06.zip.

Rodriguez:2008:FTS

[RPM^{+}08] Gabriel Rodríguez, Xoán C.

Pardo, María J. Martín, Patricia González, and Daniel Díaz. A fault tolerance solution for sequential and MPI applications on the Grid. Scalable Computing: Practice and Experience, 9(2):101– 109, June 2008. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol09/no2/SCPE_9_2_03.pdf;
 http://www.scpe.org/vols/
vol09/no2/SCPE_9_2_03.zip.

Roth:2006:DCB

[RPP06] Volker Roth, Jan Peters, and Ulrich Pinsdorf. A distributed content-based search engine based on mobile code and Web service technology. Scalable Computing: Practice and Experience, 7(4):101– 117, December 2006. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol07/no4/SCPE_7_4_06.pdf;
 http://www.scpe.org/vols/
vol07/no4/SCPE_7_4_06.zip.

Reddy:2012:NSM

[RPR12] Neelima Reddy, Raghavendra Prakash, and Ram Mohana Reddy. New sparse matrix storage format to improve the performance of total SPMV time. Scalable Computing: Practice and Experience, 13(2):159–171, June 2012. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/777.

Rivera:2005:GEI

[RS05] Wilson Rivera and J. Seguel. Guest Editors’ introduction. Scalable Computing: Practice and Experience, 6(3):v, September 2005. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/vols/vol06/
                                        no3/vol06no3introduction.
                                        html.

Rose:2007:GWI

[RSM07] Annkatrin Rose, Eric A. Stahlberg, and Iris Meier. Genome-wide identification and comparative analysis of coiledcoil proteins. Scalable Computing: Practice and Experience, 8 (2):167–171, June 2007. CO-DEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/
                                        vols/vol08/no2/SCPE_8_2_
                                        03.pdf;   http://www.scpe.
                                        org/vols/vol08/no2/SCPE_8_
                                        2_03.zip.

Rozman:2008:PMC

[RSM^{+}08] Igor Rozman, Marjan Sterk, Jaka Mocnik, Borut Robic, and Roman Trobec. Performance measurements of computing networks. Scalable Computing: Practice and Experience, 9(2): 143–150, June 2008. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol09/no2/SCPE_9_2_07.pdf;
http://www.scpe.org/vols/
vol09/no2/SCPE_9_2_07.zip.

Rzonca:2009:OEP

[RSS^{+}09] Dariusz Rzońca, Jan Sadolewski, Andrzej Stec, Zbigniew Świder, Bartosz Trybus, and Leszek Trybus. Open environment for programming small controllers according to IEC 61131-3 standard. Scalable Computing: Practice and Experience, 10(3): 325–336, September 2009. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol10/no3/SCPE_10_3_
08.pdf;   http://www.scpe.
org/vols/vol10/no3/SCPE_
10_3_08.zip.

Rak:2015:EPC

[RTV15] Massimiliano Rak, Mauro Turtur, and Umberto Villano. Early prediction of the cost of cloud usage for HPC applications. Scalable Computing: Practice and Experience, 16(3):303–320, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1103.

Rybicki:2016:ESS

[RvSV16] Jedrzej Rybicki and Benedikt von St. Vieth. An extensible software-as-a-service solution for distributed infrastructures. Scalable Computing: Practice and Experience, 17(2):91–102, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1158.

Rahimi:2007:PPA

[RWL07] Shahram Rahimi, Michael Wainer, and Delano Lewis. A performance and programming analysis of Java communication mechanisms in a distributed environment. Scalable Computing: Practice and Experience, 8(2):209–226, June 2007. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol08/no2/SCPE_8_2_
07.pdf;   http://www.scpe.
org/vols/vol08/no2/SCPE_8_
2_07.zip.

Saqib:2007:CAI

[Saq07] Nazar Abbas Saqib. Complexity analysis for 4-input/1-output FPGAs applied to multiplier designs. Scalable Computing: Practice and Experience, 8(4):411–422, December 2007. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol08/no4/
SCPE_8_4_07.pdf;   http:/
/www.scpe.org/vols/vol08/
no4/SCPE_8_4_07.zip.

Sanford:2007:MAB

[SBB^{+}07] J. Sanford, J. Baumgardner, W. Brewer, P. Gibson, and W. ReMine. Mendel’s accountant: a biologically realistic forward-time population genetics program. Scalable Computing: Practice and Experience, 8 (2):147–165, June 2007. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol08/no2/SCPE_8_2_
02.pdf;   http://www.scpe.
org/vols/vol08/no2/SCPE_8_
2_02.zip.

Sharma:2017:ISI

[SBBS17] Kavita Sharma, Suman Bala, Himani Bansal, and Gulshan Shrivastava. Introduction to the special issue on secure solutions for network in scalable computing. Scalable Computing: Practice and Experience, 18(3):iii–iv, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1299.

Schojer:2005:ASM

[SBH05] Peter Schojer, László Böszörményi, and Hermann Hellwagner. An adaptive standard meta-data aware proxy cache. Scalable Computing: Practice and Experience, 6(2):93–104, June 2005. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol06/no2/SCPE_6_2_
09.pdf;   http://www.scpe.
org/vols/vol06/no2/SCPE_6_
2_09.zip.

Schewe:2011:CCC

[SBL^{+}11] Klaus-Dieter Schewe, Károly Bosa, Harald Lampesberger, Ji Ma, Mariam Rady, and Boris Vleju. Challenges in Cloud computing. Scalable Computing: Practice and Experience, 12(4):385–390, December 2011. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/747.

Symeonidis:2007:MAI

[SCKM07] Andreas L. Symeonidis, Kyriakos C. Chatzidimitriou, Dionysios Kehagias, and Pericles A. Mitkas. A multi-agent infrastructure for enhancing ERP system intelligence. Scalable Computing: Practice and Experience, 8(1):101–114, March 2007. CODEN ???? ISSN 1895-1767. URL http://www.

                                        scpe.org/vols/vol08/no1/
                                        SCPE_8_1_09.pdf;   http:/
                                        /www.scpe.org/vols/vol08/
                                        no1/SCPE_8_1_09.zip.

Sridharan:2019:PSI

[SD19] R. Sridharan and S. Domnic. Placement strategy for intercommunicating tasks of an elastic request in fog-cloud environment. Scalable Computing: Practice and Experience, 20(2):335–348, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1526.

Szejnfeld:2010:VTT

[SDK^{+}10] Dawid Szejnfeld, Piotr Dziubecki, Piotr Kopta, Michal Krysinski, Tomasz Kuczynski, Krzysztof Kurowski, Bogdan Ludwiczak, Tomasz Piontek, Dominik Tarnawczyk, Małgorzata Wolniewicz, Piotr Domagalski, Jarosław Nabrzyski, and Krzysztof Witkowski. Vine Toolkit — towards portal based production solutions for scientific and engineering communities with Grid-enabled resources support. Scalable Computing: Practice and Experience, 11 (2):161–172, June 2010. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol11/no2/SCPE_11_2_
07.pdf;   http://www.scpe.
org/vols/vol11/no2/SCPE_
11_2_07.zip.

Sedighi:2014:FTS

[SDZ14] Art Sedighi, Yuefan Deng, and Peng Zhang. Fariness of task scheduling in high performance computing environments. Scalable Computing: Practice and Experience, 15(3): ??, ???? 2014. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/1020.

Schott:2011:GDG

[SE11] Bernhard Schott and Ad Emmen. Green desktop-grids: Scientific impact, carbon footprint, power usage efficiency. Scalable Computing: Practice and Experience, 12(2):257–264, June 2011. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/718.

Sadki:2016:RCP

[SE16] Souad Sadki and Hanan El Bakkali. Resolving conflicting privacy policies in Mhealth based on prioritization. Scalable Computing: Practice and Experience, 17(3):207–226, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1181.

Sedighi:2007:EV

[Sed07] Art Sedighi. Editorial: Virtualization. Scalable Computing: Practice and Experience, 8(2):i–ii, June 2007. CODEN ???? ISSN 1895-

                                        1767.  URL http://www.
                                        scpe.org/vols/vol08/no2/
                                        vol08no2editorial.html.

Sen:2010:EUP

[Sen10] Jaydip Sen. An efficient and user privacy-preserving routing protocol for wireless mesh networks. Scalable Computing: Practice and Experience, 11(4):345–358, December 2010. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/666.

Saleem:2016:ISI

[SG16] Kashif Saleem and Amjad Gawanmeh. Introduction to the special issue on reliability and security of e-health information systems. Scalable Computing:

Practice and Experience, 17(3): iii–iv, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1177.

S:2018:EDS

[SGG18] Prabu S, Gpinath Ganapathy, and Ranjan Goyal. Enhanced data security for public cloud environment with secured hybrid encryption authentication mechanisms. Scalable Computing: Practice and Experience, 19(4):351–360, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1422.

Singh:2019:TRP

[SGJ19] Parminder Singh, Pooja Gupta, and Kiran Jyoti. Triangulation resource provisioning for Web applications in cloud computing: A profit-aware approach. Scalable Computing: Practice and Experience, 20(2):207–222, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1510.

Singh:2019:RAS

[SGJN19] Parminder Singh, Pooja Gupta, Kiran Jyoti, and Anand Nayyar. Research on autoscaling of Web applications in cloud: Survey, trends and future directions. Scalable Computing: Practice and Experience, 20(2):399–432, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1537.

Suri:2006:EMD

[SH06] Niranjan Suri and Henry Hexmoor. Editorial: Mobility in distributed systems. Scalable Computing: Practice and Experience, 7(4):i–ii, December 2006. CODEN ???? ISSN 1895-1767. URL http://www.

                                        scpe.org/vols/vol07/no4/
                                        vol07no4editorial.html.

Slizik:2008:SMN

[SH08] Peter Slízik and Ladislav Hluchý. Simulation models of natural disasters and service-based visualization of their datasets in the medigrid project. Scalable Computing: Practice and Experience, 9(2): 95–100, June 2008. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol09/no2/SCPE_9_2_02.pdf;
                                         http://www.scpe.org/vols/
                                        vol09/no2/SCPE_9_2_02.zip.

Shampine:2009:VSO

[Sha09] L. F. Shampine. Vectorized solution of ODEs in Matlab. Scalable Computing: Practice and Experience, 10(4):337– 345, December 2009. CO-DEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/
                                        vols/vol10/no4/SCPE_10_4_
                                        01.pdf;   http://www.scpe.
                                        org/vols/vol10/no4/SCPE_
                                        10_4_01.zip.

Srinivas:2007:DMD

[SJ07] A. Vijay Srinivas and D. Janakiram. Data management in distributed systems: A scalability taxonomy. Scalable Computing: Practice and Experience, 8(1): 115–130, March 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no1/SCPE_8_1_10.pdf;
http://www.scpe.org/vols/
vol08/no1/SCPE_8_1_10.zip.

Sanders:2010:CEI

[SK10] Jason Sanders and Edward Kandrot. CUDA by Example: an Introduction to Generalpurpose GPU Programming. Addison-Wesley, Reading, MA, USA, 2010. ISBN 0-13-138768-5. xix + 290 pp. LCCN QA76.76.A65.

Shrivastava:2017:PAA

[SK17] Gulshan Shrivastava and Prabhat Kumar. Privacy analysis of Android applications: Stateof-art and literary assessment. Scalable Computing: Practice and Experience, 18(3):243–252, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1304.

Sharma:2019:ERE

[SK19] Deepak Kumar Sharma and Deepika Kukreja. Establishing reliability for efficient routing in opportunistic networks. Scalable Computing: Practice and Experience, 20(1):27–40, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1479.

Simecek:2014:TBS

[SL14] I. Simecek and D. Langr. Treebased space efficient formats for storing the structure of sparse matrices. Scalable Computing: Practice and Experience, 15(1):??, ???? 2014. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/962.

Seyfari:2017:PPV

[SLK17] Yousef Seyfari, Shahriar Lotfi, and Jaber Karimpour. PVL: Parallelization and vectorization of affine perfectly nestedloops considering data locality on short-vector multicore processors using intrinsic vectorization. Scalable Computing: Practice and Experience, 18(1): 67–89, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1234.

Singh:2008:MAS

[SMGC08] David E. Singh, Alejandro Miguel, Félix García, and Jesús Carretero. Mobile agent systems integration into parallel environments. Scalable Computing: Practice and Experience, 9(2):127–134, June 2008. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol09/no2/SCPE_9_2_
05.pdf;   http://www.scpe.

org/vols/vol09/no2/SCPE_9_
2_05.zip.

Scaife:2006:EPP

[SMH06a] Norman Scaife, Greg Michaelson, and Susumu Horiguchi. Empirical parallel performance prediction from semanticsbased profiling. Scalable Computing: Practice and Experience, 7(3):1–8, September 2006. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol07/no3/SCPE_7_3_
01.pdf;   http://www.scpe.
org/vols/vol07/no3/SCPE_7_
3_01.zip.

Scaife:2006:PSM

[SMH06b] Norman Scaife, Greg Michaelson, and Susumu Horiguchi. Parallel Standard ML with skeletons. Scalable Computing: Practice and Experience, 7(2): 87–92, June 2006. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol07/no2/SCPE_7_2_07.pdf;
http://www.scpe.org/vols/
vol07/no2/SCPE_7_2_07.zip.

Singh:2019:DTS

[SNKS19] Simar Preet Singh, Anand Nayyar, Harpreet Kaur, and Ashu Singla. Dynamic task scheduling using balanced VM allocation policy for fog computing platforms. Scalable Computing: Practice and Experience, 20(2):433–456, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1538.

Slota:2010:PLB

[SNP^{+}10] Renata Slota, Darin Nikolow, Stanislaw Polak, Marcin Kuta, Mariusz Kapanowski, Kornel Skalkowski, Marek Pogoda, and Jacek Kitowski. Prediction and load balancing system for distributed storage. Scalable Computing: Practice and Experience, 11(2):121–130, June 2010. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol11/no2/SCPE_11_2_
03.pdf;   http://www.scpe.
org/vols/vol11/no2/SCPE_
11_2_03.zip.

Song:2009:CLT

[Son09] Ha Yoon Song. Combination of localization techniques for mobile sensor network for precise localization. Scalable Computing: Practice and Experience, 10(3):307–324, September 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no3/
SCPE_10_3_07.pdf;   http:
//www.scpe.org/vols/vol10/
no3/SCPE_10_3_07.zip.

Sierra-Pajuelo:2016:POA

[SPPGP16] Fernando Sierra-Pajuelo, Abel Paz-Gallardo, and Antonio Plaza. Performance optimizations for an automatic target generation process in hyperspectral analysis. Scalable Computing: Practice and Experience, 17(1):1–12, ????

2016. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1146.

Sharif:2010:EEE

[SPR10] Atif Sharif, Vidyasagar M. Potdar, and A. J. D. Rathnayaka. ERCTP: End-to-end reliable and congestion aware transport layer protocol for heterogeneous WSN. Scalable Computing: Practice and Experience, 11(4):359–371, December 2010. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/667.

Siddavatam:2017:TVM

[SPSK17] Irfan A. Siddavatam, Sachin Parekh, Tanay Shah, and Faruk Kazi. Testing and validation of Modbus/TCP protocol for secure SCADA communication in CPS using formal methods. Scalable Computing: Practice and Experience, 18(4):313–330, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1331.

Shanmugasundaram:2018:SHM

[SR18a] Vengadeswaran Shanmugasundaram and Balasundaram Sadhu Ramakrishnan. Significance of hierarchical and Markov clustering in grouping aware data placement for data intensive applications having interest locality. Scalable Computing: Practice and Experience, 19(3):245– 258, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1375.

Somula:2018:SMC

[SR18b] Rama Subbareddy Somula and Sasikala R. A survey on mobile cloud computing: Mobile computing + cloud computing (MCC = MC + CC). Scalable Computing: Practice and Experience, 19(4):309–337, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1411.

Somula:2018:RRL

[SR18c] Ramasubbareddy Somula and Sasikala R. Round robin with load degree: An algorithm for optimal cloudlet discovery in mobile cloud computing. Scalable Computing: Practice and Experience, 19(1):39– 52, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1392.

S:2018:EAS

[SS18] Karthi S and Prabu S. Execution analysis of spatial data storage indexing on cloud environment. Scalable Computing: Practice and Experience, 19(4):339–349, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1421.

Sanchez:2018:ESB

[SSBO18] David Sanchez, Oswaldo Solarte, Victor Bucheli, and Hugo Ordonez. Evaluating the scalability of big data frameworks. Scalable Computing: Practice and Experience, 19(3):301–307, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1402.

Steghoefer:2015:CPM

[SSE^{+}15] Jan-Philipp Steghoefer, Hella Seebach, Benedikt Eberhardinger, Michael Huebschmann, and Wolfgang Reif. Combining PosoMAS method content with Scrum: Agile software engineering for open selforganising systems. Scalable Computing: Practice and Experience, 16(4):333–354, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1127.

Sharma:2019:ZBE

[SSM19] Rajan Sharma, Balwinder Singh Sohi, and Nitin Mittal. Zonebased energy efficient routing protocols for wireless sensor networks. Scalable Computing: Practice and Experience, 20(1): 55–70, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1432.

Schmidt:2008:BRA

[SSS08] Kay-Uwe Schmidt, Roland Stühmer, and Ljiljana Stojanovic. From business rules to application rules in rich Internet applications. Scalable Computing: Practice and Experience, 9(4):329–340, December 2008. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol09/no4/SCPE_9_4_09.pdf;
                                         http://www.scpe.org/vols/
                                        vol09/no4/SCPE_9_4_09.zip.

Stocker:2008:SKT

[SST08] Alexander Stocker, Markus Strohmaier, and Klaus Tochtermann. Studying knowledge transfer with Weblogs in small and medium enterprises: An exploratory case study. Scalable Computing: Practice and Experience, 9(4):243–258, December 2008. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol09/no4/SCPE_9_4_02.pdf;
                                         http://www.scpe.org/vols/
                                        vol09/no4/SCPE_9_4_02.zip.

Skala:2016:ISI

[STA16] Karolj Skala, Roman Trobec, and Enis Afgan. Introduction to the special issue on distributed computing with applications in bioengineering. Scalable Computing: Practice and Experience, 17(2):iii–iv, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1155.

Stpiczynski:2006:ISI

[Stp06] Przemyslaw Stpiczyński. Introduction to the special issue: Large scale computations on Grids. Scalable Computing: Practice and Experience, 7(2): v, June 2006. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/vols/vol07/
no2/vol07no2introduction.
html.

Snyder:2015:SSO

[SV15] Paul L. Snyder and Giuseppe Valetto. SODAP: Self-organized topology protection for superpeer P2P networks. Scalable Computing: Practice and Experience, 16(3):271–288, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1101.

Singh:2019:BBC

[SV19] Nikita Singh and Manu Vardhan. Blockchain based echeque clearing framework. Scalable Computing: Practice and Experience, 20(3):511–526, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1506.

Swany:2005:NSC

[SW05] Martin Swany and Rich Wolski.

Network scheduling for computational Grid environments. Scalable Computing: Practice and Experience, 6(3):85– 94, September 2005. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol06/no3/SCPE_6_3_08.pdf;

                                         http://www.scpe.org/vols/
                                        vol06/no3/SCPE_6_3_08.zip.

Sadi:2016:CAA

[SY16] Samy Sadi and Belabbas Yagoubi. Communicationaware approaches for transparent checkpointing in cloud computing. Scalable Computing: Practice and Experience, 17(3):251–270, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1184.

Thain:2005:PTU

[TL05] Douglas Thain and Miron Livny. Parrot: Transparent user-level middleware for dataintensive computing. Scalable Computing: Practice and Experience, 6(3):9–18, September 2005. CODEN ???? ISSN 1895-1767. URL http://www.

                                        scpe.org/vols/vol06/no3/
                                        SCPE_6_3_02.pdf;   http:/
                                        /www.scpe.org/vols/vol06/
                                        no3/SCPE_6_3_02.zip.

Taher:2012:ECA

[TNL^{+}12] Yehia Taher, Dinh Khoa Nguyen, Francesco Lelli, Willem-Jan van den Heuvel, and Mike Papazoglou. On engineering Cloud applications — state of the art, shortcomings analysis, and approach. Scalable Computing: Practice and Experience, 13(3):215–231, September 2012. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/794.

Tracy:2018:CMC

[TOC18] Fred Thomas Tracy, Thomas C.

Oppe, and Maureen K. Corcoran. A comparison of MPI and co-array FORTRAN for large finite element variably saturated flow simulations. Scalable Computing: Practice and Experience, 19(4):423–432, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1468.

Tray:2015:SSU

[TOWC15] Fred T. Tray, Thomas C. Oppe, William A. Ward, and Maureen K. Corcoran. A scalability study using supercomputers for huge finite element variably saturated flow simulations. Scalable Computing: Practice and Experience, 16(2):153–168, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1087.

Tosic:2008:AAP

[TRI08] Sasa Tosić, Milos Radovanović, and Mirjana Ivanović. APP: Agent planning package. Scalable Computing: Practice and Experience, 9(1):39–49, March 2008. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol09/no1/
SCPE_9_1_05.pdf;   http:/
/www.scpe.org/vols/vol09/
no1/SCPE_9_1_05.zip.

Tudruj:2009:ISI

[Tud09] Marek Tudruj. Introduction to the special issue. Scalable Computing: Practice and Experience, 10(2):i–ii, June 2009. CODEN ???? ISSN 1895-1767. URL http:/

                                        /www.scpe.org/vols/vol10/
                                        no2/vol10no2introduction.
                                        html.

Thoai:2005:APS

[TV05] Nam Thoai and Jens Volkert. An approach to providing small-waiting time during debugging message-passing programs. Scalable Computing: Practice and Experience, 6(2): 1–12, June 2005. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol06/no2/SCPE_6_2_01.pdf;
                                         http://www.scpe.org/vols/
                                        vol06/no2/SCPE_6_2_01.zip.

Tracy:2016:SSI

[TWC16] Fred T. Tracy, Lucas A. Walshire, and Maureen K. Corcoran. Sensitivity study of input parameters for seepage flow simulations using parallel computers. Scalable Computing: Practice and Experience, 17(1): 47–60, ???? 2016. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1149.

Venticinque:2010:MHP

[VAM^{+}10] Salvatore Venticinque, Rocco Aversa, Beniamino Di Martino, Renato Donini, Sergio Briguglio, and Gregorio Vlad.

Management of high performance scientific applications using mobile agents based services. Scalable Computing: Practice and Experience, 11 (2):149–159, June 2010. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol11/no2/SCPE_11_2_
06.pdf;   http://www.scpe.
org/vols/vol11/no2/SCPE_
11_2_06.zip.

Varga:2015:IPB

[Var15] Laszlo Varga. On intentionpropagation-based prediction in autonomously self-adapting navigation. Scalable Computing: Practice and Experience, 16(3):221–232, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1098.

Villari:2013:EFF

[VCFP13] Massimo Villari, Antonio Celesti, Maria Fazio, and Antonio Puliafito. Evaluating a file fragmentation system for multi-provider cloud storage. Scalable Computing: Practice and Experience, 14(4):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/index.php/
scpe/article/view/932.

VanDamme:2008:DLC

[VCV08] Céline Van Damme, Tanguy Coenen, and Eddy Vandijck. Deriving a lightweight corporate ontology form a folksonomy: a methodology and its possible applications. Scalable Computing: Practice and Experience, 9(4):293– 301, December 2008. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol09/no4/SCPE_9_4_06.pdf;
                                         http://www.scpe.org/vols/
                                        vol09/no4/SCPE_9_4_06.zip.

R:2018:PSS

[VE18] Belfin R V and Grace Mary Kanaga E. Parallel seed selection method for overlapping community detection in social network. Scalable Computing: Practice and Experience, 19(4):375–385, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1429.

Vudragovic:2018:VSD

[VIJ^{+}18] Dusan Vudragovic, Luka Ilic, Petar Jovanovic, Slobodan Nickovic, Aleksandar Bogojevic, and Antun Balaz. VI-SEEM DREAMCLIMATE service. Scalable Computing: Practice and Experience, 19(2): 215–221, ???? 2018. CO-DEN ???? ISSN 1895-

                                        1767.  URL https://www.
                                        scpe.org/index.php/scpe/
                                        article/view/1396.

Verma:2019:AAM

[VK19] Jyotsna Verma and Nishtha Kesswani. AMIGM: Animal migration inspired group mobility model for mobile ad hoc networks. Scalable Computing: Practice and Experience, 20(3):577–590, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1574.

vonLaszewski:2005:TRG

[vLAV05] G. von Laszewski, B. K.

Alunkal, and I. Veljkovic. Toward reputable Grids. Scalable Computing: Practice and Experience, 6(3):95–106, September 2005. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol06/no3/SCPE_6_3_09.pdf;
http://www.scpe.org/vols/
vol06/no3/SCPE_6_3_09.zip.

Valero-Lara:2017:THE

[VLKJ17] Pedro Valero-Lara, Ezhilmathi Krishnasamy, and Johan Jansson. Towards HPC-embedded. case study: Kalray and message-passing on NoC. Scalable Computing: Practice and Experience, 18(2):151–160, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1287.

Valero-Lara:2016:MTC

[VLNP^{+}16] Pedro Valero-Lara, Poornima Nookala, Fernando L. Pelayo, Johan Jansson, Serapheim Dimitropoulos, and Ioan Raicu. Many-task computing on many-core architectures. Scalable Computing: Practice and Experience, 17(1):32–46, ???? 2016. CODEN ????

ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1148.

Valero-Lara:2017:ISI

[VLZAP17] Pedro Valero-Lara, Mawussi Zounon, Maksims Abalenkovs, and Fernando L. Pelayo. Introduction to the special issue on high performance computing solutions for complex problems. Scalable Computing: Practice and Experience, 18(2): iii–iv, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1282.

Varghese:2010:ECA

[VM10] Blesson Varghese and Gerard McKee. Exploring carrier agents in swarm-array computing. Scalable Computing: Practice and Experience, 11 (1):53–62, March 2010. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol11/no1/SCPE_11_1_
05.pdf;   http://www.scpe.
org/vols/vol11/no1/SCPE_
11_1_05.zip.

vanNieuwpoort:2005:SSE

[vNMKB05] Rob van Nieuwpoort, Jason Maassen, Thilo Kielmann, and Henri E. Bal. Satin: Simple and efficient Java-based Grid programming. Scalable Computing: Practice and Experience, 6(3): 19–32, September 2005. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol06/no3/SCPE_6_3_

03.pdf;   http://www.scpe.
org/vols/vol06/no3/SCPE_6_
3_03.zip.

Vallurupalli:2007:ABM

[VP07] Vaishali Vallurupalli and Carla Purdy. Agent-based modeling and simulation of biomolecular reactions. Scalable Computing: Practice and Experience, 8(2): 185–196, June 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no2/SCPE_8_2_05.pdf;
 http://www.scpe.org/vols/
vol08/no2/SCPE_8_2_05.zip.

Vazquez-Poletti:2012:ISI

[VPPL12] José Luis Vázquez-Poletti, Dana Petcu, and Francesco Lelli. Introduction to the special issue: Selected papers from the international workshop on clouds for business and business for clouds. Scalable Computing: Practice and Experience, 13(3): iii, September 2012. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/791.

Velliangiri:2019:TFB

[VRK19] S. Velliangiri, R. Rajagopal, and P. Karthikeyan. Trust factor based key distribution protocol in hybrid cloud environment. Scalable Computing: Practice and Experience, 20(2):349–364, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1529.

Wang:2008:DWN

[WH08] Yang Wang and Thomas Hornung. Deep Web navigation by example. Scalable Computing: Practice and Experience, 9(4): 281–292, December 2008. CO-DEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.org/
                                        vols/vol09/no4/SCPE_9_4_
                                        05.pdf;   http://www.scpe.
                                        org/vols/vol09/no4/SCPE_9_
                                        4_05.zip.

Wiandt:2015:ACM

[WKS15] Bernat Wiandt, Andras Kokuti, and Vilmos Simon. Application of collective movement in real life scenarios: Overview of current flocking solutions. Scalable Computing: Practice and Experience, 16(3):233–248, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1099.

Wallace:2014:RCE

[WMVP14] Richard M. Wallace, Patrick Martin, and Jose Luis Vazquez-Poletti. Regulated conditionevent matrices for cloud environments. Scalable Computing: Practice and Experience, 15(2):??, ???? 2014. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/index.
                                        php/scpe/article/view/982.

Woo:2005:SAJ

[Woo05] Jongwook Woo. Static analysis for Java with alias representation reference-set in high-performance computing. Scalable Computing: Practice and Experience, 6(1):125–139, March 2005. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/vols/vol06/
SCPE_6_1_10.pdf;   http:/
/www.scpe.org/vols/vol06/
SCPE_6_1_10.zip.

Woo:2007:BRC

[Woo07] Jongwook Woo. Book review: The Comparison of J2EE and .NET for Enterprise Information Systems. Scalable Computing: Practice and Experience, 8(1):131– 140, March 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no1/SCPE_8_1_11.pdf;
http://www.scpe.org/vols/
vol08/no1/SCPE_8_1_11.zip.

Wei:2007:DTD

[WS07] Xinfa Wei and Kaoru Sezaki.

DHR-trees: A distributed multidimensional indexing structure for P2P systems. Scalable Computing: Practice and Experience, 8(3):291–300, September 2007. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol08/no3/SCPE_8_3_07.pdf;
http://www.scpe.org/vols/
vol08/no3/SCPE_8_3_07.zip.

Wu:2009:PAO

[WTLS09] Xingfu Wu, Valerie Taylor, Charles Lively, and Sameh Sharkawi. Performance analysis and optimization of parallel scientific applications on CMP clusters. Scalable Computing: Practice and Experience, 10(1):61–74, March 2009. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol10/no1/SCPE_10_1_
06.pdf;   http://www.scpe.
org/vols/vol10/no1/SCPE_
10_1_06.zip.

Weyns:2005:SDT

[WTV05] Danny Weyns, Eddy Truyen, and Pierre Verbaeten. Serialization of distributed threads in Java. Scalable Computing: Practice and Experience, 6(1):81–98, March 2005. CO-DEN ???? ISSN 1895-

1767.  URL http://www.
scpe.org/vols/vol06/SCPE_
6_1_07.pdf;   http://www.
scpe.org/vols/vol06/SCPE_
6_1_07.zip.

Wang:2006:CPM

[WZS06] Kai Wang, Jun Zhang, and Chi Shen. A class of parallel multilevel sparse approximate inverse preconditioners for sparse linear systems. Scalable Computing: Practice and Experience, 7(2):93– 106, June 2006. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol07/no2/SCPE_7_2_08.pdf;
 http://www.scpe.org/vols/
vol07/no2/SCPE_7_2_08.zip.

Xhafa:2009:ISI

[XB09] Fatos Xhafa and Leonard Barolli. Introduction to the special issue: Simulation in emergent computational systems. Scalable Computing: Practice and Experience, 10(1): i–ii, March 2009. CODEN ???? ISSN 1895-1767. URL http:

//www.scpe.org/vols/vol10/
no1/vol10no1introduction.
html.

Yang:2009:PEW

[YBI^{+}09] Tao Yang, Leonard Barolli, Makoto Ikeda, Giuseppe De Marco, and Arjan Durresi. Performance evaluation of a wireless sensor network for mobile and stationary event cases considering routing efficiency and goodput metrics. Scalable Computing: Practice and Experience, 10(1):99–109, March 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no1/
SCPE_10_1_09.pdf;   http:
//www.scpe.org/vols/vol10/
no1/SCPE_10_1_09.zip.

Yang:2017:ASD

[YBW^{+}17] Zhengyu Yang, Janki Bhimani, Jiayin Wang, David Evans, and Ningfang Mi. Automatic and scalable data replication manager in distributed computation and storage infrastructure of cyber-physical systems. Scalable Computing: Practice and Experience, 18(4):291–312, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1330.

Yang:2018:AAD

[YBY^{+}18] Zhengyu Yang, Janki Bhimani, Yi Yao, Cho-Hsien Lin, Jiayin Wang, Ningfang Mi, and Bo Sheng. AutoAdmin: Automatic and dynamic resource reservation admission control in Hadoop YARN clusters. Scalable Computing: Practice and Experience, 19(1):53–68, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1393.

Yang:2017:ATM

[YRL17] Zhiliu Yang, Zachary I. Rauen, and Chen Liu. Automatic tuning on many-core platform for energy efficiency via support vector machine enhanced differential evolution. Scalable Computing: Practice and Experience, 18(2):117–132, ???? 2017. CODEN ???? ISSN 1895-1767. URL https:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/1284.

Zalewski:2009:ISI

[Zal09] Janusz Zalewski. Introduction to the special issue: Realtime distributed systems and networks. Scalable Computing: Practice and Experience, 10(3):iii, September 2009. CODEN ???? ISSN 1895-1767. URL http:/

                                        /www.scpe.org/vols/vol10/
                                        no3/vol10no3introduction.
                                        html.

Zahra:2019:ASS

[ZC19] Syed Rameem Zahra and Mohammad Ahsan Chishti. Assesing the services, security threaths, challenges and solutions in the Internet of Things. Scalable Computing: Practice and Experience, 20(3):457–484, ???? 2019. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1544.

Zhang:2011:HMR

[ZD11] Chen Zhang and Hans De Sterck. HBaseSI: Multi-row distributed transactions with global strong snapshot isolation on Clouds. Scalable Computing: Practice and Experience, 12(2):209–226, June 2011. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/index.php/scpe/
article/view/699;   http:
//www.scpe.org/index.php/
scpe/article/view/715.

Zeginis:2013:UCM

[ZDB^{+}13] Dimitris Zeginis, Francesco D’Andria, Stefano Bocconi, Jesus Gorronogoitia Cruz, Oriol Collell Martin, Panagiotis Gouvas, Giannis Ledakis, and Konstantinos A. Tarabanis. A user-centric multi-PaaS application management solution for hybrid multi-cloud scenarios. Scalable Computing: Practice and Experience, 14(1): ??, ???? 2013. CODEN ???? ISSN 1895-1767. URL http:

                                        //www.scpe.org/index.php/
                                        scpe/article/view/824.

Zhang:2007:OBP

[Zha07]  Yu Zhang.  Observation-

based proactive communication in multi-agent teamwork. Scalable Computing: Practice and Experience, 8(1):63– 77, March 2007. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol08/no1/SCPE_8_1_06.pdf;
                                         http://www.scpe.org/vols/
                                        vol08/no1/SCPE_8_1_06.zip.

Zhang:2006:DMM

[ZL06] Yang Zhang and Edward A. Luke. Dynamic memory management in the Loci framework. Scalable Computing: Practice and Experience, 7(3):27– 37, September 2006. CODEN ???? ISSN 1895-1767. URL

                                        http://www.scpe.org/vols/
                                        vol07/no3/SCPE_7_3_03.pdf;
                                         http://www.scpe.org/vols/
                                        vol07/no3/SCPE_7_3_03.zip.

Zefreh:2016:TST

[ZLKK16] Ebrahim Zarei Zefreh, Shahriar Lotfi, Leyli Mohammad Khanli, and Jaber Karimpour. Tiling and scheduling of three-level perfectly nested loops with dependencies on heterogeneous systems. Scalable Computing: Practice and Experience, 17(4):331–350, ???? 2016. CODEN ???? ISSN 1895-

                                        1767.  URL https://www.
                                        scpe.org/index.php/scpe/
                                        article/view/1205.

Zimeo:2013:DSS

[ZOBC13] Eugenio Zimeo, Gianfranco Oliva, Fabio Baldi, and Alfonso Caracciolo. Designing a scalable social e-commerce application. Scalable Computing: Practice and Experience, 14(2):??, ???? 2013. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/index.
php/scpe/article/view/845.

Zedadra:2015:EAA

[ZSJF15] Ouarda Zedadra, Hamid Seridi, Nicolas Jouandeau, and Giancarlo Fortino. An energyaware algorithm for large scale foraging systems. Scalable Computing: Practice and Experience, 16(4):449–466, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1133.